A Complete Bibliography of Publications in Technology and Culture: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

23 May 2024 Version 1.03

Title word cross-reference

e [Bel21, Gui22, Nèg22a, RR21b, Ver22b, Wil21b].

&c [Wil20].

12 [Low20]. 15th [Cha21c]. 1600 [Gag21]. 16th [Di 21, Gui22, Nèg22a]. 1750 [Ver23]. 1790s [Cha22c]. 17th [Gou21]. 17th-early [Gou21]. 1800 [And21]. 1830 [Cop24]. 1830s [Hom22]. 1838 [Ach22]. 1848 [Ros20]. 1850 [Chu23, Die22, Gie22]. 1850s [Cha22c]. 1865 [Bri20]. 1870 [Plo23]. 1870s [Fer24]. 1880 [Jon22a]. 1880s [Bon21b, Lev21, Per23b]. 1890s [Mon23]. 1895 [Bec23]. 18th [Bel21, Cha21c, Ryd20]. 1900 [Kvi23a, Ten22]. 1900s [Cha22b, Got20]. 1906 [Sco23]. 1910 [Tom21]. 1910s [Per23b, Yi22]. 1914 [Lag24]. 1915 [Mat24]. 1920 [Cat21, Dav23b]. 1920s [Eri21, Yi22]. 1930 [Fou23, MF20, Mcg20]. 1930s [Ber23, HG24, Mel23]. 1935 [Rob20]. 1939 [Cam21a, Rog21, Sch20d]. 1940 [Ang23, Cie21, Gro21, Mai23, PZ23, Suá21].

1940s [Chu22b, Fer24]. 1942 [Han22, SG21]. 1945 [Gid24, Gum24, Yan20, vdS22]. 1946 [Gue22b]. 1949 [Dan21]. 1950 [KdS24, Pea22, Wen24]. 1950s [AH20, Cig23, Ign22, Koc23, Mon22, Mor24, KM23]. 1952 [Coh21]. 1955 [Zen22]. 1960 [Cie23, Ram22a]. 1960s [Lau21, McC23, San20, Zan22, De 23a]. 1970 [Cha21a]. 1970s [Ign22, Lev21, Mor24, Tat21, Aug20, Hof23]. 1973 [Wal21a]. 1975 [Her20, Wol23]. 1980 [Cha20, Sie22]. 1980s [Koc23, Mil24, Yoo20, Lau21]. 1982 [Zai23]. 1986 [Gut23]. 1986/87 [Gut23]. 1990 [Poo23]. 1990s [Kif24a, Yoo20]. 1991 [Kra23]. 1992 [Mac24]. 1998 [Mel23]. 19th [Di 21, Gab22, Kri24, Mic21, Ryd20, Ver22b, Wil21b].

2-D [Ott21b]. 2000 [Del21, Rom22]. 2000s [Got20]. 2002 [Kui22]. 2010s [San20]. 2015 [Dis21a]. 2019 [Mor22]. 2020 [BBC^{+}21, Fri22b, Han23, Pie22]. 2023 [Ano23e]. 2024 [Ano24b, Ano24a]. 2050 [Tri20]. 20th [Als20, Bel21, Gab22, Gou21, Mic21, Wij22]. 21st [Nèg22a, Wil21b]. 21th [Ver22b].

3 [Ban21]. 3-D [Ban21].

47th [BBC^{+}21]. 4Mb [Yoo20].

66 [PW24]. 68 [Zai20].

7-Day [Røs22b]. 70s [Ber23]. 74 [Jam23]. 75 [Syl22].

80s [Chu22b]. 87 [Gut23].

âge [HP21a].

Aaron [Cop24, NOB^{+}20]. Aashique [Mel23]. Abbate [Sch23a]. Abel [Coa24, Low23]. Abena [Ade21]. Abigail [Rup24]. Abnet [Rus24b]. Absent [Tho20]. Absorption [Bea21]. Abstractions [Sch23a]. Acceleration [Het23]. Access [Han23]. Accessible [Zdr21]. Accident [Ion21]. Accidents [Her20]. Accounting [Sta20]. Accra [Har20b]. Accumulated [Bar21b]. Achim [See22]. Acid [Ste20]. Acoustic [Fin22, Lin24a]. Acoustics [Bea21]. acque [Ber24, Bia22]. Acquired [His23a]. Across [Tho24, Ber21e, Bro21b, Ger21c, San20]. Act [Hil23b]. Action [KMÅ20, KB22b]. Activism [Mil22]. actores [Zul24]. Actors [Kai21, Zul24]. AD [Whi23]. Adam [Bar22, Hay22, Swa21]. Adams [Lan23b, Vai23]. Adar [Opp22]. Addis [Cha22a]. Admiralty [Har23a]. Advantage [Ros20]. Advocacy [BMSS24]. Aebischer [Kri21]. Aeroplane [Mel23]. Aesthetic [Fou23]. Aesthetics [Han20, McC22]. Affected [Vik20]. Afkhami [Cla21]. Africa [Ade21, And22, Cos20, Moo20, Bee23, Den20a, Kif24b, Mor24, d’A20, Duq24]. African [BF24, Gro23, Map23, Fil23]. Africanizing [Twa20a]. Afrika [And22]. AfroNet [SM21]. After [Ell22, Ade21, Bea20, Kra21b, See20b, Str22a, Wal21a]. Afterlives [Kno22b]. Aftermath [Kno20]. against [Die21, Rua24]. Agata [Lab21]. Age [AD22, Cla21, Del21, Het23, Hor23, Jac23b, Kan24, LM20, Lin24a, Lin21, Loe24, O’D22, Rob20, Rob22c, Tho22, Tho23a, HP21a, Rie22, AL20, Cor23a, Van20b, Bim22, Gab23, Lin22, Lud23, OA23, Sch23c, Sla23, Ves20, War23, Bee23]. Agency [Hum24, Kaa23, Coo23]. Agent [Kin23]. Agents [Rom22]. Ages [AD22, Cam23]. agitate [Ber24]. Agrarian [Das24]. Agricultural [Dim24, Lip22, Zul24]. Agriculture [Hah23, Mal24, Hay22]. Agroindustrial [Vai23]. Agronomy [Sei24]. agropecuaria [Zul24]. Aguiar [Mat21]. Aguilar [Sol24b]. Agustí [Bar23, Hof21]. Ahmed [Mel23, Zem21]. AI [Fru21, Kag20, Kli23]. Aid [Zou22, Rey22]. AIDS [Gre22, Mon22]. Air [Cru21, Esp23, FR21, Han23, Kob22c, Spi20, Tho23b, Ven24, Cha21a]. Air-Conditioning [Cha21a]. Aircraft [Zai23, Zia23]. Airplane [Als20, Sch20d]. Airspace [Cru21]. Akteure [And22]. al [Lóp22, Lóp24, VÁ21]. al. [Bau23, Bee23, Bit21, Cha22b, Cha21c, Cig23, Die22, Duh22, Fla23, Fre22, Gar20b, Goo21, Gué22a, Hoc23, Kai21, Kos21, Lal21b, Mar23a, Mat21, Pom20, Ram22c, Spi21, Tat22, Tri20, VÁ21]. Alain [Ver22b]. Alan [Abn23, Fru21]. Alarm [HH21]. Albert [Kar23]. Albertini [Cur23]. Albrecht [Bee21]. Albro [NG20]. Aldebert [Fla23]. Alerta [Gil24a]. Alessandra [Mil22]. Alex [Mil23a, Rad23]. Alexander [Dav23a, Ell22, Gab23, Mar21a, NG21, Bea20]. Alexandrine [Yam23a]. Alexey [Vee22]. Alexis [HP21a]. Algeria [Zou22]. Algorithm [Pet24b]. Algorithmic [Bin23]. Algorithms [Hum24]. Alice [Hva24]. alimentaires [Wil21b]. alimentation [Wil21b]. Alison [Jac23b]. Alix [Lal21b]. Allah [Hal21a]. Allais [Sta23a]. Allen [Rik21]. Allende [Rag24]. Alles [Dis20]. Alliance [Lun22]. Allied [See20b]. Allison [Wen22]. Alltag [Aug21]. Allyson [Wån23]. Aloft [Zyl22]. Alonso [Tho24]. Alternative [Tat21, HP21a]. alternatives [HP21a]. Altice [O’D20]. Altitude [She20]. Aluminum [Jør21]. Am [O’D20]. Amahia [Fre20]. Amanda [Pry23]. Amateurs [vM20]. Amber [Che23]. ambiental [Bar23]. Ambivalence [Mch24]. America [Ach22, Bou22, Dae21, Jac21, Kin23, Kru22, Pet24a, PÑ22, SD23, Zdr21, Cas24, Dae24, Mur20, Oso22, PW24, RCCAM24, Spi20, Bur23, Cha21b, Ger21b, GC24, Jan23a, Kat23, Kay20, Lif23, Mei22b, O’D22, Zyl22, Bou21]. American [Atw23, BF24, Bix24, Bla20b, Bro21a, Cha21a, Coh21, Dal22, Ham24, Hay22, How23, Huh24, Lal21a, Lau21, LM20, Mal24, Mue22, Mul21, Nor21, O’S20, Ott21a, Pin21, Pos23, Rob22a, Ros20, Rus24b, Sca20, Sch20d, Sim20, Spi24, Tar23, Vet23, Wel21, Zai23, vdL22, Bro22, Pop20, TT22, Bri20, Don24]. American-Led [Zai23]. Americans [Isr23, Kno22a]. Ames [Löv23]. amid [SD24]. Amir [Cla21]. Amish [Nol23]. Amphitheater [Mur20]. Amplified [Haj21]. Amsterdam [Wac23]. Amusement [Abn23]. Amy [Kin23, Zyl22]. Ana [Di 21, Don23, Zap20]. Analog [Hoc23]. Analogy [Duh22]. Analysis [Eng22]. Anarchy [Ove22]. Anatomy [Plo23]. Ancient [Gué22a, Kal20, Yan22]. Andean [Ram22c]. Anderson [Als20]. Andra [SD23]. André [Nèg22a, Sou24a]. Andrea [Far24]. Andreas [Ham21, Nel22, Sch22]. Andrés [GC24]. Andrew [Ass23, Emm21, Lou21, Mei22a, Sla23, Vol22]. Andy [Dis21a]. Anfängen [Bee21]. Angela [Ree22, Wac23]. Angeles [Cam24]. Angels [Pui22]. Anglo [Mue22]. Anglo-American [Mue22]. Angola [Kui22, Per23b]. Angry [Hes20b]. Animal [Mit22]. Animals [Nov23a, Sad24]. Animating [RW23]. Anker [Koc21]. Ann [Cur23, Gol24, Kos21]. Anna [His23a]. Anne [Cha21b, Hah23, Hva24, Koc23, Nèg22a, Par20b, Pou23a]. Anne-Francoise [Nèg22a]. Anne-Marie [Pou23a]. années [Has22]. anni [How22]. ante [Cam22]. Antebellum [Har21]. Anthony [Lip22]. Anthropocene [Lin21, Kap20]. Antibiotics [Mue22]. Anticoncepción [Lab21]. Antiquarians [Bon23]. Antiquity [Tay23]. Antoine [Mic21, Rav20]. Antonin [Wal22]. Aparecida [Mac22a]. Apartheid [Mch24]. Apocalypse [Mit23]. Apocalyptic [Wel21]. Apollo [Ell22]. Apostle [Ver21]. Appalachia [Str22a]. Appeal [His23b]. Appearance [Pen22]. Apple [Loe24]. Approach [WAB^{+}23]. April [Ano20e, Ano20f, Ano21f, Ano22e, Ano23c, Ano24b]. Arab [El 21]. Aral [Kam22]. arcade [Hai21, Abn23, Abn23]. Archaeologies [Mar21a, Plo20]. Archaeology [Hol21, Loo22, Bar21a]. Archeological [Bea20]. Archipelago [Bin24]. Architects [Hal20, MK23, RR21a]. Architectural [Bea21, Mar21a, Moo21, Spi21, Suá21, Hal20]. Architecture [Cha21a, Cro23, Del21, Fin22, FR21, Gli23, How23, Mei22b, Par20b, Zyl22]. Archives [De 23b, ZK22]. Archivists [Edi22]. Arctic [Bou22]. Arenas [Kai21]. Ares [Dal22]. Argentina [Als22, Mat21, SD24]. Ariel [SG21]. Arm [Bro21a]. Arne [Cig23]. Arnold [Ale23, Bla21]. Arrived [Mac22a]. Art [Bot21, De 23a, Ham21, Hol21, Kal20, McC20, McC22, Tys24, Daz23, Luc20a, Per23a, RR21a]. Arte [McC23]. Artefacts [Low20, De 23b]. Artemis [McN24]. artifices [How22]. Artificial [Bri20, Ves23b]. artificij [How22]. Artillerymen [dCC22]. artilleurs [dCC22]. Artisan [Gun24]. Artisanal [Sto23, Dup21]. Artisans [Hom22, Gui22, RR21a]. artísticas [Zap20]. Artists [Ger21b, Tys24]. Arts [Cue22, Dup21, Mor23, Sch23c]. Asare [Ade21]. Ash [Mar22a]. Asha [Kan23a]. Ashley [War23]. Asia [Dav23b, Han22, Kam22, RO23, Hag21, Fru20, Kal23]. Aspiration [Bee23]. Aspirations [Kai21]. Assembly [Arc23a]. Assessment [Mcg21]. assistance [Gou21]. Assisted [Fre22]. Associations [HF21, Mil24]. Assured [Han23]. Astrid [Uts21]. Astroculture [Ell22]. Astronaut [Ros22a]. Astronauts [Lau21]. Astronomical [Wit21]. Asynchronic [Ban21]. Atacama [GMR24]. Ataíde [Hva24]. Atari [O’D22, Šve21]. Atkinson [Haj21]. Atlantic [Pol21, Sto22, Ber21e, Ash21]. Atmospheric [Tsu23]. Atom [Cig23, KJ23, Kru22, Pol24]. atomarer [Aug21]. Atomgeschäfte. [Kir22].

Atomic [Bin24, Coo23, Hög22, Jac23b, Kru20, Pea22, Aug21, Ade21, Kno22a, OA23]. Atoms [RO23]. Atomville [MK23]. attelé [Gui22]. Atwood [Ask22]. Audible [Yam23a]. Audiences [Kus23a]. Auditory [Kvi23b, Mar21c]. Audrey [Cai22]. Augmented [Aks23]. Augusta [Yu21]. Augustine [Pre21]. aujourd’hui [Mar24]. Aural [Yam23c]. Aurality [Goo21]. Ausgestrahlt [Gut23]. Austerity [Pos23]. Austin [Sch21c]. Australia [Lan23b]. Austria [Uts21]. Autarchia [Cal22]. Autarky [Cal22, Kra21b, Rey22]. Author [Ano20a, Ano21a, Ano22a, Ver23]. Authority [Har20b, Ram22c, Ign21]. Auto [Pos23]. Auto-Centric [Pos23]. Automata [Wac23]. Automated [Kaa23]. Automation [Rus24a, Sch23b, Zac23]. Automobiles [Ess21]. Automobilism [Sie22]. Automobility [Muk22]. Automotive [Mil24]. Autonomy [Kuo23]. Autonorama [Lad22]. Autorité [Pre23a]. autrement [Nèg22a]. autres [Wil21b]. Autumn [Fou23]. aux [Daz23, Gra22]. avant [Cha22b, McC20]. Avant-Garde [McC20]. Aviation [Als22, Bed22, Lal21a, Mel23, Pea22, vdL22, Nis23, Tap23]. AVISTA [Luc20a]. Awards [Ano20b, Ano21b, Ano22b, Ano23a]. Ayuntamiento [Mat24]. Ayurveda [Nai22].

B [Dav23b, HF21, Kri21, Lou21, Pre23a, Rei21b, Str22b, SD23, ZJ22]. B.C. [Ram22c]. Back [Pat24]. Backroom [Zea22]. Badezimmer [Sta22]. Baekeland [Fri22a]. Bakelite [Fri22a]. Bakhtiar [Kar24]. Bala [Jon23]. Balbi [Nel22]. Balkan [Kan24]. Balkans [Don23]. Ballim [Mch24]. Balloons [Cru21]. Baltimore [Chu21]. Bambuk [d’A20]. Bangham [Ber21a]. Banks [War23]. Baptiste [Jon23]. Barak [Brü21]. Barbara [Eva21, Mar23a, Tel20, Lóp24]. Barbenheimer [Old24]. Barber [FR21]. Barbie [Mai24]. Bardoneschi [Gui22]. Barriers [Sca20]. Barron [Daz23, Sol24b]. Base [Bou22]. Basin [Kam22]. Bassett [Pra20]. Basu [Rai23]. bathroom [Sta22]. Baton [MSH20]. Batteries [Hir23]. Battle [Rua24]. Battleship [Pom20]. Baumgärtel [Mac24]. Baxter [Ham20]. Bay [Jon22b]. Be [vdS22, Sch21b]. Beach [Boo22, Cam24]. Beam [PZ23]. Bearers [Gün21]. Beaune [Duh22]. Beautiful [Mon22]. Became [Cie21, Koc21, Kon21, Loe24, Ven24, Lev20]. Beck [McC20]. Beckett [Opp22]. Become [Nov23a]. Becoming [Hor22]. Bedeutung [Jan23b]. Beese [And22]. Before [Bin23, Cha22b, FR21, Nag21, Zyl20, Buc23]. Begets [Kob22c]. beginning [Bee21]. Beginnings [McD22]. behind [Kan24, Wac22, Low20]. Behrisch [Har23a]. Beijing [Zha20, Lav21]. Being [Tri20]. Bek [Poo23]. Belfanti [Ber22b]. Belief [PZ23]. Beliefs [Wel21]. Bell [Smi23]. Belle [Bot21]. Belt [Her20]. Bender [DeC22, Dis21b]. Benefit [Sch20c]. Benevolence [Fru20]. Bengal [Muk21, vdS22]. Benjamin [Den20b, His23a, Kre21, Mcg21, Per23a, Rus20, ZJ22, The22]. Benoît [Bel21]. Benoit [Sch20a]. Bentham [Rei21a]. Berliner [NG21]. Bernadette [Gil24b]. Bernard [Le 23]. Bernasconi [Bau23]. Bertillonnage [EL20].

Bertilorenzi [Cal22]. Bertrand [Mit23]. Bertucci [Dup21]. Berveglieri [How22]. Bess [GB22, Zdr21]. Betrayed [Mac22b]. Between [Di 21, HP21b, Cas21, Gui22, Pou23b]. Bewegung [Eng20]. Beyond [Bim22, Cue22, Lan23a, OY23, Cam24, Gué22a, Fri22a, Kai23, Lau23]. Bhabha [Kar24]. Bharat [Pra23]. Biagioli [Wal22]. Bian [Ben22]. Bicycle [Car21, Poo23, Sme20, Tur20, Wie21]. Big [Ber21c, Mey20]. Bigelow [Wen22]. Bihar [Das24]. Bihari [Nai22, PW24]. Bildfabriken. [Hol22]. Bildschirmtext [Kub22]. Bill [Cha22a]. Billing [VF24]. Billington [Dae21]. Bind [Cat21]. Binder [Pet24b]. Biography [Cho22]. Biological [Wil23]. Biology [Ber21c]. Biopolitics [Ada20]. Biowarfare [Lip22]. Birdsong [Mun23]. Birth [Bou21, Ger21a, HG24, Mul21]. Birthing [San20]. Births [San20, Che23]. bis [Bee21, Kub22]. Bishop [McC20]. Bismarck [Pom20]. Bits [Man20]. Bittel [Mak23]. Black [Fou23, Jon22b, Jon23, Moo20, Mor24, SM21, Fer23b, Jon23, SM21]. Blair [Kos21]. Blake [Ask22]. Blanc [Ald20, Buc23]. Blanchard [Lif23]. Blary [Pou23a]. Blau [NG21]. Bleached [Pag23]. Blond [Cha21c]. Blood [Ber21a]. Bloom [Pos23]. blue [NG21]. Boats [Bar21a, Jam22]. Bob [Kar23]. Bodies [Ash21, Bun23, Pat24, Yu21]. Body [Ber21b, Bix24, Hor22, Val21]. Bolton [Bou23]. Bomb [MK23]. Bombing [Bla20b]. Bonan [Ber24]. Book [OY23, Vir22]. Books [Cas24, Lin24a, Cha22b]. Booß [Aum21]. Borderland [Şer24]. Borderlands [Mun20]. Borders [Tho24]. Bored [Hes20b]. Borlée [Gab22]. Bösch [Sch20b]. Bot [Ver22b]. Botanical [dZ21]. Bottomley [Mei22a]. Boudreault [Yam23a]. Boudreault-Fournier [Yam23a]. Boulton [Fer23a]. Bound [Cur23]. Bousquet [Rav20]. Bovine [Arn23b, Kan23b, Nov23a, WJ23]. Box [HH21]. Boyer [Kap20]. Braided [Nai22]. Braille [CW24]. Brain [Rad23]. Brandau [Ell22]. Branson [Isr23]. Brass [Hom22]. Braunschweig [Bee21, Bee21]. Bray [Ale23]. Brazelton [Yu21]. Brazil [Cru21, Buc23]. Bread [Pag23]. Break [Che23]. Breast [Fra21]. Breeding [Cha21b]. Brescia [Per21b]. brevets [Pou24]. Breyfogle [Smi20]. Brian [Aru23, Coo20, Lud23, Wri21]. Brice [Kou24, dCC22]. Brick [Don22]. Bricolage [Pat24]. Bridging [Del21]. Bright [HG24, Kan20]. Brigitte [Rie22]. Brilliant [San21]. Brings [Dis21b]. Britain [Cam21a, Kno22b, Mcg20, Tuc22, Zai20, Bon21b, Jam22, McG24, Pag23, Røs22b, TS20]. Britannia [Abn23]. British [Abn23, Cox22, Cra21, Das24, Giu22, Han22, Plo23, Pol21, Rei21a, Sch20d, Smi23, Tho23a, Zai23, Eam21, Ram23, War20, Cop24, Dan21]. Broadcasters [Jon23]. Broadcasting [Tho24]. Broadside [Plo23]. Brockmann [Ach22]. Broeckmann [Ham21]. Bronze [Rob22c]. Brooklyn [BMSS24]. Brothers [Tel20]. Brown [DiM22, PW24]. Brownell [Deg22]. Bruce [Giu22, HF21, Sim20]. Bruno [Ber21c]. Bruyerre [Gra22]. Bruyninckx [Mun23]. Bryce [Lal21a]. Builders [Zel22]. Building [Bec23, Cal21, Cha22a, Cha22c, Cli20, Har24, Les24, Mar22a, Sha22a, CW24, How23]. Built [Ber21d, Hva24, Par20b, Ram22b, Sch23b]. Bulb [TT22]. Bulbs [Vik20].

Bulgaria [Bek24]. Bulgarian [Kan24]. Bull [Lun23]. Bullet [Low23]. Bundesrepublik [Gut23, Kir22, Kra23, Kub22, Sch20b, See22]. Burbano [GC24]. burden [Cam22]. Bureaucracy [HP21b, Sei24]. Bureaucrats [Hom22]. Burger [Ram22c]. Burma [Han22]. Burnham [McC22]. Burning [Bla20a]. Burrichter [Rie22]. Bus [Poo23]. Business [Arc23b, Dav23b, Fri22a, Kag20, Mai23, Wie21, Muk21]. businessmen [Mat24]. Buthmann [Mac22b]. Butsch [Arc20]. Butterworth [Wor24]. Button [Par20a]. Buy [His23b].

C [Boo22, Cie21, Ell22, Hof23, Sim20, Syl22, Ver21, Vet23, Wri21, KdS24, Mac22a, Zai20]. ca [Hom22]. Cabinet [Wir23]. Cable [Giu22]. Cain [But23]. Cairo [Ism23]. Caitlin [Sta20]. Calculating [Mcg21]. calentamiento [VÁ21]. California [KM23]. Callapez [Mei23]. cambio [Cas21]. came [Sch21b]. Camelot [Las23]. Camera [Arn23b, Per23b]. Camp [Str22a, Bou22]. Campaign [Røs22b]. Campaigns [Rog23]. Campbell [Arc23b, HF21]. Campbell-Kelly [Arc23b]. Camprubí [VÁ21, Mit23]. Can [Fru21, Hol21, McG24]. Canada [Ang23, Gag21, Sch23c]. Canadelli [Low20]. Canadian [Fer24]. Canal [Far24, Has22]. Canale [Far24]. Cancer [Fra21]. Cannady [Sto23]. cannon [AD22]. Canon [AD22]. Capabilities [Cha22c, HP20]. Capacity [AG24]. Capital [Bee23, Li22, Mun20, Eng20, Sou24b, Kay23, Sou24b]. Capitalism [Chu21, Jan23a, Lou21, Mai23, Mai24, MO22, Rus22, Zha21, Ram22b]. Caplan [McC23]. Captivating [The22]. Cara [Ger21a]. Carbonization [Brü21]. Carcasses [Den23]. Carceral [The22]. Care [KB22a]. Caring [Kan23b]. Carioca [Mac22a]. Carl [Hai21, Jør21, Mod21]. Carla [Mak23, San21]. Carlo [Ber22b, Wie21]. Carlos [Pik21]. Carmen [Di 21]. Carol [Hes20a]. Carolyn [Lin21, Tho21]. Carroll [How23]. Cars [Nor21]. Carter [Cox22]. Carvais [Nèg22a]. Carvalho [Mei23]. Cases [Cal22]. Casey [Hur20]. Casi [Cal22]. Casini [Pat24]. Casper [Har23b]. Cassette [Ass23]. Castillo [Boo24]. Castro [Dim23, Ver21]. Catalan [GL23]. Cateforis [De 23a]. Cater [Hur20]. Catherine [Ves23b]. Caton [Bit21]. Cattle [Arn23b]. Cavalry [Sch20d]. Cave [Kli23]. Caverns [Lal21b]. Caves [Lal21b]. cdrom2 [Hai21]. Ceaseless [Fer24]. Cecilia [Lóp24]. celik [Mar21a]. Cell [Ber21d]. cellars [Lal21b]. celliers [Lal21b]. Cellular [Arc23b]. Center [Hög22]. Central [Ach22, Cos20, Den20a, Kal23, Kam22, PÑ22, Sar22]. Centric [Pos23]. centuries [Bel21, Boo24, Gab22, Gui22, Mic21, Nèg22a, Ver22b, Wil21b, Di 21, Ryd20]. Century [Bae23, Bas21, Bec23, Bou22, Bru23, Cha21b, Chu21, Dae24, Gol24, Hal21b, Hof21, Ing22, Lal21a, Lav21, Lin24a, Low20, Pos23, Rei21b, Rik21, Tuc22, Wac23, Cha21c, Gou21, Huh24, RR21b, Bon23, FV20, Sch21a, Sch23c, TS20, WAB^{+}23, d’A20, Als20, Cub23, Dae21, Dav20, DeC22, Ham21, Kai23, Kri24, Par20b, Pik21, Ren22, Sta23a, Wal21b, Wij22, Bix24]. Ceramics [SC20a, Gil20]. Céramique [SC20a]. Ceruzzi [Heß22]. Cesarean [Got20, Rob22a]. Cesty [Šte22b]. Chachereau [Pou24]. Chaim [Dav20]. Challenge [HM20, HF21, Hun20, Cam22]. Challenges [Van20b, AL20]. Challenging [HHK^{+}20]. Chambers [Fin22]. Chandler [Sch21c]. Chandrashekar [ZP24]. Chanel [AD22]. Chang [Cli20]. Change [Cha20, Lev20, Lin21, Luc20a, Per21b, Rom22, San20, Zan22, Zyl22, Cas21]. Changed [Kag20, Tho21, Tho22]. Changing [Coo20, Hes20b, TS20]. Characteristics [CW24]. Charged [Hir23]. Charisma [Löv23]. Charles [Lif23, Tho23b, Zel22]. Charlie [Dan21]. Charlton [SM21]. Charney [Han22]. Chasing [Zac23]. Chastain [SD23]. Chat [Nag21]. Chaussées [Cha22c]. che [San21]. Chemical [Tho21, Tho22, Vai23, Cal22]. Chemicalization [Hay22]. Chemicals [Kin23, Rom22]. Chemistry [Che22, HL22, Hof21]. Chemists [Tho22]. Chen [Dua23, Naq22, Hom22]. Chernobyl [DiM22, Kno22b, Edi20b, Rin20, Sch20c]. Chevaux [Gui22]. Chicago [MF20]. Child [Löv23]. Childbirth [Shi20]. Chile [Boo24, Lóp24, HG24, Lóp24, Mar23b]. Chileans [Rag24]. chimico [Cal22]. China [Bec23, Cie23, Gie22, Gro21, Hom22, Lou21, Lue21, Rob22c, Sco23, Ste22a, Wu23, Zak23, Zan22, Zen22, Zha21, Don22, GX24, Got20, HG22, Kif24b, Li24, San20, Sch21a, SC20a, Yi22, Zou22, Ben22, Rei21b, Tan21a, Yao21, Yu21, ZP24]. Chinatown [Pen22]. Chinese [CW24, Cli20, Dua23, Pen22, Sha22b, Yan22, Kif24a]. Chinese-Funded [Kif24a]. Chloë [Emm21, Gué22a]. Choi [Sch21c]. Choi-Fitzpatrick [Sch21c]. Cholera [Cla21]. Chornobyl [Wen20]. Choubey [Kan23a]. Chris [Ber22a]. Christian [Aum21, Kai23]. Christie [Cra21]. Christina [Cam24, Gar24, Gli23]. Christine [Ger21b, Mak23, Wag23]. Christopher [Kre21, PW24, Zai20]. Christos [Bol21]. Chronicles [Lie23]. ch’un [Bec23]. Chuño [Oso22]. Ciência [Sou24b]. ciencia [Dal22, Mat21, Zul24, Cam22]. Científica [Sou24b]. científico [VÁ21]. Cine [Bae23]. Cine-Mobility [Bae23]. Cinema [Bro22, Coa22, Don23, Gar20b, Neg22b, Pen22, Cra21, Smi23, Tel20]. Circuit [Hoc23]. Circulating [Ryd20]. Circulation [Dav23b, Gun24, Ign21, JR23]. Circumpolar [Gil20]. cité [Fon23b]. Cities [Dav20, Muk22, Pik21]. Citizen [Loz22, Yam23c, Yu21]. Citizens [Hei23, Kin23, Kno22a]. Citizenship [Zyl20]. Citroën [Den20a]. City [Chu21, Edw23, Fre20, Hal21b, Mat24, NG20, War23, Wen24, dWB22, Boo24, Gal22, Ste22a]. Ciudad [Mat24, Boo24]. Civic [Les24]. Civil [Cha22a, Har23b, PW24, Wri21, Yam23c, Zyl20, vdL22, Bed22]. Civilians [Tho23a]. Civilization [Ale23, OY23, Ram22c]. Civilizing [Den20a]. Claas [Mue22]. Claimed [vM20]. Clair [Lin24a]. Claire [Gul20, Gul20]. Claiton [Het23]. Clark [Gid24]. Clarke [Fin22]. Class [Als22, Cie21, ZK22]. Classes [Sar22]. Classics [Zem21]. Claude [Fon23b]. Claude-Pierre [Fon23b]. Claudio [Het23]. Clayson [Bot21]. Clean [Hir23, Tho23b]. Clément [Lal21b]. Climate [FR21, Gär20a, Lin21]. Climatic [Tsu23]. Clock [Syl22]. Cloth [Ram23]. Cloud [Mai24]. Clouds [Sto24]. Coal [Brü21]. Coals [Kob22a]. Coats [PW24]. Cobbold [Tho21]. Code [Wal22, Le 23].

Codebreaking [HP20]. Codes [Arc23a]. Codification [HP21b]. Coefficient [Bea21]. coffee [Bee21]. Cohen [His23a]. Coin [Hom22, Šve21]. Coin-Operated [Šve21]. Cold [Aug21, Bab22, Bed22, Bin24, Bro21b, Ell22, Gär20a, Ger21b, Gil24b, Hai23, Han20, Kan24, Kui22, Lau21, Mit24, Pou23b, RO23, Rey22, Rib23, SD23, Tys24, VÁ21, Wil22a, Zai20, Abe21, Har23b]. Cold-War [Rey22]. colectivo [Boo24]. Colin [Ham20, Van20a]. Collaboration [De 23a, Rof20, Vol22]. collapse [Aum21]. Collecting [ZK22, Ber21c]. Colleen [Ang23]. Collision [Sla22]. colloque [Die21]. Colonality [Fra21]. Colonial [Arn23b, Arn23a, Bee23, Das24, Don22, Gre21, Hag21, Hva24, Ism23, Lan24, MO22, Rei21b, Roy21, Wen22, vdS22, And22, Kan23a, Har20b]. Colonialism [Sim24, Amm22]. Colony [Sei24]. Color [Car21, Che22, Kan20, Mur20, Ves23b]. Colorado [Mal24]. colorants [Wil21b]. Colored [Zyl20]. Colossus [HP20]. Com [Erd21]. Come [Sch20a]. Coming [Chu23]. Commercial [Mel23, Pop20]. Committee [BBC^{+}21, ŁPO^{+}20]. Commodities [Pre23b]. Communicating [Sch22]. Communication [FV20]. Communications [Gum24, Joh23, Kir23, Lau23]. Communist [Ďur23, Zha21, vM20, GX24, Mit22]. Communities [Les24, McG24]. Community [Dae24, Lau21, Nol23, dWB22, Lin24b]. companies [Cal22, Jud22, Ver22b]. Companion [Kos21]. Company [Muk21]. Comparative [Aeb23]. Compendium [Mar23a]. Competition [Gum24]. Complex [Lev21, Zan22, Mil23a]. Complicity [OA23]. Computational [CC21, WAB^{+}23]. Computer [Cer20, Loe24, McC23, Sch20b, Sch21b, SC20b, Sla23, Tat22, War20, War23, vM20, San21, O’D20]. Computereinsatz [Kra23]. Computerisierung [Kra22]. computerization [Kra22]. Computernützung [Sch20b]. Computers [Bro22, Gil24b, Plo20]. Computing [Bro22, Cor23b, Fer23b, Kan24, Sch23a, Cor23b, Heß22]. Con [Bur23, Zul24]. Concept [Wil21a, ZJ22, Nye20b]. Concise [Mas20]. Concrete [Bee23]. Conditioning [FR21, Cha21a]. Confessions [Pre23a]. Confinement [KM23]. Conflict [Fla23]. Conflicted [Tar23]. Cong [Dua23]. Congo [Amm22]. Conjuring [Gid24]. Connected [Ber21d]. Connecting [Bon21a]. Connections [Pre23b]. Connective [Mil22]. Conquering [Tay23]. Conscience [OA23]. Consciousness [Nye21]. Consequence [Rob22a]. Consequences [Ung22]. Conservation [GF21, Kob22c]. Conservatoire [Cue22]. Consortium [Yoo20]. construcción [Bar23]. Constructing [GX24]. Construction [Fer24, Hus23, Huh24]. Consulting [Pui22]. Consumer [Hal21a, Ism23, Mil24]. Consumerism [Fra20, Zha21]. Consuming [Nor23]. consumption [Rei21c, Bla20a]. Contact [Chu22a]. Contagion [Cla21, Bon21b, Ada20]. contaminants [Wil21b]. Contamination [Lin22]. Contemporânea [Sou24b]. contemporary [Sou24b]. Contest [Zia23]. Contestations [Pag23]. Contested [Hag21, Wan21, Wu23]. Contextualizing [HP20]. Continuity [Pol21]. Contraception [Lab21, Mas20]. Contraceptive [Ign22]. Control [Aks23, Gär20a, Gum24, Hoc23, McC23, Boh21, Cur23, Tsu23]. Controllers [Sol24a]. Controlling [Kus23a]. Conventions [Wor24]. Convergence [Mei22a]. Conversations [Ber23, She20]. Cook [Twa20b, Twa20b]. Cooperated [Has21]. Cooperation [Fla23, Şer24]. Coping [San20]. Copper [Ber22a]. Córdoba [Whi23]. core [Hai21]. Coreen [Wil22b]. Cork [Cri22]. Corn [Her20]. Cornel [Ten22]. Corona [dWB22]. Corporate [Chu21, Don24, Les24, O’S20]. Corporation [Mar21b, Ram22b]. Corporations [Wij22]. Corporeality [Gar20b]. Corruption [Eam21]. Corry [Pel24]. Cortada [Hei20]. Cosmetics [Gro21]. Cosmonauts [Hil23a]. Cosmos [Hil23a]. Cossart [dCC22]. Costa [Hum24]. Cotton [Ram23]. Could [Les20]. Counterside [Pat23]. Countries [Jud22]. Course [Sla22]. Courts [Mea21]. Cover [Hil23a, Kif24b, Kus23b, Nov23b, Sol24a, Yam23b]. COVID [Gro22]. Cows [Kaa23]. Craft [Bra20, Che22, Hil23a, Hub24, Lan23b, Mon23]. Crafted [Yi22]. Crafting [Kan23b, Sto23]. Crafts [MO22, Kan23a]. Craftsmen [Has21, Kan23a, Sch21a, Gui22]. Craig [Bla20b, Jon22a, Wir23]. Crawford [Gli23]. Creager [Ree22]. Created [Hal20]. Creating [Bau20, Dae24, Esp23, dWB22]. creation [Sch21b]. Creative [Tys24]. Creativity [Lan23b]. Creveld [vD21]. Crime [Erd21]. Criminalization [Lud23]. Crises [KMÅ20]. Crisis [Hah23, Wal21a, Wil21a, Cro23, Her23]. ’critical [Ver22a, Mod21, Nye20b, Pie22]. Cronqvist [Har23b]. Crook [Tom21]. crop [Soe21]. Cropper [Pel24]. Cropscapes [Bra20]. Crossborder [Fla23]. Crossing [Sla22]. Crouy [AD22]. Crouy-Chanel [AD22]. Crow [BF24, Car21]. Crowdfunding [Vet23]. Crowds [Kus23a]. Crowdsourced [ZK22]. Cruz [Mai23, Mei23]. Cryptocurrency [Vet23]. Cryptography [FV20]. Csikszentmihalyi [Mor23]. Cuba [Loz22]. Cuchí [Mat24]. cucina [Sca21]. cuirassés [Daz23]. Cuisine [Gil20]. Cultivating [Mon22, GF21]. Cultural [Ing22, Low23, Mei22a, Mit23, Pel24, RO23, Rus24b, Ber22b, Sca21]. Culturale [Ber22b, Sca21]. Culture [Arc20, Ass23, Ben22, Bin23, Bla23, Don23, Eri21, Hal21b, Hum24, Map23, Nor20, Par20b, Pen22, Sca23, Sub22, Zem21, Rei21c, Jac23b, Kan23a, Pol24, Sto22, Tys24]. Cultures [Bau23, El 21, Per23a, Ryd20, Gar24, Wis22]. Cummings [Rad23]. Curation [Cor23a]. Curators [Edi22]. Cure [Pra23]. Current [Cha22a, Bac22]. Curry [Cha21b, Hah23]. Curtain [Bro21b, vM20, Kan24]. Cyber [War23]. Cyberia [Kan24]. Cybernetic [Rag24, Aum21]. cybernétique [Tru24]. Cycling [Car21, Fri22b, Mor24, Cox22, Fri22b]. Cymene [Kap20]. Cyrus [Bro21a, Hof23]. Czechoslovak [Kvi23b]. Czechoslovakia [Ďur23, vM20].

D [Als20, Arc23b, Arn23a, Bim22, Del21, Gär20a, Mel23, Ott21b, PW24, Le 23, SM21, Zyl22, Ban21]. Dadabhoy [Kar24]. Dagen [Pos23]. Daggett [Ger21a]. Dahlquist [Gar20b]. Dalton [Ber21b]. Dam [Buc23, Hus23, Huh24]. Dams [RCCAM24]. Dana [Wel20]. Daniel [Arc23b, Ell22, Fal24, FR21, Hei23, Jac23a, Kay20, Kir22]. Daniels [Kat23]. dans [Bau23, Duh22, Fon23b, Lal21b]. Danube [Şer24]. Dar [Deg22]. Dark [Bri20]. Darwin [Kru22]. d’assistance [Gou21]. Data [Ber21c, Fou23, Pat24, Sch23b, Sim24]. Daughton [Amm22]. David [Ald20, Bee23, Ber21b, Bim22, Chu21, Chu23, Cri22, Dae21, Dav23b, De 23a, Fra20, Gär20a, Gun24, Han23, Mal24, MF20, Mil20a, Pin21, Plo20, Sch21b, Sör23, Tar23]. Davide [Bin24]. D’Avignon [Duq24]. Dawn [OA23, Ott21a]. Day [Røs22b, Bea21]. DDR [Kra23, Mac22b]. deals [Kir22]. Death [Löv23]. D’eau [Bel21]. d’eaux [RR21b]. Debate [Ful24, Map23, Par20b, SV21]. Debating [Joh23]. Debatte [Gut23]. début [Gou21]. débuts [Pou24]. Decisions [Sub22]. Decline [Ďur23]. Decomposed [Yam23a]. deCourcy [Bon21a]. Deep [Cho22, Ful24]. Defence [Har23b]. Defense [Eam21, Yam23c]. Deficiency [KM23]. Defoliation [Kin23]. Degani [Edw23]. Degraded [Rus24a]. Dekker [Fri22b]. del [Ber24, Ber22b, Cal22, Mat24, Lóp22]. Delayed [Mar21c, Kal20]. Delgado [Cam22]. del’informatica [San21]. della [Ber24, How22]. dell’innovazione [How22]. Delta [Swa21, Mil23a]. Delusion [Hen21]. DeMaagd [Wån23]. demain [Mar24]. Demands [Sch21a]. Demidowicz [Fer23a]. Democracy [AL20, Bou21, Bre23, Mil24, Zai20]. Democratize [Sch21c]. Denied [Bla21]. Denise [Gab22]. Denmark [Uts21]. Department [Wan21]. Depletion [Lev21]. Deploying [Les20]. Derbyshire [Arn23a]. Derek [Her20]. Derksen [vE22]. Derouet [Mic21]. desarrollismo [Lóp22]. Desde [Dal22]. Desert [GMR24]. Design [FR21, GB22, Haj21, Jør21, Pie22, Pom20, Rag24, Šve21, Bix24, Bla23, Bru23, Cha22a, LM20, Zdr21, Mar21a]. Designers [Als20, Hal20]. Designing [KB22a, Kob22a, Les24, Abe23, Che23]. Designs [Fer23b, Sta23a]. désindustrialisation [Fon23a]. Destruction [Sta23a]. Determinism [Ves23a]. Dethlefs [Aeb23]. deutschen [And22]. Deutschland [Kir22, Kra22, Kub22]. Deutschlands [Soe21]. Development [Bec23, Bur23, Dah23, Das24, Har21, Kir23, Rei21b, See20a, Tri20, Wol23, And22, Aud22, Fil23, PÑ22, Yan22]. developmentalism [Lóp22]. Devine [Yam23a]. Dhan [Pet24a]. d’Hélène [Nèg22a]. Diagnostic [Fra21]. dial’nice [Šte22b]. Diana [Sca20, Wen24]. Diane [Coa22]. Diasporic [Lue21]. Diaz [Bur23]. Dick [Cor23b, Sch23a, Zel22]. dictatorship [Mat21]. Did [Mil24]. Didier [Bou21]. Diego [Mat21]. Dietrich [Ram22a]. Difference [Wil22b]. Different [Sme20, GC24]. Difficult [DiM21]. Diffusion [Tan21b, VF24, Gab22]. Digital [Gab23, GM22, Hal20, Lin24a, Loz22, Lud23, Neg22b, Rog23, Ves20, Zea22, BBC^{+}21, Kra22, Sch20b, Sch21b, Hoc23, Bit21]. Digitale [Sch20b]. digitaler [Sch21b]. Digitalisierung [Kra23]. Digitality [McD22]. Digitally [RW23]. Digitize [Lud23]. Dihal [Kli23]. Dillon [Kli23]. Dimensions [Sin23]. DiMoia [Fru20]. d’ingénieurs [Cha21c]. d’invention [Pou24]. Diplomacy [Abe21, AL20, Lun22, Nel22]. Dirk [Dio24, Dis20]. Disability [Dae24, GB22, Wil22b, Zdr21]. Disabled [Dae24, McG24]. Disappearing [Kru20]. Disaster [HG24, Kno22b, Plo23, Pos23, Wen20]. Disasters [Kno20]. Discontinuity [OY23]. Discordant [Jac23b]. Discourses [Tuc22].

Discovery [Har23a]. Discrimination [Mcg21]. Disease [Fra21, GA21, Tho22]. Diseases [Bon21b]. Disenchantment [Lan23a]. Display [Aks23]. Displaying [Low20, SG21]. Disposal [Bor23]. Disrupted [Fer24]. Dissemination [Rai22]. Dissensuous [Wån23]. Distant [Cam21a]. Distributions [Fra21]. Diversity [Gar24]. divided [Kra22]. Dixie [Hur20]. Do [MS21, Wil23, Sch20a]. Dockyards [Rei21a]. Doctor [Sta23b]. Doctoring [Nai22]. Doctors [How23, OA23]. Doctrine [MS21]. documenta [Mac24]. Documentation [De 23b]. Documenting [Col20, Kno20]. Documents [Tho20]. Dolores [Pre21]. domestic [Sca21]. domestico [Sca21]. Dominic [Kap20]. Donna [Mas20, San21]. Don’t [Koe23]. Donzé [Jud22]. Doomsday [Syl22]. Dóra [Bro21b]. Dorin [Fru21]. Doron [Rog21]. Dorondel [Dah23]. Dorothea [Rie22]. Dorothy [Naq22]. Dot [Erd21]. Doucet [Gab22]. Douet [Cro23]. Douglas [See20b, Ung22]. Dourish [Man20]. Dove [Ade21]. Down [Sca20, Wil20]. Draining [Mar22a]. DRAM [Yoo20]. dramatic [Cas21]. dramático [Cas21]. Dream [Low23]. Dreaming [HG24]. Dreams [Den20b, Kam22, Kou24]. Drilling [Zha20]. Drinking [Wei20]. Driscoll [Mil23b]. Drive [Nor23, Yoo20]. Drivers [Har20b]. Driving [Lad22]. Drone [Sch21c, Rav20]. Drucker [Mas20]. Duarte [Bia22, Di 21, Wis22]. Duckworth [Emm21, Gué22a]. Dunbar [Cam24, Gar24]. Dunbar-Hester [Cam24, Gar24]. duo [Hai21]. Dupree [Cur23]. durante [Cal22]. Durco [Šte22b]. durée [Bau23]. Durham [Sto24]. During [Het23, Lau21, Cal22, Cub23, Har24, Ram22c, Šte22b]. Dustin [Rus24b]. Dusty [Gue22b]. Dutch [Fri22b]. Dutfield [Pie22]. Duval [De 23a]. Dwelling [Cie23]. Dyes [Che22, Tho21]. Dynamism [Shi20]. Dystopia [Ell22].

Early [Ben22, Dae24, Dav23a, Fru21, Gar20b, Gli23, Har24, JR23, Koc23, Luc20b, Mar22a, Ram22c, Rei21c, Ros22a, Tan21b, Tap23, Tel20, Wen22, Wij22, Wor24, Gou21, NG21, Car21, HP21b, Ima22, Don23]. Ears [Yam23c]. Earth [CC21, Lin22, Tho22, Vol22]. Easing [Mul21]. East [Brü21, Dav23b, Mit22, Muk21, Pou23b, Hag21]. Eastern [Cha20, Dah23, Kan23a]. Eat [Kag20]. Eaton [Gre21]. Eberspächer [See22]. échelle [Cha22b]. Echo [Fin22]. Eclectic [RCCAM24]. Ecobiography [Bar22]. Ecohistory [Gro23]. École [Cha22c]. Ecological [Dah23]. Ecologics [Kap20]. Ecology [Smi20, Yam23a]. Economic [Arc23b, Arn23a, Gat20, Lag24, Hay22]. Economics [Tho23b, Wis22]. Economist [Ism23]. Economizing [Zou22]. Economy [Muk21, Emm21, Rad23, Str22b]. Ecuador [Gid24]. ed [Aeb23, Arc23a, Bar23, Bas21, Bau23, Bea20, Bee23, Bia22, Bit21, Bla23, Bon21a, Boo24, Bor23, Cal22, Cam21b, Cam22, Cha22a, Cha22b, Cha21c, Cig23, Cli20, Col20, Cop24, Cor23b, Cur23, Dah23, Dal22, Dav23a, Dav23b, DeC22, De 23a, Di 21, Die21, Die22, Dim23, Duh22, Ell22, Emm21, Fla23, Fon23b, Fox21, Gab22, Gag21, Gar20b, Ger21c, Gil20, GB22, Goo21, Gou21, Gué22a, Gun24, Hår20a, Har23b, HP21a, Het23, His23a, Hva24, Jan23a, Jan23b, Kai21, Kai23, Kal23, Kli23, Kos21, Lal21b, Lie23, Lóp22, Lóp24, Low20, Loz22, Luc20a, Mak23, Mar21a, Mar24, Mar23a, Mat21, McN24, Mea21, Mei23, Mic21, Mor23, Muk24, Naq22, Nèg22a, Nel22, Opp22, Par20b, Per23a, PZ23, Pet24a, Pin21, Pou23a, Pry23, Ram22c, Ree22, Rei21c]. ed [Rie22, Rik21, Rom22, Rup24, Ryd20, Sch23a, Sch20b, SS21, Sin23, Sme20, Smi20, Sol24b, Sou24b, Sto23, SD23, Syl22, Tat22, The22, Uts21, VÁ21, Ver22b, Ves20, Wal22, Whi23, Wis22, Wri21, Yam23a, Yan22]. Edge [Swa21]. Edges [AG24]. Editorial [Old20]. Editors [Zem21]. Edmond [Daz23]. eds [Fru20]. Education [Mur20]. Edward [vdL21]. Effective [Kif24a]. Effects [Bel23]. Efficiency [Kob22c, Pui22]. Eglė [Bab22]. Egypt [Ass23]. Ehardt [Wag23]. Eigensinnige [Aud22]. eighteenth [RR21b, TS20]. Eighteenth-Century [TS20]. Eike [Kai23]. Eike-Christian [Kai23]. Einat [Opp22]. Elaine [Mak23]. Elder [Syl22]. Electric [Edw23, Haj21, Hur20, Vik20, TT22, Moo21]. Electrical [Giu22, Spi20, Li22]. Electricity [Aug21, Bou23, Coh21, Kay20, Kri24, Mch24, Plo20, Rog21, vdS22, Boy21]. Electrification [Spi24]. Electrified [Yan20]. Electrifying [Wen24]. Electroacoustics [Hor23]. Electron [Wil23]. Electronic [KB22b]. Electronics [Hen21, Rua24]. Elemental [Sto24]. Elena [Low20]. Elevated [Zyl22]. Elisa [Bon21a]. Elisabeth [Coo23, Jan23b, Rom22, Tan21a]. Elision [DiM21]. Elizabeth [Cie23, Gil24a, GB22, Zyl20]. Elvira [Mei23]. Emanuel [Die22]. Emballages [Wil21b]. Embargo [Wal21a, Str22a]. Embodied [Ber21e, Esp23, Gol24]. Embodiments [Sch23a]. Emergence [Mun20, O’D22, Rog21, Sie22, SC20b, Wil21a]. emergency [Aug21]. Emerging [Muk22]. Emily [Deg22]. Emma [Ste20]. Emmanuel [AD22, Bou21]. Emoji [Coa24]. Emotion [Coa22, KMÅ20]. Empire [Ale22, Ash21, Brü21, Col21, Den20a, Gro21, Kam22, Li22, Mit24, Muk21, Wal21b, Gün21, Per23b, Sei24, Giu22, Naq22, Col21, Sha22b]. Empowered [Mac20]. empresarios [Mat24]. Empty [Fut22]. Ems [Nol23]. Enchantment [Gus24]. Encounters [Gid24]. Encyclopaedia [Mei21]. End [Rus24a]. Endangered [Hah23]. Enduring [Gre22, Zan22]. energía [Lóp24]. énergie [Bel21]. energies [HP21a, HP21a]. Energiewende [Sta22]. Energopolitics [HG24, Kap20]. Energy [Bou23, Cig23, Coo23, Gag21, Ger21a, HM20, Hir23, Hög22, Hur20, Jac23a, Kay20, Kob22a, Kob22c, Pea22, Sta22, Str22a, Wel20, Wen24, Lóp24, Rup24]. Engaging [Cig23]. Engelmann [Bol21, Gre22]. Engine [Mil20b, Hai21]. Engineer [Ism23, Wal21b]. Engineering [Aka22, Bos20, Cas24, Cha22a, Dae21, Dim24, GX24, Gil24a, Gol24, Hof23, Lie23, Mac20, Mod21, SC20a, Ver21, Wei20, Wis22, vE22, vdS22, Lóp22, Fru20, Gär20a, Hal21b, Jac21, Jon22a, Şer24, Sha22a]. Engineers [Bar21b, Bon23, Dim23, Jac23a, Lau21, Mon23, Mun20, Tys24, Ver22b, Cha21c, Mic21]. Engines [GC24]. England [Tom21, Ven24, Mar22a]. English [Muk21, Wor24]. Enhancement [Ban21, Kes21]. Enlightenment [Ach22, Dup21, Sto23]. Enough [Ion21]. Entangled [Ger21c].

Entanglement [AL20]. Enterprise [Han23, Muk21]. Entertainment [Kus23a]. Enthusiasm [Ima22]. entre [Cas21, Gui22]. Entrepreneurial [Hei23]. Entrepreneurship [vdS22]. Entreprises [Ver22b]. Entstehung [Sch21b]. Entwicklung [Aud22, Aum21]. Environment [Deg22, Gär20a, Gat20, HHK^{+}20, Hva24, Jac23a, Muk24, Sei24, SD23, Wen24, vdL21, Ree22]. Environmental [Bla21, CC21, Fre20, Her23, Het23, Koc21, Lun22, Mil24, Ram21a, Sör23, Str22a, Tho23b]. Environments [Smi20]. éoliennes [Gra22]. EPA [Tho23b]. Epidemic [Bro21b, Gre22]. Epistemic [BHJI^{+}22]. Époque [Bot21]. Equivalence [EL20]. Era [Don24, Lin24a, Rei21c, Tur20, ZP24, Neg22b, Wel21]. Eric [Don24, Mar22a, Nye20b, Bar21a]. Erika [Har23a]. Erisman [Tap23]. Ernst [Rom22]. Eroine [San21]. Error [O’D20]. Ersatzstoffe [Jan23b]. Erschließung [And22]. Escalator [Hor22]. Escalator-Legged [Hor22]. Espada [Mat24]. Espagne [RR21b]. España [Lab21, Lóp22, VÁ21]. Essay [Hil23a, Kus23b, Man20, Sol24a, Yam23b]. Essays [Bot21, Mod21]. Establishing [Chu22a]. Estados [VÁ21]. estatales [Zul24]. estere [Cal22]. Estudios [Boo24]. État [dCC22]. Eternal [Hal21b]. Eugene [Coh23]. Eugenia [Gro21]. Eurasia [Ger21c]. Eurasian [Smi20]. Europa [And22]. Europe [And22, Cha21c, Cig23, Fla23, Har23b, Has22, Lal21b, PZ23, Sch22, Uts21, Whi23, HHK^{+}20, HP21b, Luc20b, Bou23, Dah23, Dav23a, Gun24, Mea21]. European [Lag24, AL20, Mil24, MS21, Kou24]. Evangelists [Ren22]. Evans [Ber22a, Lal21a]. Events [Sch22]. Every [TT22, Hei23]. Everyday [Lav21, Mor24, The22, Aug21]. Everything [Dis20]. Evolution [Als20, Neg22b, Nye21, Cha21b]. Evolve [Fru21]. Ewa [Ves20]. Ewan [Mar24]. Exchange [Kif24a]. Exhibit [Cue22, Low20]. Exhibits [Cor23a]. Existential [Wel21]. expansion [Mic21]. Expedition [Den20a]. Experience [Mar22b, Mcg20]. Experiences [Ert21, KMÅ20]. Experiencing [Bon21a]. Experiment [Fou23, Ste22a]. Experimentation [Lan24]. Experiments [Ber21c, Tsu23]. expert [Aud22]. Experten [And22]. Expertenwissen [Aud22]. Expertise [Ber21e, Don22, Her20, Ryd20, SD23]. Experts [And22, Ess21]. Explode [MS21]. Exploitation [Dan21, Rob22b, See20b, Pou23a]. Explosive [Ham20]. Expo [Sch23c]. export [Kir22]. Exposed [Gro23]. Exposition [Sad24]. Exposure [AG24]. Express [Low23]. Extinct [Mar23a]. Extinction [Hah23]. Extraction [Lif23]. Extractive [Fra21]. Exuberance [Sie22]. Eye [Chu22a, His23b, Rav20]. Eyes [Kes21, Ros22a].

F [Ham20, Las23, Spi24]. F-15 [Las23]. F-16 [Las23]. Fabriano [NG20]. fabrication [Gab22]. face [Die21, HP21a]. Faced [Tho23a]. Facing [Fla23, HP21a]. factories [Hol22]. Factory [Rob22b, TT22]. Faeeza [Mch24]. Fahlstedt [Pen22]. Failed [Ber21e, Per21a, Zou22, Coo20]. Failure [Ves23a]. faire [Bau23]. Fairs [Low20]. Faith [PZ23]. Fake [Ald20]. Fall [Hei20, Mac23, Mit22]. Fallan [Bla23]. Fallout [Her23]. Falls [Jac23a, Mac20, Mon23]. Family [Cie23, Her20, O’D20, Och23]. Famine [Tho22]. Famous [Jac23a]. Fantasies [Yao21]. Far [Bim22]. Farbøl [Har23b]. Faria [Hva24]. Farman [Kal20]. Farming [Her20]. Farms [Spi24, Wil22a]. Fascism [Cal22]. Fascismo [Cal22]. Fat [KdS24]. fatto [San21]. Fay [Cur23]. Fear [HL22]. Federal [Kir22, Kub22, Uts21]. Federico [Zul24]. Feed [KdS24]. Feedback [Mar21c]. Feelings [Hes20b]. feet [Zul24]. Fehrenbach [Hol22]. Fehring [Lie23]. Fellows [Mar22b]. Fellowships [Ano21b, Ano22b, Ano23a, Gro22]. Female [Bea21, San21]. Feminism [Mai24]. Fens [Mar22a]. fer [Fon23a]. Ferber [Fre22]. Fernandez [Hes20b, Mai23]. Ferng [Sto23]. Ferreiro [Del21]. Ferrocarril [Cas21]. feu [Bel21]. ficción [Dal22]. Fickers [Sch22, Nel22]. Fiction [Mey20, Rie22, Dal22, Ves23b]. Fictional [Rof20]. Field [Kai23, Mun23]. fields [Soe21, Jac23b]. Fifteenth [Sch21a]. Fifteenth-Century [Sch21a]. Fifty [Luc20a]. Fight [HL22, Zea22, Zyl20]. Fighter [Las23]. Filing [Wir23]. Filippone [Ger21b]. Film [Bae23, Col20, Pen22, Per23a, Pop20]. fin [Gou21]. Financial [Rus22, Sch23b, Vet23]. Finland [Kaa23]. Finn [Kno20]. Finnish [Pou23b, Ert21]. Finstein [Zyl22]. Fire [HH21, Kay20, Ten22, Bel21, Tat22]. Firepower [Roy24]. firmes [Jud22]. First [Den20b, Low23, Mal24, Ram22c, BBC^{+}21]. Fisher [Che23, Cli20]. Fishery [Lev21]. Fishkin [Cli20]. Fishmeal [Bac22]. Fitzpatrick [Sch21c]. Fixing [Li24, Jac23a]. Flambard [Pou23a]. Flat [Den20b]. Flat-screen [Den20b]. Fleischman [Mit22]. Flight [Mel23, Pos23, RO23]. Flood [Buc23]. Floor [Yi22]. Flops [Coo20]. Florence [Wil21b]. Florent [Ver22b]. Florian [Pui22]. Floriana [Gui22]. Flour [Pag23]. Flow [Sca20]. flows [Dis20]. Fluss [Dis20, Huh24]. Flux [Neg22b, Huh24]. Flying [Las23]. Fonoscopy [Kvi23b]. fontaniers [RR21b]. Food [Fro23, Gol24, KdS24, Mar23b, Mue22, Ree22, Ves23b, Wil22a, Wil21b, HL22, His23a, Tho21, Lal21a]. Force [Esp23, Han23]. Forced [Zel22]. Foreign [Cha22c, Rey22, Zha20, Cal22]. Forensic [Rai23]. Forest [Amm22, Kra21b]. Forestry [Zak23]. Forged [Tys24]. forges [Bel21]. Forgotten [Tap23, Tur20]. Forino [Sca21]. Form [Gar20b]. Formation [Li24]. Forming [Har24]. Forms [EL20]. Forschung [Cam21b]. Fortschritt [See22]. Fossil [Bla20a, Ger21a]. Fotografie [Hol22]. Fought [Tho22]. Foundations [SD24, Aka22]. Foundry [Fer23a]. Fountains [Bos20, Fre20, Wac23]. Four [Cam23]. Four-Volume [Cam23]. Fournier [Yam23a]. Fragile [SD24]. fragmentation [Mic21]. Framing [Per23b]. francais [RR21b]. francaise [Bel21, Pre23a]. France [TS20, Tru24, Bau20, Bot20, Kuo23, Dup21, Kno22b, Pel24]. Francesca [Ale23]. Francisco [Pen22, Sol24b]. Franco [Lóp22]. Francois [Fon23b, HP21a, Lin22, Pou23a]. Francoise [Nèg22a]. Francoist [Cam22]. Frank [Sch20b, Tho22]. Frankreich [Gut23]. franquista [Cam22]. Fraser [Per23a]. Frawley [Har21]. Fred [Tap23]. Frederik [Huh24]. Fredrik [Li22]. Free [Bre23, Sca20]. Free-Flow [Sca20]. Freedom [McC23, Nor21]. Freely [Spi20]. Freemartin [BN23]. French [Amm22, Bel21, RR21b, Le 23, Kay20, Kra21b]. Frens [Mar23b].

Frens-String [Mar23b]. Frequencies [Kui22]. Fresh [Mac22a]. Fría [VÁ21]. Friendliness [Fer23b]. Fries [Hår20a]. Front [Mul21, Zan22]. Frühen [Rei21c]. frühneuzeitlichen [NG21]. Frumer [Wit21]. Fuel [Bla20a, Gag21, Kob22c, Kra21b]. Fueling [Wen24]. Fuels [Ger21a]. Fuller [Plo23]. Fumigation [Bol21]. Fun [Wac22]. Funded [Kif24a]. Fungi [Rua24]. Funk [Sme20]. Funke [Rei21c]. furrow [Zul24]. Furuhata [Tsu23]. Future [Cur23, Die22, Dim24, Hol21, HG24, Per21b, Rik21, SEO20, vD21, Kub22, See22, Cor20, DiM22, Hir23, Mar21b]. Futures [KMÅ20, War20]. Futurologie [See22]. futurology [See22]. Fuxi [Yan22].

G [Dal22, Dav23b, Loo22, Löv23, Mac23]. Gabel [Syl22]. Gabriele [Nel22]. Gabrys [CC21]. Galan [Bar23, Hof21]. Galili [Rog21]. Galina [Opp22]. Galloway [Gab23]. Galluzzi [Kuc21]. Galviz [Pik21]. Gamble [Kru22]. Game [Coo20, Koc23, Šve21]. Game-Changing [Coo20]. Gamers [Van20a]. Games [O’D22, RR21b, vM20]. Gaming [McD22, vM20]. Gan [Yan22]. Garcia [Arc23b, Bia22, Loz22, Wis22]. Garcia-Swartz [Arc23b]. Garcon [Nèg22a]. Garde [McC20]. Garden [KB22b, O’S20]. Gardikas [GA21]. Gareth [Wol23]. Garrioch [Gun24]. Garza [Dim23]. Garzke [Pom20]. Gas [Lif23, Tho23a]. Gatejel [Şer24]. Gauchet [Has22]. Gaudillière [Ree22]. GDR [Mac22b]. gemeenschap’ [Lin24b]. Genay [Gus24]. Gender [Bea21, Ber23, Bix24, Coa22, Fil23, Kaa23, Koc23, Lee24, Och23, Sch23b, Lab21]. Gendered [Gol24, Jam22, Mak23, Mai23]. Gendering [Wil23]. General [TT22]. Generations [Bar21b, San20]. género [Lab21]. Genetics [Ber21a]. geniali [San21]. Gente [Sol24a]. Geobodies [d’A20]. Geoff [DeC22, Erd21]. Geoghegan [Le 23]. Geology [Duq24]. George [Fer23a]. Georgia [Tay23]. Geppert [Ell22]. Geraci [Ram21b]. Géraldine [Daz23]. Gerardo [Bur23]. German [And22, Dan21, Pop24, See20b, Wen24, Den23]. Germany [Gum24, Kir22, Kra22, Kub22, Mit22, Pre21, Sch20b, See22, Soe21, Uts21, HH21, Sch23b, HF21]. Germs [Cur23]. Gershberg [Bre23]. Gerth [Zha21]. Geschichte [Bee21, Jan23b]. Gesellschaft [Dis20, Sch20b]. geteilten [Kra22]. Giacomo [Ber24, Cos20]. Gianenrico [Bau23]. Gibbs [Gil20]. Gibson [Rog23]. Gijs [Sie22]. Gilde [Lin24b]. Giles [Tur20]. Gillon [Ves20]. Ginger [Bru23]. Giraud [Bia22, Wis22]. Girls [Rob22b]. Girten [Cop24]. Giuntini [Far24]. Giving [Pat24]. Glamorized [Rob22b]. Glass [Hub24, Yan22]. Glinsky [Kar23]. Global [Ale22, Arc20, Ber22a, Bla20a, Boo22, Brü21, Bun23, Cam21a, Chu23, Dio24, Ert21, Fre22, Hai23, Hal21a, Hei20, Her23, HP21b, Hun20, Jon22a, Kif24a, Kob22b, Kru22, Lau23, Mac22a, Mel23, Pre23b, Ram22b, Rob20, Sca20, SV21, Ung22, Ver21, VF24, Wij22, Cha22b, Huh24, Jud22, VÁ21, Erd21, Lue21, Rin22]. globale [Huh24, Jud22]. Globalisation [Zai23]. Globalization [Bed22, Ste20, Bas21]. Globalizing [KdS24, Sie22]. Glyphosate [Vai23]. Go [Rog23, dZ21]. Godfrey [Hun20]. Godin [Sch20a]. Gogh [Mac24]. Going [GM22]. Gold [Duq24, Pie22]. Goldstein [Lav21]. Golubev [Vee22]. Gomez [Gue22b, Lab21]. Gone [Deg22]. Gong [Mei21]. González [Ber21d].

Good [Sch21c]. Gooday [Mcg20]. Goods [Lev20, Sta22]. Google [Cor20]. Gopakumar [Muk22]. Gordon [Cli20]. Göta [Has22]. Gouarné [Die21]. Gourhan [Sou24a, Sou24a]. Governance [Cur23, Fri22b, Wij22]. Governing [AL20, Tom21]. Government [Hei23]. Govind [Muk22]. Grace [Fil23]. Graeme [Mcg20]. grafx [Hai21]. Graham [Pie22]. Grand [Als20, AL20, BMSS24, Van20b]. Grande [Mal24]. Granja [RR21b]. Grasping [Val21]. Grassroots [Ima22]. Grayzel [Tho23a]. Great [Het23, Mul21, Pos23, Zac23, Pag23]. Greco [Tay23]. Greece [GA21]. Green [Ung22, Rin22]. Greene [Sta23b]. Greening [Lun22]. Greenland [Bou22]. Greenlees [Ven24]. Gregg [Col20]. Gregory [Pre23a]. Greifswald [Aug21]. Grelon [Nèg22a]. Grgić [Don23]. Grijalva [Cas21]. Griset [Sch22]. Gross [Den20b]. Grossman [Mun20, Sim24]. Ground [Deg22, Gil24a]. group [Mic21]. Growth [NOK22, Cow21]. Gruschetsky [Pet24a]. Guan [Mei21]. Guangzhou [Per21a]. Guerra [VÁ21]. Guerres [dCC22]. Guffey [GB22]. Gugerli [Sch21b]. Guggenheim [GMR24]. Guide [DiM22, Wil20]. Guild [Lin24b]. Guillem [Bar23]. Guillem-Llobat [Bar23]. Guins [Šve21]. Guise [Fru20]. Guitar [Haj21]. Gulf [Har21]. Gun [Cos20]. Gunner [Moo20]. Gutenberg [Cor20]. Gutta [Hun20]. Gwyn [Chu23].

H [Bou22, Cli20, Ger21c, Lie23, Mar22a, Pea22, Pom20, Ree22, Rob22a]. Habit [Hor22]. Habits [Sla22]. Hachez [Wil21b]. Hachez-Leroy [Wil21b]. Hacked [Mil22]. Hackers [Van20a]. Hacking [Cru21, Erd21, Gar24]. Haddad [Hes20a]. Hagen [Kub22]. Hagood [Boh21]. Hahn [Eva21]. Haigh [Heß22]. Hailian [Hom22]. Halevi [Rob20]. Half [Dav20, Mik20]. Half-Life [Mik20]. Hall [Gro22, Mat24, Dan21]. Halliday [Wil20]. Halvorson [Tho23b]. Hamblin [Kru22]. Hamilton [GF21, Wil22a]. Han [Bae23, Mor23]. Hand [Mei22b, Ram23]. handicrafts [Dua23]. Handloom [MO22, Ram22a]. Hands [Ber21e, Sch23b]. Hankins [Las23]. Hanlon [Cop24]. hanno [San21]. Hans [Aeb23, Hår20a, Lin24b]. Hanshu [Mor23]. Haptics [Plo20]. Harari [Nye21]. Harm [Dis21b]. Harmonization [Kou24]. Harnessed [Gui22]. Harnessing [Sch20d]. Harp [Gro23]. Harrison [Rup24]. Harry [Tri20]. Hartmann [Aud22]. Harvest [Her20]. Harvey [HL22]. Hashim [Spi21]. hasta [Dal22]. Hath [Hal21a]. Haumann [Ver22a]. Haupstadt [Eng20]. Haw [Jac21]. Hay [Kin23]. Hazardous [Chu22b, Hol21, Wil21a, Rom22]. Hazards [Her20]. Head [Aks23]. Head-Up [Aks23]. Headshots [Tho20]. Headstrong [Aud22]. Healing [Suá21]. Health [Cla21, Esp23, Gid24, Gue22b, Ign21, PW24, Ree22, Suá21, Tom21]. Hearing [Goo21, Mcg20, Mon22]. Heathcott [Sin23]. Hehmeyer [Aka22]. Heidenblad [Sör23]. Heidi [Gum24]. Heike [Bor23, Cam21b]. Heine [Kai23]. Heinrich [Aud22]. Heirloom [O’S20]. Helen [Cha21b, Hah23, Hun20]. Hélène [Nèg22a]. Helge [Pin21]. Helped [Ham20, VF24]. Hendricks [Rik21]. Henk [Fri22b]. Henk-Jan [Fri22b]. Henke [Kre21]. Henken [Loz22]. Henning [Pat23, Sch20d]. Henry [Bou22].

Héricher [Pou23a]. Heritage [DiM21, Hag21]. Herman [KB22b]. heroines [San21]. Herrmann [Fox21, Mei21]. Hervé [Gab22]. Hess [Mil20a]. Heßler [Cam21b]. Hester [Cam24, Gar24]. Heuristic [Eng22]. Hibakusha [Bun23]. Hid [Kay20]. Hidden [Pre23b]. Hide [Kan23b]. Hier [Mar24]. High [Lad22, NOK22, Pie22, San20, She20, NG21]. High-Growth [NOK22]. High-Tech [Lad22, San20, NG21]. Highland [Gid24]. Hightech [NG21]. Hightech-Material [NG21]. Highways [Zyl22, Šte22b, Pry23]. Higuchi [Her23]. Hilaire [Bas21, Fon23b, Cha22b]. Hilck [Lau21]. Hill [Zai20]. Hillard [Mac23]. Hindi [Tho24]. Hindi-Urdu [Tho24]. Hindmarch [Kay23]. Hindmarch-Watson [Kay23]. Hinduism [Ram21b]. Hintz [Don24]. Hippel [Tho22]. Hirsh [Spi24]. Hirshberg [Mit24]. Hisano [Kag20]. Hispanic [dCC22]. hispanique [dCC22]. Histoire [Bau23, Jud22, Bel21, Has22, HP21a, Pre23a, Sou24a, Tru24]. Historia [Lóp24, Sou24b, Dal22]. Historian [Kag23]. Historians [Ano21c, Edi20b, McG24]. Historical [AL20, Boo24, Cha22a, Chu22a, Eng22, HM20, Jør21, Kos21, Les20, Shi20, Sin23, Cam21b]. Historico [Mod21]. Historico-Philosophical [Mod21]. históricos [Boo24]. Histories [Fro23, Gre22, Lan23b, Moo23, Mor23, Ram22a, Rup24, Sch23a, Sme20, Sto23, GB22, ZJ22, Kai21, vE22]. historischer [Cam21b]. History [Abn23, Aka22, Ald20, Ano21b, Ano22b, Ano23a, Arc23b, Aru23, BBC^{+}21, BHJI^{+}22, Bim22, Bla20a, Bla23, BE20, Bol21, Bra20, Bri20, Cai22, Cam21b, Cam23, Cas24, Cer20, Chu23, Cig23, Coo23, Cor20, Cos20, Deg22, Di 21, Dio24, Dis21a, El 21, EL20, Fal24, Fil23, Fre20, Gag21, Gus24, Haj21, Ham20, Han23, Heß22, Het23, Hir23, Hög22, HG22, Ing22, Ion21, Joh23, Kan20, Kay20, Kli23, Kra21a, Lif23, Lin22, Lóp24, ŁPO^{+}20, Lou21, Low20, Low23, Luc20a, Mak23, Mar21c, McC21, Mei23, Mei22a, Mit22, Mue22, Nel22, Nis23, Nye20b, Nye21, Old24, Oso22, Ott21b, Par20a, Pel24, Pie22, Pop24, Pou23b, Pre23b, PÑ22, Pui22, Rai23, Ram21a, Rey22, RCCAM24, Rob22a, RW23, SS21, Sim20, Sla23, Sör23, Sta23b, Suá21, Swa21, Twa20a]. History [Van20b, Ven24, Vet23, Wie21, Wir23, Yan22, Zan22, Zdr21, dZ21, vD21, Bar21a, Bee21, Bel21, Ber22b, Dal22, Has22, HP21a, How22, Huh24, Jud22, San21, Sca21, Soe21, Sou24a, Sou24b, Ver22a, Aeb23, Ale23, Arc20, Ber22a, Boo22, But23, Col21, Cox22, Dru20, Fer23a, Fre22, Gat20, Lue21, Mas20, Moo21, Pin21, Pom20, Rus20, Rus24b, Smi20, Bar21a, Muk24]. Hobby [HF21]. Hochman [Aru23]. Hochschule [Fox21]. Hockenberry [Arc23a]. Hoenig [O’S20]. Hollis [Bot21]. Hollywood [Sha22a]. Holy [Val21]. Homburg [Rom22]. Home [Cie21, Cie23]. Homebrew [McD22]. homeland [Ber24]. Homi [Kar24]. Hommage [Nèg22a]. homme [Pou23a]. Hong [Mel23]. Honking [Har20b]. Hoof [Pui22]. horlogére [Jud22]. Horn [Pat23]. Horns [Har20b]. Horowitz [Dis21a]. Horses [Gui22]. Hospital [Cur23, San20, Suá21]. Hot [Cru21, Koc23]. House [Cie23, Hoc23, TT22]. Household [Hei23]. Housing [Bee23]. Houston [See20a]. Howard [Plo23, Ver23]. Howe [Kap20]. Hsien [Bec23]. Hsien-ch’un [Bec23]. Huacuja [Tho24]. Huai [Li24]. Huang [CW24]. Hub [Mel23]. Hubka [Cie21]. Hugh [Lau23]. Hughes [Atw23, Ove22]. Hultzsch [Par20b]. Human [Ber21a, Ros22a, ZJ22, vE22]. humanidad [Lóp24]. Humanities [Kvi23a, Lin21, Gul20]. humanity [Lóp24]. Humans [Kaa23, Sch21c]. Humboldt [Bac22]. Hundred [Fut22, How22]. Hungary [Bro21b, Gat20]. Hunger [Fut22]. Hungry [Mar23b]. Hunt [Giu22]. Hurricane [Sim24]. Hurt [Ung22]. Hush [Boh21]. Hutková [Muk21]. Hvattum [Par20b]. Hwasook [Lee24]. Hyacinth [Røs22b]. Hybrid [De 23a]. Hydraulic [Bon23, Hus23, Mac20, Bel21]. hydraulique [Bel21]. Hyperchromatic [Kes21].

Ian [Arn23a, Cra21]. Ibarra [Lóp24]. Iberian [Di 21, Wen22]. IBM [Hei20]. Ice [Bou22]. ICOHTEC [BBC^{+}21, ŁPO^{+}20]. Icon [Hei20]. Idea [Fin22, Rad23, Fal24]. Ideal [Lal21a]. Ideals [Lan24]. Ideas [Dae21, Str22b]. Identification [Kvi23b]. Identity [Fra20, Tho20, Isr23]. If [Mar21b]. Ignacio [Hum24, PÑ22]. Ignaciuk [Lab21]. ignorancia [Bar23]. II [Loe24, Cri22, Ham20, dZ21]. Illing [Bre23]. Illuminated [Bot20]. Illusory [Lad22]. Iluminated [Bot21]. im [Aug21, Aum21, Dis20, Eng20, Fox21, Hol22, Huh24, Jan23b, Kra22, Mac22b, Sta22]. Image [Hol22, LM20, Pat24, SG21]. Image-Making [LM20]. Images [Ott21b, Rob22b, Yam23b]. Imaginaries [Har23b, Sla24]. Imaginary [Don23]. Imagination [McC20, The22]. Imaginative [Kli23]. Imagined [OY23]. Imaging [Fra21, Zel23]. Imma [Sca21]. Immeasurable [Sim24]. Immigration [Loo22]. Impact [Arn23a, Cho22, GMR24]. Impacted [Rai22]. imperative [Cam21b]. Imperial [Şer24, Smi20, Ber21e, HH21, Kon21, Giu22, Han22, Jam23]. Imperialism [Cla21, Don22, Fru20, Per21a]. Important [Ven24]. Imported [Yi22]. imprese [Cal22]. Impressions [Šve21]. Improvement [Arn23b]. Improving [JR23]. Improvised [Hei23]. Incident [SC20b]. Inclusion [Gar24]. Incorporation [Vet23]. indépendance [Pre23a]. Independence [Kuo23, Ade21]. Independent [Don24, Tho20]. Index [Ano20a, Ano20c, Ano21a, Ano21d, Ano22a]. India [Arn23a, Cha20, Col21, Das24, Hai23, Jam23, Kan23a, Lan23a, Muk21, PW24, Ram22a, Rey22, Arn23b, Kan23b, Lan24, Rai22, Roy21, Cho22, Gre21, Lou21, MO22, Mel23, Muk24, Rai23, Ram21a]. Indian [Cho22, Muk22, Ott21a, Ram22b, Ram21b, Sar22, Ram23, Pra20]. Indigenous [Sha22b, Wen22]. Industrial [Atw23, Bec23, Del21, Eva21, Gou21, Hah23, Hay22, Kob22b, Lan24, LM20, Lin22, McC20, Mil20a, Mil23a, Tan21b, Zan22, vdL21, HP21a, Har21, Lan23b]. Industrialisation [Pou23b, Bel21]. Industrialisierung [Ver22a]. Industrialism [Gro21]. Industrialization [Bac22, Bel21, Ber24, Ver22a]. industrializzazione [Ber24]. Industrie [Hol22, Jud22]. industriel [HP21a]. Industriels [Gou21]. Industries [Ven24]. Industry [Cri22, Ham24, Hol22, Kob22a, Kob22c, Koc23, Loo22, MO22, Mun20, Pie22, Ram22a, Ros20, Sch23b, Wie21, Zai23, Jud22, Arc23b, Cer20, Gus24, Mac23].

Ines [KdS24]. Infection [Cur23]. Infectious [Bon21b]. Infertility [BN23]. Information [Bel23, EL20, Kan24, Kay23, Kno20, Kuo23, McC22, Le 23, Sch22, Sla23, Aum21, Man20, Wir23, Kos21, Yao21]. Informationsverarbeitung [Aum21]. infraestructuras [Pet24a]. Infrared [Fra21]. Infrastructure [Bek24, Cam24, Deg22, Edw23, Eng22, Fri22b, Hal21b, Hei23, Sin23, SC20b, Has22]. Infrastructures [Cal21, Eng20, Fla23, Kai23, Kno20, Kre21, Yam23a, Pet24a, Dio24]. Infrastrukturen [Eng20]. Ingegnosi [How22]. ingeniería [Lóp22]. Ingénieurs [Ver22b]. Ingenious [How22]. Ingenuity [Edw23, Dav23a]. Ingrid [Aka22]. Initiative [Eam21]. Initiatives [Lun22]. Inland [Bia22]. innovación [Mat21]. Innovated [Dae24]. Innovation [AH20, BE20, Cha21b, Cho22, Dae21, Esp23, Gro21, Gun24, Har23a, Ign22, Les20, Loz22, McG24, Pol21, Sch20a, Sch23c, Yoo20, How22, Mat21, Wu23, Sch20a]. Innovations [Luc20b, Rus22]. Innovative [Dav20]. Insanity [Hen21]. Inscriptions [Dac20]. Insecurities [Jam22]. Insider [Sim20]. Insight [Dae21]. Inspectors [Coo23]. Installing [Muk22]. Instant [Kal20, Ste22a]. Instituciones [Zul24]. Institutional [HP20, SD24]. Institutions [Wis22, Zul24]. Instrumental [Cho22]. Instruments [Has21]. insurance [Kra22]. Insurgent [Cru21]. Integration [Lag24, PÑ22, Wit22]. Intellectual [Bau20, Zem21, Sou24a]. Intellectuals [Bar21b]. intellectuelle [Sou24a]. Intelligence [Hal20, Nye21]. Intelligent [Kli23]. Interaction [Ram22c]. Interconnected [Sin23]. Interfacing [Plo20]. International [Arc23b, Coo23, ŁPO^{+}20, Nel22, Pea22, Sad24, Şer24, Aud22, BBC^{+}21]. internationales [Aud22]. Internationalism [Cam21a, Pea22]. Internet [Mei22a, Nol23, PÑ22, Sca23, Sch23c, SM21, Nag21, Bin23, Nel22]. Interpreting [Ott21b]. Interrupted [Ber23]. Interstate [Lut20, Zyl22]. Interstates [Pry23]. intervencionista [Mat24]. interventionist [Mat24]. Interwar [Coa22, Den20a, Hor22, Wil22b, Šte22b]. Intimate [Cho22, Swa21]. Intrigue [Cri22]. Introducing [Hil23b, Nov23a, Old24]. Introduction [Fro23, Luc20b, Tan21b, Twa20a, Rin22]. Inuit [Fra21]. Invasion [Bea20]. Invent [Zia23]. Invented [Mar21b, McG24]. Inventing [Wal21a, Str22b]. Invention [Cam21a, Hal21a, Isr23, Mac24, Sou24a, Fri22a]. Inventions [TS20]. Inventor [Dae24]. Inventors [Don24]. Invisible [Kay20, Zia23, Sme20]. invisibles [Bar23]. Iqbal [Mel23]. Iran [Cla21, Ask22, Ras22]. Iranian [Ras22]. Irby [Tay23]. Ireland [Kir23]. Irlinger [Eng20]. Iron [Arn23a, Bro21b, Cli20, Kan24, vM20]. Irving [Sol24b]. Isabel [Tho24]. Isabelle [Die21]. Isenstadt [Moo21]. Islam [Rin22, Rob20]. Islamic [Zem21, Hal21a]. Island [Aug20]. Isolation [Ada20]. Istanbul [TT22]. Itaipu [Buc23]. Italia [Cal22, Far24]. Italian [Kuc21, Mor22]. Italy [Ber22b, Cal22, Far24, Uts21, VF24, Bin24, McC23, Mil22]. Itineraries [Ger21c, SD23]. IVF [Fre22]. IX [Mac24]. Izdebski [Bar22].

J [Amm22, Ber21c, Ber21d, Cam21a, Dav23a, Giu22, Gum24, Hes20a, Hes20b, Kar24, Kui22, Mas20, Mei22a, Mil20a, Nye20a, Plo23, Rav20, Sca23, Sim24, Syl22, Ver21, Wen24]. Jack [McC22]. Jacob [Buc23, Kru22, Ryd20]. Jacobs [Bun23]. Jacqueline [Rob22a]. Jaczko [Pre23a]. Jahrhundert [Huh24]. Jaime [Dav20]. Jaime-Chaim [Dav20]. James [Cro23, Dim23, Hei20, Hir23, Jon23, OA23]. Jan [Fri22b]. Jane [BF24]. Janeiro [Mac22a]. Janet [Sch23a, Sub22, Ven24]. January [Ano20d, Ano21e, Ano22d, Ano23b, Ano24a]. Japan [Coa24, Miy22, Yan20, Abe21, Has21, Kob22c, Mon22, Shi20, Tan21b, Wit22, Abe23, Coa22, Fra20, Wit21]. Japanese [Nis23, Och23, Don22, Ima22, Kob22a, Zha20]. japanicus [Och23]. Japonaise [Eri21]. Jara [Mit23]. Jaroslav [vM20]. Jarrige [HP21a, Lin22, Fon23b]. Jason [Kal20, Rus24a]. Javier [Mit23, Whi23]. Jayita [Hai23]. Jayram [Pra23]. Jean [Ree22]. Jean-Paul [Ree22]. Jeanne [Suá21]. Jefferson [Lud23]. Jeffrey [Cer20, Hen21, Pet24b, Ren22]. Jenks [Vol22]. Jenna [Sto22]. Jennifer [CC21, Coh21, Och23, Ros22a, Sto23]. Jenny [Ber21a]. Jeremy [Bri20, Hoc23, Sta23b]. Jerry [Zac23]. Jessamyn [Low23]. Jesse [Lan23b]. Jesús [Ign21]. Jeux [RR21b]. Ji [Mei21]. Jia [Zen22]. Jill [Mar21b]. Jim [Car21]. Jinping [ZP24]. Joachim [Aeb23]. JoAnne [Jon22a]. Job [Kif24b, Poo23]. Jobs [Zac23]. Joel [vdL21]. Joeri [Mun23]. Johan [Kai21]. John [Als20, Bla21, Jac21, Kat23, LM20, Mar21a, McC20, Mel23, O’S20, Rik21, SS21, Sto24]. jokō [Rob22b]. Jolly [Bon21a]. Jonas [Kal23]. Jonathan [Coa24, HL22, Jam23, Sar22, Sin23, Tuc22]. Jones [De 23b, Wol23]. Jordan [Gil20, Hei23, Gut23]. Joris [Fri22a]. José [Dav23a, Sol24b]. Joseph [Cha21a, Fin22, Sin23]. Josephson [Pol24]. Josh [Jan23a]. Joshua [Fil23, Lav21, Mar23b]. jours [Jud22]. Joy [Amm22]. Joyce [Hol21]. Joyner [Jon23]. Jr [Pom20]. Jr. [Als20, Dae21, OA23]. Ju [Zan22]. Juan [Ste22a]. Judging [Eri21]. Judith [dZ21]. Julia [Kal23, Rei21c]. July [BBC^{+}21, ŁPO^{+}20, Ano20g, Ano21g, Ano22f, Ano23d]. Jun [Naq22]. Junction [Ade21]. Jungle [Pre23b]. Jupiter [Kes21]. Jürgen [Nis23]. Justice [SM21, Pry23]. Justin [Dim23, Lin24a, Ver21]. Jütte [Fal24].

Kaffee [Bee21]. Kai [Naq22]. Kaijser [Cig23]. Kalkstein [Ver22a]. Kalmbach [Kno22b]. Kalten [Aug21]. kam [Sch21b]. Kansas [Fre20]. Kanta [Kli23]. Kao [Mei21]. Karen [Mcg20]. Karena [Kno22b]. Karin [Lau21, Bim22]. Karl [Zha21]. Karolina [Muk21]. Karsten [Hår20a, Pop24]. Karuka [Sha22b]. Karuna [Ram22a]. Kasper [Kra22]. Kate [Ash21, DiM22]. Katerina [GA21]. Katherine [How23, Sch21c]. Kathke [Cat21]. Katie [Kay23]. Katko [Ert21]. Katowice [ŁPO^{+}20]. Katrin [Gut23]. Katrina [Dis21a]. Kazakh [Sei24]. Kelley [Col20]. Kelly [Arc23b]. Ken [McC21]. Kenneth [Jan23a]. Kerim [Yan20]. Kernfusion [Mey20]. Kernkraftwerk [Aug21]. Kerr [Ada20]. Kevin [Gil20, Mil23b]. Key [Pol21]. Keyser [Ves23b]. Khan [Str22b]. Khipus [RCCAM24]. Kideckel [His23a]. Kilimanjaro [Dis21b]. Killed [Tho22]. Kim [Gid24, Pen22, Bae23, Gie22]. King [Kag23]. Kingdom [Che22]. Kirchhelle [Mue22]. Kirchhoff [Uts21]. Kirsten [Bee23]. Kiryushina [Opp22]. Kisacky [Suá21]. Kit [Atw23, Ott21b]. Kitchen [Kag23, Sca21]. Kjetil [Bla23]. Klein [Rie22]. Knapp [Sad24]. Know [Ben22]. Knowledge [Bar21b, Bon21b, Dis21b, Don22, Duq24, Esp23, Gou21, GL23, Gün21, Gun24, Hal21b, Huh24, JR23, Kai21, Kai23, Kif24a, Rai22, Ryd20, SV21, SS21, Sei24, Str22b, Tan21b, Wen22, Whi23, ZJ22, Aud22, Gab22, dCC22, Mak23, Sör23, Kat23]. Knowledge-Making [JR23]. Knowledges [Ger21c]. Ko [Naq22]. Kodachrome [Pop20]. Koeppe [Mea21]. Köll [Tan21a]. Kolonialtechnik [And22]. Kong [Mel23]. Konrad [Mei21]. Konsum [Rei21c]. Korea [Bae23, Yoo20, Lee24, OY23]. Kouli [Lag24]. Koyagi [Ras22]. Kraft [NG21]. Krajewski [Joh20]. Krakow [Bar22]. Krebs [Bor23]. Kreditwirtschaft [Kra23]. Krieg [Aug21]. Krige [Kat23, SS21]. Kristian [Bou22]. Kristin [Bac22, Cop24]. Kristina [Bix24]. ’Kritischer [Ver22a]. Kultur [Rei21c]. Kulturgeschichte [Wag23]. Kulturpflanze [Soe21]. Kurlat [Dal22]. kybernetischen [Aum21]. Kyle [See20a, Yam23a].

L [Ada20, Coh21, Dru20, Fin22, Gro23, How23, OA23, Pre21, Ram22c, Tay23, Vol22]. Laak [Dio24, Dis20]. Laan [Nye20a]. Lab [Kai23, Sim20]. Labor [Aks23, Ber21e, Gol24, Lee24, Li24, Twa20b, Zan22, Atw23, Gro23, Rus24a]. Laboratory [Kon21, Wil23, Cho22, Sim20, KM23]. Laborers [Hom22]. Laborie [Lag24]. Labour [Sar22, Zel22]. Labyrinths [Wac23]. LaCouture [Cie23]. Ladies [Lau21]. Lady [Lau21, Lau21]. Ladyem [Koc23]. Laine [Loe24]. Lamata [Zap20]. Lamps [Bot20]. L’analogie [Duh22]. Lancashire [Ven24]. Land [Gus24]. Ländliche [Aud22]. Landline [AH20]. Landscape [Kru20, vdL21, GF21, Hal20]. Landscapes [Ach22, Ash21, El 21, Nor23, Per23b, Zel23, Tar23, GA21, Wel20]. Language [Cas24, Gil24b, Wen22, WAB^{+}23, Pet24b]. Languages [Cha22c]. L’animal [Pou23a]. Lantern [Bon21a, Neg22b, PZ23]. Laptop [Löv23]. Larissa [Bas21]. Larrie [Del21]. Larsson [Sör23]. Last [Gal22]. lastre [Cam22]. Late [Cub23, Hal21b, Kai23, Kri24, Vee22, Gou21, Cam23, Har20b, Sei24]. Late-Colonial [Har20b]. lateinamerikanische [Huh24]. Latin [Dal22, GC24, Huh24, Pet24a, Pin21, SD23, Cas24, Oso22, RCCAM24]. Latinoamérica [Pet24a]. latinoamericana [Dal22]. Latter [Bea21]. Latter-day [Bea21]. Lauer [Jan23a]. Launch [GX24, Han23]. Laura [Gol24]. Lauren [Mit24, Sto23]. Laurent [Kou24]. Lavelle [Rei21b]. Law [Bro21a, Bur23]. Lawrence [Sim20]. Layered [Sie22]. LCDs [Den20b]. Lean [Gro21]. Learn [McG24]. Learning [Bro22, Cai22]. Lebensadern [Dis20]. Led [Den20b, Zai23]. Lee [Ess21, Han20]. Legacy [Bla21, Löv23]. Legal [Eri21]. Legged [Hor22]. Leisure [BE20]. Lemberg [Sca20]. Lemercier [Gul20]. L’énergie [Mar24]. Lenk [PZ23]. Lenka [Hol22]. Lenny [Neg22b]. Lens [Ott21a]. Lente [Cor23b]. Leo [Fri22a]. Léonard [Lag24]. Leonardo [Jon23, Sca23]. Leong [Mak23]. Leor [Rob20]. Lépinay [Wal22]. Lepore [Mar21b]. Leroi [Sou24a, Sou24a]. Leroi-Gourhan [Sou24a, Sou24a]. Leroy [Wil21b]. Leslie [PW24]. Lesson [Aks23]. Lessons [Hir23, Twa20b]. Let [vdS22]. Lethal [Ald20]. Letters [Coh21]. Levasseur [Ros22a]. Leviathan [Mch24]. Lewis [Coa22]. Li [Zan22]. Liberalisation [Bou23]. Liberation [KJ23]. Liberatory [The22]. Library [GM22, Gro22]. Lieberman [Coh21]. Lies [Sch20c]. Life [Ask22, Bim22, Cam24, Coh21, Cub23, Gre21, Jac21, Jon23, Les24, Löv23, Mik20, Mon22, Sar22, Vee22, Kar24, The22, Van20a, Cub23]. lifeblood [Dis20]. Lifelines [Dio24]. Lifeways [Gil20]. Light [Bri20, Mon23, Moo21, PZ23, Rup24, TT22, Vik20, vdS22]. Lighthouses [Lev20]. Lighting [Bot21, Wal21a]. Liliane [Bas21, Cha22b, Fon23b]. Limestone [Ver22a]. Limiting [Ell22]. Limits [Pra23, Sta23b]. Linda [Gro22]. Lindsay [McC23, Nol23, Ros20]. Lindtner [Wu23]. Line [Car21, Mal24, Per21a, Tat21]. Lines [Coh21]. Lingold [Bit21]. Lino [Mit23, VÁ21]. Lintsen [Tri20]. L’invention [Sou24a]. Lipartito [Jan23a]. Lipton [Neg22b]. Lira [KM23]. Lisa [Wel21]. Lisboa [Sou24b]. Lisbon [Sou24b]. Listeners [Aru23, Coh23]. Listening [Cam21a, Mon22, Mun23]. Literature [Cop24, Per23a]. Liu [Lou21, Yao21]. Livermore [Sim20]. Lives [Cho22, SM21]. Livestock [Sad24]. Living [Fou23, WAB^{+}23, Hum24]. livre [Cha22b]. Liz [Moo20]. Llave [Whi23]. Llobat [Bar23]. Loath [Ver23]. Lobo [Hva24]. Local [Gro21]. Localized [JR23]. Lockhart [Roy24]. Loewy [LM20]. Logan [Rey22]. Logistics [Arc23a]. Loire [Gui22]. Loma [Loo22]. London [Dav20, Kay23, Pik21, War23, Wil20, Hor22, Ada20]. Lonely [Hes20b]. Long [Bro21a, DeC22, De 23a, Gab23, Gal22, Hof23, SD23, Wri21, Hal21b]. Longo [Gil24b]. longue [Bau23, Pre23a]. Look [Kag20, Koe23]. Looking [Hil23b]. Lopes [Mac22a]. López [Whi23, Cam22, Dal22, Pik21]. Lorek [SD23]. Lorenzo [Cam22]. Lori [Sch20d]. Lorkowski [Sta22]. Lorraine [Fon23a]. Loss [Mcg20]. Lost [Kru20, NOK22, Rob22c, ZJ22]. Lost-Wax [Rob22c]. Lothar [Ryd20]. Louise [Ber22a]. Louisiana [Bla21]. Love [Bro22]. Low [Miy22]. Lower [Şer24]. Lucas [Pry23]. Luce [Ash21]. Luci [Sha22a]. Lucia [Sta23a]. Lucie [Gus24]. Lucifers [Bri20]. Lukas [Bol21, Gre22]. Luke [Hes20b]. Luminita [Şer24]. Lungs [Gue22b]. luogo [Sca21]. Lupack [Tel20]. Lure [Boo22]. Luxury [McN24]. Lynteris [Bol21]. Lyons [Ing22]. Lyubomir [Bek24].

M [Cam22, Cha21a, Cop24, Han20, Hay22, Hof23, Kin23, Mun20, Nye20a, Pet24b, Wu23]. Ma [Sco23]. Maarten [vE22]. Macao [Per21a]. Macedo [Sou24b]. Macfarlane [Jac23a]. Machine [Ban21, Löv23, Mar21c, Naq22, Ram23, WAB^{+}23, Cor23a, Con22, Ham21]. Machinery [Yi22]. Machines [Bos20, Bru23, Cai22, Fer24, Fru21, Hal21a, Mai23, Sad24, Sch21c, Twa20b, Wac22, Kli23, Kuc21, Šve21]. Mack [Boh21]. Mackowski [Cub23]. Macola [Cos20]. Made [Ber22b, Brü21, Cha21b, Fer23a, Smi23, San21]. Maggi [Mor22]. Magic [Bon21a, Neg22b, PZ23, Tel20, Wor24]. Magnet [Rad23]. Magnetic [Pat24].

Magnusson [Dac20]. Maine [Mac23]. Mainstream [Rog23]. Maintaining [WJ23]. Maintenance [Bor23, KB22a, Kre21]. Maize [Hah23]. Majo [Het23]. Majsova [PZ23]. Make [Che23]. Makers [Pat24]. Making [Ber21a, Ber21c, CC21, Dav23a, Goo21, Gro21, Has21, Her23, How23, Ism23, JR23, Kay23, Lee24, LM20, Mar21c, Mar23b, Mel23, Poo23, PW24, Sta23a, Tom21, Wei20, Bra20, Bou23, Cer20, GB22, Mea21, Naq22, Sad24, Tys24, Wit21, vdL21]. Maladjustment [Ign22]. Malaria [GA21]. Mali [Twa20b, Gol24]. Mallea [Fre20]. Man [Koc23]. Management [Aka22, Di 21, Dis21b, EL20, Kus23a, Sta20]. Managing [FV20, Mcg20]. Mandelman [Swa21]. Maniac [Mar21c]. Maniac-Making [Mar21c]. Manning [Jam22]. Männistö [Sme20]. Männistö-Funk [Sme20]. Manu [Sha22b]. Manual [DiM22]. Manuel [Lóp22, Zap20]. Manufactory [Fer23a]. Manufacture [SC20a]. Manufacturing [Har21, Luc20b, Ros20]. Mao [ZP24]. Maoist [Li24]. Mapping [Lev21, RW23, Gre22]. Marcaida [Dav23a]. Marcelo [Boo24]. marche [Pre23a]. Marching [Rog23]. Marcio [Het23]. Marco [Ber22b, Cal22, Loo22]. Mardones [Boo24]. marées [Mar24]. Margaret [And21, Lun23, Wen22]. Margins [Rog23, Nol23]. Marguerite [Cur23]. Mari [Par20b, Wie21]. Maria [Ham24, Mac22a, Mei23, Ign21]. Maria-Aparecida [Mac22a]. Marie [Cur23, Har23b, Pou23a]. Marín [Di 21]. Marina [Gar20b]. Mario [Kat23, Wal22]. Marissa [Kui22]. Maritime [Bol21]. Mark [Gal22, Mor23, Opp22]. Market [Mac22a, Bee21, Mat24, Zak23]. Marketed [Dae24]. Marketing [Wie21]. Marketplace [Ham24]. Markets [Bou23, HP21b, War20]. Märkl [Mey20]. Markus [Joh20]. Marques [Mei23]. Marr [Dav23a]. Marta [Sou24b]. Marti [Aug21, Aug21]. Martin [Arc23b, Die22, Kai23, Kra23, vD21]. Martina [Cam21b, Naq22]. Martocci [Zul24]. Martyn [Bon21a, Fra20, Ing22]. Marvelous [Sto24]. Marvels [Mea21]. Mary [Bit21, Yu21]. Marzola [Sha22a]. Masculine [Hes20a]. Mask [Tho23a]. Mass [Yi22, GX24, Fra20, Yu21]. Massenmedium [MF20]. Masses [Ass23]. Master [TS20]. Masters [Sta20, PW24]. Matala [Pou23b]. Material [Aka22, Dac20, Fal24, Map23, NG21, Rei21c, Rob20, Soe21, Ver22a, Hol21]. Materialities [DeC22, Man20]. Materiality [Kre21, Vee22]. Materials [Ger21c]. Materielle [Rei21c]. Maternal [Got20]. Mathias [Eng20]. Matt [Hes20b]. Matter [Dav23a, Fou23, RR21a, SM21]. Mattered [Cal21]. Matters [Oso22, Rus22, Nor23]. Matthew [Ada20, Arc23a, Dis21b, Vir22]. Maura [Cub23]. Mauricio [Lóp24]. May [Mar21a]. Maya [Kam22]. Mazierska [Ves20]. McCray [Tys24]. McDivitt [Koc23]. McGarity [Boy21]. McGuire [Wil22b]. McIlwain [SM21]. Meades [Abn23]. Meanings [Kno22b]. Measurement [Wit21]. Measuring [Wil22b]. Meat [Mac22a]. Mechanical [Dup21, Lie23, Ves23a]. Mechanics [Bos20]. Mechanization [Hor22]. Media [Atw23, Boh21, Bon21b, Bre23, Coa24, Fra20, GC24, Hai21, Joh23, Mac24, Mar21c, Pui22, Sch21c, Tsu23, Zai20, Yam23a, Arc23a, Mil23b, Pat23, Sca20, Sto24, Ass23]. mediale [Gut23].

Mediating [Sco23]. Mediation [Bro22, Yao21]. Medical [How23, Kif24a, Mur20, PW24]. Medicalization [San20, Shi20]. Medicine [Lue21, Muk24, PW24]. Medieval [Luc20a, NG20, Wal21b, Lal21b]. médiévale [Lal21b]. Medium [Sto22, vM20]. medzivojnovom [Šte22b]. meeting [BBC^{+}21]. Megacities [Cow21]. Megan [Gre21, Kno20]. Meiji [Rob22b, Wit22]. Meiji-Period [Rob22b]. Meiske [Kai23]. Meiton [Li22]. Melanie [McD22, Smi23]. Melina [Pet24a]. Melzer [Nis23]. Memex [Les20]. Memorabilia [Ove22]. Memoriam [NOB^{+}20]. Memories [Jac23b, ZK22]. Memory [Cor23a, Hag21, Wen20]. Menaces [Wil21b]. Meng [Zak23]. Meniketti [Loo22]. mercado [Mat24]. Mercelis [Fri22a]. Merchant [Pol21, Lin21]. Merchants [Hom22]. Merit [Ang23]. Meserve [Lun23]. Mesoamerica [Bea20]. Mestizo [Ber21b]. Metallurgy [Wen22]. Metalworking [Rob22c]. Meteorological [Sim24]. Methods [Gul20]. Métiers [Cue22]. Meuse [Gui22]. Mexican [Ber21d, Gus24]. Mexico [Mai23, Mat24, Sol24b, Wen24, Mat24, Ber23, Mon23, Ber21b, Dim23, Ver21, Sol24b]. MfS [Aum21]. Micaela [Jon23]. Michael [Coo20, Edw23, Fer23b, Han22, Har21, His23a, Las23, Mac23, Mei22b, Mor23, O’D22, Ove22, Rus20, Sch23c, Str22a, Wol23]. Michal [Šte22b]. Michel [Ver22b]. Michelle [Che23]. Microelectronics [Bro21a]. Microfilm [Bel23]. Microorganisms [Cow21]. Microscope [Wil23]. Microscopes [Wil23]. Mid [Bix24]. Mid-Century [Bix24]. Midcentury [Han20]. Middle [AD22, Brü21, Cam23, ZK22]. Middle-Class [ZK22]. Midlands [Fer23a]. Midwives [Got20]. Mighty [Mac23]. Mignon [Uts21]. Migrant [Ber21e]. Migrations [Wal22]. Mike [De 23b]. Mikiya [Ras22]. Milburn [Van20a]. Mile [Aug20]. Militarization [Mit24]. Militarizing [Ell22]. Military [Bou22, Han22, Kon21, McC20, Mil23a, Rav20, Rua24, Wal21b]. Military-Industrial [McC20, Mil23a]. Milking [Kaa23]. Mill [Loo22]. Millar [Che23]. Millennia [Ram22c]. Miller [KB22b]. Millions [Tho24, Tho22]. Mills [Wit22]. Mimesis [Coa24]. Minds [Ber21e, Bru23]. Mines [MS21]. Ming [GX24]. Miniature [Bel23]. Mining [Bar21b, GMR24, Gue22b, Hom22, MS21, Mun20, Wen22]. Miniskirt [Bea21]. Miño [Cas21]. Mint [Fer23a]. Mircea [Ram22b]. Miriam [BF24]. Misa [Sca23]. Miskell [Ber22a]. Miss [KJ23]. Mission [Den20a, Sto22]. Missionaries [Mon23]. Mississippi [Swa21]. Missouri [Fre20]. Mitcham [Mod21]. Mitman [Col20]. Mizuno [Fru20]. Mobile [Arc23b, Hor22, VF24, Gué22a]. Mobile-Phone [Arc23b]. Mobilités [Cha21c]. Mobilities [Car21]. Mobility [Cha21c, Die22, Gun24, Har20b, Muk22, Nor20, Poo23, Ras22, Sch21a, SEO20, Sie22, Zyl20, Zyl22, Boo24, Bae23]. Mobilized [Mil20a]. Mobilizing [Den20a, Naq22, Yam23c]. model [Aud22]. Models [Cha22a, Mac20, Sch20a, War20, Sch20a]. Modem [Mil23b]. Modern [Abe23, Ass23, Ben22, Cha21a, Cie21, Cla21, Cox22, Dav23a, Fer24, GA21, GB22, Goo21, Har24, Heß22, HP21b, His23a, How23, Hur20, Lee24, Luc20b, Mar22a, Mar23b, Mel23, Mul21, Pop24, Ram21a, Rei21c, SC20a, Suá21, Tan21b, Ves23b, Wen22, Wen24, Yan20, Yu21, vdL21, Gra22, Lal21b, NG21, Wei20, Moo20, FR21, Rob20, Zyl22]. moderne [Lal21b]. modernen [NG21]. modernes [Gra22]. modernidad [Dal22]. Modernisierung [Hår20a]. Modernism [Bix24, Han20, Wån23, Les24, Mei22b]. Modernities [Got20, Pik21, SG20, Ten22]. Modernity [Ber21b, Don23, HF21, Jam23, JR23, Luc20b, Ram21b, Tom21, Wan21, Dal22]. modernización [Cam22, Cas21]. Modernization [Dav23b, Den23, Har24, Kan24, Cam22, Cas21, Dua23]. Modesto [Ver21]. Mody [Bro21a, Hof23]. Molard [Fon23b]. Molly [Hal20]. Mom [Sie22]. Moment [Ale22]. monarchie [dCC22]. monarchy [dCC22]. monde [Cha22b]. Mondschein [McC21]. Monopolies [War20]. Montaño [Wen24]. Montgomerie [Sto22]. montres [Jud22]. Monuments [Sta23a]. Moog [Kar23]. Moon [Bim22]. Moore [Bro21a, Fru20, Rup24, NOB^{+}20]. Moorman [Kui22]. Moped [BE20]. moral [San20]. Morgan [Löv23]. Moriguchi [RO23]. Morning [Jon23]. Morris [Miy22, Bla20b]. Morriss [Rei21a]. Mortality [Got20]. Morton [Bee23, Hir23]. Most [Hol21, Jac23a, How22]. Motherhood [Che23]. Motion [Bos20, Neg22b, Ras22]. Motor [Hår20a]. Motors [Fil23]. moulins [Gra22]. Mountain [Gie22, Gie22]. Move [Tur20, Boo24, Vai23]. Moved [Rog23]. Movement [Tat21, Eng20, Wri21]. Movements [Mil20a, Sch21c, Zai20]. Moves [SS21, See20a]. Movie [Smi23]. movilidad [Boo24]. movimiento [Boo24]. Moving [SG21, Wac23, Ess21]. Moyen [AD22]. Mozambique [Per23b, Tho20, Bee23]. Mr. [Ves23a]. Mud [Bra20]. Mud-making [Bra20]. mujeres [Lab21]. Muka [Hub24]. Mukharji [PW24, Nai22]. Mullaney [Tat22]. Muller [vdL21]. Mulry [Ash21]. Multinational [Mai23]. multinationals [Cal22]. multinazionali [Cal22]. Multispecies [Kaa23]. München [Eng20, Fox21]. mundial [VÁ21]. Munich [Eng20, Fox21]. Munns [Bim22, Gär20a]. Murdoch [Fer23a]. Murphy [Jon22a]. Murray [Kan20]. Musealization [Wen20]. Museum [Cue22, Kon21, GL23, De 23b]. Museums [Low20]. Music [Per23a, Ves20, Yam23a, Yam23a]. Musical [Mon22]. Muslims [Gre21]. Musterschüler [Aud22]. Myllyntaus [Sme20]. Mystics [Gün21].

N [Emm21, Gué22a, Han23, Jon22a, Ree22]. Nada [Tho20]. Nai [CW24]. Naked [Lau21]. Nam [Lee24]. Nanny [Gie22]. Nanoscale [Cho22]. Não [Tho20]. Narratives [Jac23b, Kli23, Les20, Wan21, Nye20a]. NASA [Cub23, Sub22, Wri21]. Natalie [KM23]. Natalija [PZ23]. Natasha [Aug20]. Nathan [O’D20, Sou24a]. Nation [Ale22, Aug20, Cat21, Har24, Wu23, Li22, Och23, Wel20]. National [Bou21, Cue22, Fox21, Fra20, Isr23, Kat23, Sim20, Eng20, Kir23]. Nationalism [Ram21b]. Nationalsozialismus [Fox21]. Nationalsozialistischen [Eng20]. Nations [Sha22b, Jud22]. Native [Ott21a]. Natives [AH20]. NATO [Lun22]. Natural [Lif23, Mil20b]. Nature [Ach22, Bla23, Dah23, Eri21, GF21, Mac20, Naq22, Rei21b, Smi20, TS20]. nautical [Bar21a]. Navajo [Wel20]. Naval [Del21, Jam22, Plo23, Rei21a, Daz23]. Navigating [And21]. Navigators [Har23a]. Navy [Wol23]. Nawata [Aeb23]. Nazi [Ove22, See20b, Zel22]. Neal [Sad24]. Necaxa [Mon23]. Negated [Zha21]. Neglected [Moo23]. Negotiating [Wan21]. Nègre [RR21a]. Neighborhood [Pen22]. Neil [Cox22]. Neither [Ale22]. Neoliberal [Han20]. Nerdinger [Fox21]. Nerds [Rog23]. Netherlands [Poo23, Tri20]. Network [Sto22, Kir23, Ber21d]. Networked [Cal21, Nag21]. Networks [Ach22, FV20, PÑ22, Wij22, Sto23]. Neuzeit [Rei21c]. Newman [O’D22]. News [Gum24]. Newsom [Ada20]. Newspapers [Gre21]. Next [Gal22]. Niagara [Jac23a, Mac20]. Nicholas [Pos23, Smi20]. Nickelsen [Bim22]. Nicolas [Pou24]. Nicole [Arc23a, Ott21a, Ver23]. Nielsen [Bou22, Bou22]. Nieto [Bar23, Hof21]. Nieto-Galan [Bar23, Hof21]. Nigeria [Hei23]. Nigrin [Ďur23]. Nile [Rin22]. Nina [Sta22]. Ninan [Col21]. Nineteenth [Bec23, Chu21, Dae24, DeC22, Kai23, Par20b, Pik21, Rei21b, Tuc22, Boo24, Bon23, FV20, WAB^{+}23, d’A20]. Nineteenth-Century [Bec23, Chu21, Dae24, Rei21b, Tuc22, Bon23, WAB^{+}23, d’A20]. Nintendo [O’D20]. Nitrate [GMR24]. Nixon [Opp22]. No [Amm22, Dis21b, Poo23, Swa21, Poo23]. Noémi [AG24]. Nolan [OA23, Bru23]. Nooney [Loe24]. nor [Ale22]. Nord [Has22]. Normal [Wil22b]. Normative [Shi20]. North [Arn23a, Gro23, Spi20]. Northeast [Don22]. northern [Has22, Sch23c]. Norton [Lad22]. Norway [Koc21]. nos [Jud22]. Nostalgia [Las23]. Notfall [Aug21]. Nothing [Tho20]. Novelty [McN24]. noventayochismo [Lóp22]. nucléaire [Pre23a]. Nuclear [Ade21, Bin24, Cig23, Gus24, Hai23, Her23, Hög22, Hol21, Ion21, KJ23, Kir22, Kno22a, Kru20, Kru22, Mai24, Miy22, OA23, Pre21, Pre23a, Sla24, Uts21, Wen20, Wol23, Aug21, Bun23, Pol24, Rib23]. nuevo [VÁ21]. Nuklearer [Aug21]. Nuklearexportpolitik [Kir22]. Number [Ano20e, Ano20d, Ano20g, Ano20h, Ano20f, Ano21f, Ano21e, Ano21g, Ano21h, Ano22e, Ano22d, Ano22f, Ano22g, Ano23c, Ano23b, Ano23d, Ano23e, Ano24b, Ano24a]. Numbering [Wil22b]. Numbers [Bou21]. Nurses [Mul21]. Nursing [Mul21]. Nye [Tar23]. Nylan [Mor23].

O [Hal21b]. období [Šte22b]. Obertreis [Kal23]. Objectivity [Kvi23b]. Objects [Kou24, Naq22, Mar23a]. Observatory [She20]. Obsolete [Mar23a]. Occupational [Chu22b]. Océan [Amm22, Lev21]. Oceans [Hub24]. October [Ano20h, Ano21h, Ano22g, Ano23e]. Oden [Her20]. Odom [Wri21]. Oertzen [Mak23]. oeuvre [Nèg22a]. Office [KB22a, KB22b]. Oil [Bla21, GL23, Lut20, Str22a, Wal21a, Soe21, Cam24]. Old [Dup21, Hög22]. Oldest [Mei21]. Ölfelder [Soe21]. Olive [GL23]. Olshin [ZJ22]. Omère [RR21b]. On-the-Job [Kif24b]. One [Chu22a, Löv23, Zan22, BMSS24]. Online [Mar22b, dWB22]. Oosterhoff [Dav23a]. Open [Bre23, Gar24, Nor21]. Opened [Bab22]. Opening [HH21].

Operated [Šve21]. Operation [Hil23b]. Operational [Pom20]. Oppenheimer [Sla24]. Opportunity [Hun20]. Optical [Ban21]. Oral [Lan23b]. Oranburg [Vet23]. Orange [Kin23]. orden [VÁ21]. Order [Cha21b, Dah23, McC22, Pea22, VÁ21]. Orders [Joh20]. Ordinary [Kus23b]. O’Reagan [See20b]. Organisation [Zel22]. Organization [Ram22c]. Organizations [Sub22]. orígenes [Dal22]. Origins [Bla20b, Bou23, Cra21, His23a, Lau23, Ros20, Spi24, Sto22, vdL22, Dal22]. Orleans [Jon23]. Orsini [Bin24]. Ortiz [Lab21]. Osman [Mei22b]. Osseo [Ade21]. Osseo-Asare [Ade21]. Osses [Lóp24]. Österreich [Wag23]. Östling [Kai21]. Other [Kag23, Mik20, ZJ22, Wil21b]. Otto [Kri24]. Ottoman [Gün21, Hus23]. Our [Che23, Cor20, Dio24, El 21, Tho22, Dis20]. Outer [Ell22]. outil [Gab22]. Overlooked [Spi24]. Overview [Cam23]. Own [Ber21d, El 21, Hei23, Tap23]. Oxygen [Esp23]. Özge [Har24].

P [Amm22, Bim22, Dae21, Gär20a, Hur20, Nis23, Ver22b, Wri21]. Pâris [Daz23]. Pac [Koc23]. Pac-Man [Koc23]. Pacey [Ale23]. Pacific [Lev21, Rua24, Mit24]. Packaging [Wil21b]. Packer [Hoc23]. Padovano [Bia22]. Padua [Bia22]. Pageant [KJ23]. Pain [Mul21]. Pál [Gat20]. Palace [Naq22]. Palate [Tho21]. Palestine [Li22]. Palette [Kes21]. Palisades [Lut20]. Pamela [Ger21c, Hal21b, Han20]. Pampa [Zul24]. Pan [Lal21a]. Panama [Bes22]. Pandemic [Cue22, Edi22, GM22, Rai22, ZK22]. Panic [BN23, Par20a]. Pankaj [Cho22]. Paola [Dup21]. Paolo [Kuc21]. Papal [Lun23]. Paper [Har23a, Mac23, Mak23]. Papermaking [Lan24, NG20]. Paradigms [Ban21]. Paradox [Bre23]. Parallel [Sme20]. Paris [Bot21, Dav20, Pik21, Wil20, Eri21, Wei20]. Parisi [Plo20]. Park [Lut20, Rad23, Lut20, Wan21]. Parry [Sar22]. Parta [Coh23]. Participation [Ima22]. Pascal [Fon23a, Sch22]. Pascale [Kri21]. Passing [MSH20]. Past [Bon23, Cur23, Rik21, SEO20]. Patent [Bau20, Bur23]. Patenting [Ima22]. Patents [Bel23, Str22b]. Pathways [Fri22b, Uts21]. patria [Ber24]. Patrick [Poo23, Rib23, Tys24]. Paul [Ald20, Cra21, Haj21, Heß22, Man20, Mul21, Pol24, Ree22, Roy24]. Paula [Mai23]. Paye [Mic21]. paysans [Gui22]. pc [Hai21]. pc/core/turbo/engine/grafx/16/cdrom2/super/duo/arcade/rx [Hai21]. Peace [Coo23, Kon21, Pea22, RO23, Vol22, Zai20]. Peaceful [Kru22]. Peasant [Sei24]. peasants [Gui22]. Peder [Koc21]. Pedro [Bia22, Wis22]. Pellisa [Dal22]. Peng [Rob22c, Rob22c]. Penicillin [Ign21]. Peninsula [Di 21]. Penner [Mar23a]. pensamiento [Sol24b]. Pensar [Pet24a]. Penser [Nèg22a]. pension [Kra22]. People [Als22, Dae24, Naq22, Swa21, Val21, Lin24b]. pequeño [Mat24]. Perception [Rav20]. Perceptual [Aks23, Ban21]. Perch [Lev21]. Percha [Hun20]. Pérez [Bas21, Fon23b, Mit23, Cha22b]. Perform [Sch21c]. Performance [Wor24, Kri21, Aeb23]. Perilous [Bre23]. Period [NOK22, Rob22b, Šte22b, Wil22b]. Periodical [Hva24]. Periphery [Koc21]. Perle [Eri21]. Perón [Als22]. Persistence [Bor23]. Personal [Fer23b, Loe24]. Personalized [Cai22]. Perspective [Aeb23, AL20, HM20, HP21b, Mil20b, Ert21, Jør21]. Perspectives [Bea20, Rik21, Ram22c]. Perspektiven [Jan23b]. Persuading [Bon21a]. Persuasion [Bre23]. Peru [Ram22c]. Peter [Bed22, Bee21, Bes22, Das24, Gil20, Lad22, Rei21b, Wal21b, Zia23]. Peters [Sto24]. Peterson [Kam22]. Petrocelli [San21]. Petroleum [Lut20]. Petrov [Kan24]. Pharmaceutical [Pie22]. Pharmacy [Ben22]. Philip [Gal22, Wac22, Wor24]. Philippe [Gra22, Pre23a]. Phillip [Pol21]. Phillips [Cub23, Pry23]. Philosophers [Bos20]. Philosophical [Mod21]. Philosophy [Mil20b, Mod21, Sto24]. Phone [Arc23b, Ber21d]. Phones [VF24]. Phonograph [Yan20, Ves23a]. Photograph [Chu22a]. Photographic [DeC22, Gab22]. Photographing [Ros22a]. photographique [Gab22]. Photographs [Kus23b]. Photography [Ang23, Arn23b, Col20, Pop20, Wei20, Hol22, Ott21a, Con22, DeC22]. Physical [Gar20b, Hof23, Cha22a]. Physician [How23]. Physicians [Val21]. Physics [Giu22]. Phytotrons [Gär20a]. Piave [Ber24]. Piazza [Mac24]. Picture [Neg22b]. Pierre [Fon23b]. Pierre-Yves [Jud22]. pies [Zul24]. Piglia [Pet24a]. Pigment [NG21]. Pigs [Mit22]. píldora [Lab21]. Pilgrims [Mon23]. pill [Lab21]. Pilot [Las23]. Pilots [Tap23]. Pink [Nag21]. Pioneer [Koc21]. Pioneering [Jon23]. Pipe [Kam22]. Pirani [Bla20a]. Pirate [Har20b]. pirogues [Daz23]. Pittsburgh [Rib23, vdL21]. Place [Swa21]. Places [Kai23]. Planet [CC21]. Planners [MK23]. Planning [Gli23, Har20b, Kuo23, Pos23, vdL21]. Plant [Cha21b, KB22a, Zou22, Aug21, Soe21, DiM21]. Plants [dZ21]. plaque [Gab22]. Plasticidade [Mei23]. Plasticity [Mei23]. Plastics [Mei23]. plate [Gab22]. Play [Gab23]. Playing [Wor24]. Pleasure [Par20a]. Plotch [Gal22]. Plotnick [Par20a]. Ploughshares [Lip22, Hai23]. plumbers [RR21b]. Plurality [She20]. Poisoning [Hay22]. Poisonous [Ste20]. Poland [Lab21, ŁPO^{+}20, Ign22]. Policies [Lag24, Sca20, Gou21, Mat21]. Policing [Nor21]. Policy [Bab22, Eam21, Got20, Sch23c, Yoo20, Kir22, Loz22, vdL22]. Polio [Bro21b]. Political [Aug20, Bon23, Den23, GX24, Rog23, Yam23a, Has22, Her23]. Políticas [Mat21]. Politics [Bau20, Bee23, Bix24, Boy21, Fer23b, Fri22b, Gab23, Gar24, Ger21a, Hoc23, Hof21, HH21, Lag24, Lun23, Mar22a, Mar23b, Muk22, Par20a, Pol24, See20a, Shi20, Str22a, Ung22, Wei20, Wel20, Wil23, Zac23, Cos20]. politique [Gou21, Has22]. Pollution [Kob22c, Ste20, Boy21]. Pollutions [Lin22]. Polonia [Lab21]. Ponts [Cha22c]. Popkultur [Mey20]. Popular [Mar22a, Ves20]. Portable [Rua24]. Ports [Cam24]. Portugal [Sou24b, Mei23, Sou24b]. Portuguese [Cam23, Per23b]. posdictadura [Mat21]. Possum [McG24]. Post [Bee23, Ber21b, Col21, KJ23, Kub22, Ves20, Wal22, Mat21]. Post- [Bee23, KJ23]. post-dictatorship [Mat21]. Post-Empire [Col21]. Post-revolutionary [Ber21b]. Post-Soviet [Wal22]. Postal [Kub22].

Postcolonial [AG24, Edw23]. Postsocialist [Yao21]. Postwar [Fra20, Gro23, Kai21, Kat23, Mur20, Sör23, Abe21]. Potter [Cam21a, Tuc22]. pour [Gab22]. pouvoirs [Has22]. Powell [Wel20]. Power [Ade21, Bab22, Bix24, Boy21, DiM21, Gag21, Gil24b, Hen21, Hof21, Ion21, Kap20, Koc21, Kui22, ŁPO^{+}20, Mch24, Mil23a, Miy22, Pre21, Sca20, Sla24, Tat21, Uts21, Wel20, Wij22, Wil22a, Yu21, Zai20, Zel23, Aug21, Bel21, Gra22, Has22, HP21a, Kre21, Coh21, Par20a, See20a]. Powerful [Kui22]. Powering [Brü21, Gag21, Spi24]. Powers [Coa22]. Pozharliev [Bek24]. Practical [Ach22]. Practice [Yoo20, Mar21a, Mul21]. Practices [AH20, Dav23b, DeC22, De 23a, Ger21c, Mak23, Wis22, Sou24b, Ing22, Wie21]. Pradesh [Arn23a]. Pradip [Col21]. Práticas [Sou24b]. Pre [Kan23a, Zyl22]. Pre-colonial [Kan23a]. Pre-Interstate [Zyl22]. Precarious [Yao21]. Precolonial [Map23]. Prediction [vD21]. Prehistory [Gil20, Mil23b]. Premodern [Gun24, Rie22]. Prepaid [VF24]. Preparedness [Har23b]. Prescriptive [Hal21a]. Present [Cru21, Cur23, Rik21, SEO20, Bee21, Cig23, Han20, Pop24, Pre21, Sca23, Zan22]. Preserving [Nol23]. Press [Hva24]. Prestige [Plo23]. Pretel [Pin21]. Prevention [Cur23]. Prieta [Loo22]. Print [Bon21b, Lun23, Par20b, Ver23, Zem21, Gre21]. Printed [Par20b]. Prism [Hva24]. Prison [Hoc23, Lan24]. Privacy [FV20]. Private [FV20, HF21]. Prizes [Str22b]. Problem [AG24, NOK22]. Process [Rob22c]. processing [Aum21]. Prodöhl [KdS24]. Producing [Vik20, d’A20]. Production [Kai23, Rai22, Ree22, Yi22, Gab22, Mic21, Mue22]. Productive [BN23]. Profit [Fru20]. Profits [Rei21b]. Program [Hai23, Kvi23a, CC21]. Programmata [McC23]. Programme [Lip22, ZP24]. Programming [Wal22]. Progress [Ban21, Mon23, Mue22, Ver21, See22]. Progressivism [Ver21]. Project [Hus23, Røs22b, See22, Rag24, Spi21]. Projectors [Mar22a]. Projects [Kif24a]. Projekt [See22]. Projit [Nai22, PW24]. Prometheus [Ten22]. Promise [Lad22, Rus24a, Wu23]. Promote [Esp23]. Promoting [TS20]. Propaganda [Cam21a, Lun23]. Property [Bau20, d’A20]. Prophets [Cor23b]. Propulsion [Wol23]. Protection [Tho23b]. Protests [Tat21, Kin23]. Prototype [Wu23]. Prototypes [Cue22]. Prout [Zac23]. Province [Bia22]. Provocations [Cam21b]. Provokationen [Cam21b]. Prussian [NG21]. Pryor [Zyl20]. Public [Bar21b, Cla21, Fro23, Gid24, Gre21, Gue22b, Lev20, Old24, Par20b, Tom21, Boo24]. Publics [Bot20, Mil20a]. Puissance [Gra22, HP21a]. Punish [Lud23]. Pure [HL22, Pag23]. Purest [Pie22]. Purton [Wal21b]. Pushing [Par20a]. Putting [Hög22]. Pyne [Bes22]. Pyrotechnics [Wor24]. Pyrrhic [Mue22].

Qing [Hom22, Naq22, Zak23]. Quality [Ram23]. Quantification [Bou21]. Quantitative [Gul20, Gul20]. Quest [Gär20a, HHK^{+}20, Kra21b, Rei21b]. Questions [Van20b]. Quiet [Kvi23b]. Quite [Spi20]. Quito [Cas21].

R [Cer20, Coh23, Cop24, DeC22, Gab23, Gag21, GF21, His23a, Ion21, Kre21, Lau23, Low23, Pol24, Rup24, Sto23, Tho23a, Ung22, Zai20]. R&D [Yoo20, Don24]. Race [Ber21b, Bix24, Car21, Jon22b, Lau23, Mcg21, PW24, The22, Zel23, Wil22a, Jon23]. Rachel [Par20a, Rog23, Ste20]. Racial [Bru23, Fou23, Lud23, Mcg21, PW24, SM21, Ves23b, Wen22]. Racing [Tur20]. Racism [BHJI^{+}22]. Racist [Pry23]. Radar [Coh23]. Radiation [Aug21, Aug20, Ren22]. Radio [Ale22, Cam21a, Coh23, HF21, Jon23, Kui22, Mei22a, Rua24, Yan20, Moo20, Rik21, Tho24]. Radio-Internet [Mei22a]. Radio/Black [Jon23]. Radiobilder [Wag23]. Radios [Wag23]. Radiotherapy [Mik20]. Radium [Ham24]. Rae [Sin23]. Rafał [Bar22]. Rafferty [Cur23]. Raggi [Fon23a]. Raianu [Ram22b]. Raiford [Šve21]. Rail [Loo22]. Railroad [Amm22, Sla22, BF24, Bes22, Cli20, Sha22b]. Railroads [Chu21, Tan21a]. Railway [Chu23, Ďur23, Per21a, Per23b, Per23a, Wan21, Cas21, Ras22, Arn23a]. Railways [Mor22, Pik21]. Rain [Ste20]. Rainbow [Tho21]. Raising [Van20b]. Raj [Arn23a]. Ramón [Dav23a]. Rani [Bea20]. rapeseed [Soe21]. Raps [Soe21]. Rare [Ang23]. Rasmus [DeC22]. raw [Ver22a]. Ray [Per21b]. Raymond [LM20, Pre23a]. Rayon [Ald20]. Rayos [Zap20]. RCA [Den20b]. RDX [Ham20]. Reaching [Rai22]. Reactor [Miy22]. Reactors [KJ23]. Reading [Kus23b]. Real [Coa22, Coa24]. Reality [Aks23, Sch21b]. Rebecca [Pry23]. Reception [Pop20]. Recharging [Zen22]. Record [Col20]. Recording [Mun23]. Recycling [Lav21, Emm21]. Red [Don22, HH21]. Reddy [Gil24a]. Rediscovering [Zem21]. Rees [HL22]. Reevaluation [Har21]. reflect [Cam21b]. reflexionszwang [Cam21b]. Reformation [Rob20]. Reft [Pry23]. Regele [Ros20]. Regenerating [Hur20]. Regime [Dup21]. Regimes [Nov23a]. Region [Cat21]. Regional [Ima22]. regulación [Mat24]. Regulating [BE20]. Regulation [Kat23, Mei22b, Pag23, Vet23, Mat24]. Regulations [Ess21, Gou21, Gou21]. Regulator [Pre23a]. Reich [Zel22]. Reid [Pol21]. Reinhard [Mac22b]. Reinvention [Hei20, Pat24]. Related [Ryd20]. Relation [Mil20b]. Relational [De 23b]. Relations [Ber21a]. Relationship [Tho21, Tho22]. Reliable [Roy21, SV21, Tan21b]. Relief [Fut22]. Religion [Lan23a]. Religious [Sto22]. Reluctant [Ver23, Wij22]. Remains [Aka22, Lav21]. Remedies [Ben22]. Remember [Rag24]. Remolding [Ash21]. Renaissance [AD22, Kuc21, Lun23, NG20, Sca23, Rib23, Wac22]. Rendering [Den23]. Renny [Lan23a]. Rentenversicherung [Kra22]. Rentetzi [Ham24]. Renzi [Mil22]. Repair [Bor23]. Repairing [Kre21]. Réparations [Bau23]. Replicators [Fru21]. Representations [Bot20]. Reproduce [Fru21]. Reproduction [Fre22]. Reproductive [Got20, SG20]. Repubblica [How22]. Republic [Aug21, Gun24, How22, Kir22, Kub22, Uts21, Har24]. Republik [Aug21]. Research [Cub23, Dua23, Eng22, GM22, Gro22, Rad23, Wil23, Cam21b]. Reservoir [GX24]. Resistance [GMR24, Jon23, Mor24, Jon22b]. Resnikoff [Rus24a]. Resonance [Pat24]. Resource [HHK^{+}20, Vik20]. Resources [Rei21b, Str22a]. Respawn [Van20a]. Respirators [Chu22b]. Responaut [McG24]. Responded [Edi22]. Response [HH21, Kal20, SC20b]. Responses [Sch20d]. Restoration [Ash21]. Rethinking [Oso22, SG20]. reto [Cam22]. Retooling [Hes20a]. Retzlaff [Pry23]. Reuse [Bor23, Emm21]. Review [Cas24, HG22, OY23, Abe23, Abn23, Ach22, Ada20, Ade21, Aeb23, Aka22, Ald20, Ale23, Als20, Amm22, AG24, And21, And22, Ang23, Arc20, Arc23a, Arc23b, Arn23a, AD22, Aru23, Ash21, Ask22, Ass23, Atw23, Aud22, Aug20, Aug21, Aum21, Bab22, Bac22, Bae23, Bar23, Bar22, BF24, Bar21a, Bas21, Bau23, Bea20, Bec23, Bed22, Bee23, Bee21, Bek24, Bel21, Ben22, Ber22a, Ber21a, Ber21b, Ber21c, Ber21d, Ber24, Ber22b, Bes22, Bia22, Bim22, Bin23, Bin24, Bit21, Bix24, Bla20a, Bla20b, Bla21, Bla23, Boh21, Bol21, Bon21a, Boo22, Boo24, Bor23, Bot21, Bou21, Bou22, Bou23, Boy21, Bre23, Bri20, Bro21a, Bro21b, Brü21, Bru23, Buc23, Bun23, Bur23, But23, CW24, Cai22, Cal22, Cam24, Cam21a, Cam21b, Cam22, Cas21, Cat21, Cer20]. review [Cha21a, Cha22a, Cha22b, Cha21b, Cha20, Cha21c, Che23, Cho22, Chu21, Chu23, Cie21, Cie23, Cig23, CC21, Cla21, Cli20, Coa22, Coa24, Coh23, Coh21, Col21, Col20, Con22, Coo23, Coo20, Cop24, Cor20, Cor23b, Cos20, Cow21, Cox22, Cra21, Cri22, Cro23, Cub23, Cur23, Dac20, Dae21, Dah23, Dal22, Dan21, Das24, Dav20, Dav23a, Dav23b, Daz23, DeC22, De 23a, Deg22, Del21, Den20b, DiM22, Di 21, Die21, Die22, Dim23, Dio24, Dis20, Dis21a, Dis21b, Don23, Don24, Dru20, Dua23, Duh22, Dup21, Duq24, Ďur23, Edw23, Ell22, Emm21, Eng20, Erd21, Ert21, Ess21, Eva21, Fal24, Far24, Fer23a, Fer23b, Fil23, Fin22, Fla23, Fon23a, Fon23b, Fou23, Fox21, Fra20, FR21, Fre20, Fre22, Fri22a, Fri22b, Fru21, Fut22, Gab23, Gab22, Gag21]. review [GA21, Gal22, Gar24, Gär20a, Gar20b, Gat20, Ger21a, Ger21b, Ger21c, Gid24, Gie22, GF21, Gil24a, Gil24b, Gil20, Giu22, Gli23, Gol24, GC24, GB22, Goo21, Gou21, Gra22, Gre21, Gre22, Gro23, Gro21, Gué22a, Gue22b, Gui22, Gum24, Gun24, Gus24, Gut23, Hår20a, HL22, Hah23, Hai21, Hai23, Haj21, Hal21b, Hal20, Ham20, Ham24, Ham21, Han23, Han20, Han22, Har23a, Har21, Har23b, Har24, Has22, HP21a, Hay22, Hei20, Hei23, Hen21, HF21, Her23, Her20, Hes20a, Hes20b, Heß22, Het23, Hir23, His23a, Hoc23, Hof23, Hof21, Hol21, Hol22, Hom22, Hor23, How22, How23, Hub24, Huh24, Hum24, Hun20, Hur20, Hva24, Ign21, Ing22, Ion21, Isr23, Jac21, Jac23a, Jac23b, Jam23, Jan23a, Jan23b, Jon22a, Jon22b, Jon23, Jør21]. review [Jud22, Kag20, Kai21, Kai23, Kal23, Kal20, Kam22, Kan24, Kan20, Kan23a, Kar24, Kar23, Kat23, Kay20, Kay23, Kin23, Kir22, Kli23, Kno22a, Kno20, Kno22b, Koc21, Koc23, Kos21, Kou24, Kra22, Kra23, Kre21, KdS24, Kri21, Kri24, Kru22, Kub22, Kuc21, Kui22, KM23, Kvi23a, Lab21, Lad22, Lag24, Lal21a, Lal21b, Lan23a, Lan23b, Las23, Lau21, Lau23, Lav21, Lee24, LM20, Li22, Lie23, Lif23, Lin24a, Lin21, Lin22, Lin24b, Lip22, Loe24, Loo22, Lóp22, Lóp24, Lou21, Löv23, Low20, Low23, Loz22, Luc20a, Lud23, Lue21, Lun23, Lun22, Mac22a, Mac24, Mac23, Mac22b, Mai23, Mak23, Mal24, MO22, MF20, Man20, Mar21a, Mar21b, Mar22a, Mar24, Mar23a, Mar23b, Mas20, Mat21, Mat24, McC20, McC23, McC21, McD22, McN24, Mcg21]. review [Mcg20, Mch24, Mea21, Mei21, Mei23, Mei22a, Mei22b, Mel23, Mey20, Mic21, Mil20a, Mil22, Mil23a, Mil23b, Mit23, Mit24, Mit22, Miy22, Mod21, Moo21, Moo20, Mor22, Mor23, Mue22, Muk21, Muk22, Muk24, Mul21, Mun20, Mun23, Nai22, Naq22, Nèg22a, Neg22b, Nel22, NG20, NG21, Nis23, Nol23, Nor21, Nor23, Nye20a, Nye20b, O’D20, O’D22, O’S20, Och23, Opp22, OA23, Ott21a, Ove22, Par20a, Par20b, Pat23, Pat24, Pea22, Pel24, Pen22, Per23a, PZ23, Per21b, Pet24a, Pet24b, Pie22, Pik21, Pin21, Plo20, Plo23, Pol24, Pol21, Pom20, Poo23, Pop24, Pos23, Pou23a, Pou24, Pou23b, Pra20, Pra23, Pre21, Pre23a, PW24, PÑ22, Pry23, Pui22, Rad23, RO23, Rai23, Ram22a, Ram21a, Ram22b, Ram22c, Ram21b, Ras22, Rav20, Ree22]. review [Rei21a, Rei21b, Rei21c, Ren22, Rey22, Rib23, Rie22, Rik21, Rin22, Rob22a, Rob20, Rob22c, RR21a, RR21b, Rog21, Rog23, Rom22, Ros20, Ros22a, Le 23, Roy24, Rup24, Rus20, Rus24b, Rus24a, Ryd20, Sad24, San21, Sar22, Sca20, Sca21, Sca23, Sch23a, Sch20b, Sch21b, Sch23c, Sch20d, Sch21c, Sch22, SS21, Sco23, See22, See20a, See20b, Şer24, Sha22a, Sha22b, Sie22, Sim24, Sim20, Sin23, Sla23, Sme20, Smi20, Smi23, Soe21, Sol24b, Sör23, Sou24a, Sou24b, Spi21, Spi24, Sta22, Sta23a, Sta20, Sta23b, Ste20, Ste22a, Šte22b, Sto22, Sto24, Sto23, Str22a, Str22b, SD23, Suá21, Sub22, SM21, Šve21, Swa21, Syl22, SG21, Tan21a, Tap23, Tar23, Tat22, Tay23, Tel20, Ten22, The22, Tho24, Tho21, Tho22, Tho23a, Tho23b]. review [Tom21, Tri20, Tru24, Tsu23, Tuc22, Tur20, Tys24, Ung22, Uts21, Vai23, VÁ21, Van20a, Vee22, Ven24, Ver22a, Ver21, Ver23, Ver22b, Ves20, Ves23b, Vet23, Vir22, Vol22, Wån23, Wac22, Wac23, Wag23, Wal22, Wal21b, War23, Wel21, Wel20, Wen22, Wen24, Whi23, Wie21, Wij22, De 23b, Wil20, Wil22a, Wil22b, Wil21b, Wir23, Wis22, Wit21, Wol23, Wor24, Wri21, Wu23, Yam23a, Yan20, Yan22, Yao21, Yu21, Zac23, Zai20, Zai23, Zak23, Zan22, Zap20, Zdr21, Zel22, Zem21, ZJ22, Zen22, Zha21, ZP24, Zia23, Zul24, Zyl20, Zyl22, dCC22, dZ21, vD21, vE22, vM20, vdL21, vdL22, vdS22]. Revisionist [Rus20]. Revisited [Per21a]. Revolt [Tat21]. Revolution [Bac22, GMR24, Gli23, Loz22, Mar23b, Plo23, Rag24, Ung22, Zen22, Eva21, Kar23, Zha21]. Revolutionary [Bau20, Bot20, KJ23, Ver21, Ber21b]. Revolutions [Per21b]. Rewriting [Joh23]. Reynolds [Lie23]. Rhetoric [KB22b]. Rhythm [Eng22]. Rica [Hum24]. Ricardo [Whi23]. Richard [Arc20, Dav23a, Jac21, Ram22c, Spi24]. Rida [Rob20]. Rider [Whi23]. Riding [BF24]. Rieth [Bar21a]. Rigg [Ves20]. Rigging [Erd21]. Rights [Bur23, Wri21]. Rimmington [Lip22]. Rindzeviciūtė [Bab22]. Rio [Mal24, Mac22a]. Rise [Del21, Ďur23, Hei20, Mac23, Mit22, Pet24b, Fru21, Suá21]. Rising [Nis23]. Risk [Bin24, Chu22b, HH21, Mcg21, MS21, Rob22a, Rom22, Ion21, Ree22]. risks [Gou21]. Risques [Gou21]. Rita [Wij22]. Ritchie [Boo22]. Ritual [Duq24]. River [Ber24, Fre20, Swa21, Li24, Fre20]. Rivers [Gie22, NOK22]. Riviera [Gro23]. Road [Bek24, Cli20, Nor21]. Roads [Gie22, Šte22b, Zel23]. Roaring [Zac23]. roasting [Bee21]. Robb [Das24, Gre21]. Robert [Boo22, Bun23, Nèg22a, Ram21b, Syl22, Cra21]. Roberto [Ber21d, How22]. Robertson [Och23, Wir23]. Robey [Kno22a]. Robo [Och23]. Robot [Rus24b]. Robots [Fru21, Och23]. Robyn [Duq24]. Rocha [Sol24b]. Rocio [Gue22b]. Rodgers [NG20]. Rodrigues [Di 21]. Roebling [Jac21]. Roehrlich [Coo23]. Roger [Rei21a, Tur20]. Rogers [SG21]. Rogue [Pre23a]. Rohstoff [Ver22a]. Roland [Hor23, Mil23a]. Role [Lev20]. Rolland [Ver21]. Roman [Emm21, Tay23, Wal21b]. Romberg [Kir22]. Rome [Hal21b, Lun23]. Romero [Hay22]. Ronan [Bou23, Tru24]. Rosanna [Har23b]. Rosemary [Hol21]. Rosenthal [Sta20]. Ross [Pra20]. Röstkaffeemarktes [Bee21]. Rothschild [Ste20]. Route [Jam23]. Roux [Gou21, Lin22, Tru24]. Roy [MO22]. Royal [Rei21a, Wol23, Esp23]. Rubery [Vir22]. Ruha [The22]. Rules [Jon22a]. Rural [Aud22, Buc23, Cha20, Har24, Kif24b, San20, Sla22, Spi24, Tat21]. Russell [Mit23]. Russia [Hol22, Pol24, Wal22, KJ23, Vee22]. Russian [Pol24, Sei24, Smi20]. Rüther [Bee23]. rx [Hai21]. Ryan [McC20, Pry23]. Ryukyu [Che22].

S [Ber21b, Don24, Ert21, Gum24, Har21, Her20, His23a, Lie23, NOB^{+}20, Tat22, ZP24]. sûreté [Pre23a]. Saara [Pou23b]. Sabine [PZ23]. Safe [Ion21]. Safety [Chu22b]. Saga [SD24]. Sahara [Gué22a]. Sail [Loo22]. Sailing [And21]. Saint [Pre23a, Bea21]. Sakade [Zai23]. Salaam [Deg22]. Salted [Mac22a]. Salvador [Rag24]. Samadhi [Sol24b]. Samantha [Hub24]. Samuel [Rei21a, Opp22]. San [Pen22]. Sandra [Hva24]. Sandwell [Gag21, Rup24]. Sandy [Moo21, Sim24]. Sang [Bae23]. Sanitary [Cro23]. Santamaria [Loz22]. Santesmases [Ign21]. Santiago [Boo24, Cam22, Hva24]. sapiens [Och23]. Sara [Loz22, Mei23]. Sarah [Abe23, Fre22, GF21, Kli23, Kno22a, Mun20, Nor21, Soe21]. Saraiva [Sou24b]. Saravanan [Ram21a]. Sarkar [Cha20, Hai23, Muk24, vdS22]. Sarkis [Spi21]. Satellite [GX24, Lau23]. Saturn [Kes21]. Saumitra [Rai23]. sauvegarde [Pou23a]. Savage [Bru23]. Savanna [Duq24]. savoir [Bau23]. savoir-faire [Bau23]. Savoirs [Gou21, Gab22, dCC22]. Sayer [Mcg20]. Sayf [Rad23]. Scale [TS20, Cha22b]. Scandinavia [Kai21]. Scarcity [Vik20]. Scenes [Wac22]. Schakenbach [Ros20]. Schatzberg [Nye20b]. Scheitern [Aum21]. Schiffer [Coo20]. Schilling [Ryd20]. Schippers [Lin24b]. Schlanger [Sou24a]. Schley [Chu21]. Schmidgen [Pat23]. Schmidt [Rei21c]. Schmidt-Funke [Rei21c]. Schmitt [Kra23]. Scholars [Has21]. Scholarship [Edi20a]. Schönrich [Kub22]. School [And21]. Schools [How23, PW24, But23]. Schotte [And21]. Schulze [Huh24]. Schwalm [PW24]. Schwantes [Rus20]. Schweiz [Aug21]. Science [Ach22, And21, AL20, Bru23, Cam22, Cho22, Dan21, De 23a, Dup21, Eam21, Fri22a, Giu22, Ham24, Hof21, Hor23, HG22, Low20, Luc20a, Lun22, Mat21, Mea21, Mey20, Mil20a, Muk24, Mun23, PW24, Rai23, Rie22, SS21, See20b, Sha22a, Sou24b, SD23, Sub22, Ung22, Dal22, How22, San21, Zul24, Bro21a, Ram21b, Ger21b, Lan23a, Mit23, OY23, RO23, Rei21a]. Science-Fiction [Rie22]. Sciences [Bab22, Cub23, Die21, Hub24, Kvi23a, Nai22]. Scientific [Col20, Has21, Sou24b, Ver23, VÁ21, Isr23]. Scientists [Hof23, Mac22b]. scienza [How22]. Sconce [Hen21]. Scott [Fut22]. Scott-Smith [Fut22]. Screen [SG21, Den20b, Arc20]. Screens [But23]. Sea [Fer24, Jam23, Kam22, Hub24]. Sean [Bre23, vdL22]. Search [Kuo23, Pea22, Vol22]. Seas [Del21]. Sebastian [And22, Ver22a]. Second [Dav20, Fla23, Ram22c, Rik21, See20b]. Secret [Ask22, Bou22, Ham20, Zia23]. Section [Rob22a]. Sections [Got20]. sector [Cal22, Ďur23]. Securing [Hol21, War23]. Security [Aum21, Gol24, HHK^{+}20, Kat23, SC20b, Sla23, Ten22, Mac22b]. Seduced [Ham24]. See [Nor23]. Seeds [Ful24, RCCAM24, Sei24]. Seeing [Fra21, Ott21b, Tuc22, Rog21, vD21]. Sekhsaria [Cho22, Cho22]. Selective [DiM21]. Self [Fru21, Boh21]. Self-Control [Boh21]. Self-Replicators [Fru21]. Semiconductor [SD24]. Senegal [AG24]. Sensations [Tel20]. Sense [Esp23]. Senses [Fer24, Wån23]. Sensibilities [Fer24]. Sensing [CC21]. Seo [Nor21]. Seoul [DiM21]. Separation [Eri21]. Serban [Dah23]. serene [How22]. Serenissima [How22]. Serge [Bel21]. Serial [Tel20]. Series [OY23, Sch20c]. Server [Joh20]. Service [Aum21, Joh20, Kub22, Lóp24]. Services [Cer20, Ert21]. servicio [Lóp24]. Serving [Kay23]. Seth [Pel24, Vet23]. Setting [Jon22a, Wil22b]. Settler [Sei24, Sim24]. settore [Cal22]. Seventeenth [Wac23]. Seventeenth-Century [Wac23]. Sewage [NOK22]. Sewers [Ism23, Wil20]. Sewing [Mai23]. Sex [BN23, Nag21]. Seyer [vdL22]. Sezer [Har24]. Shakespeare [Kri21, Kri21]. Shaky [Gil24a]. Shamsy [Zem21]. Shane [Wil22a]. Shanghai [Ale22, SC20a]. Shaoling [Sco23]. Shape [Bee23, Mil24, VF24]. Shaped [Don22, Jam22, Lut20, Ram23, Roy24, Rus22, Sca20, Sei24, Yan20]. Shapes [Per21b]. Shaping [Dim24, Kir23, Wie21, Sub22]. Sharra [Dru20]. Sheet [Pop20]. Shelley [Cli20, Hva24]. Shelton [See20a]. Shenzhen [Ste22a]. Shepherd [De 23a]. Shiji [Mor23]. Ship [Pol21]. Shipbuilding [Pou23b]. Shock [Dru20]. Shocks [Hög22]. Shop [Yi22]. Short [Rin22, vD21]. Show [McC22, Jon23]. Showcasing [Kif24b]. Shredding [Mac23]. Shroud [Con22]. Shukla [Kan23a]. Shulman [Dav20]. Sibylle [Aug21]. Sichtbarkeit [MF20]. Sicilia [Dav23b]. Sidney [Rus20]. Siebeneichner [Ell22]. Siebert [Naq22]. siècle [Bas21, Bel21, Cha22b, Cha21c, Gab22, Gou21, Gui22, Nèg22a, RR21b, Ver22b]. siècles [Wil21b]. siglos [Boo24]. Signal [Dac20]. Signals [Kan20]. Silent [Tel20]. Siles [Hum24, PÑ22]. Silesia [ŁPO^{+}20]. Silk [Ald20, Ber21e, Muk21]. Silva [Het23, Lóp22, Lóp24]. Silver [Gue22b]. Silvia [Dal22, Pat24, Wu23]. Simon [Bla20a, Cam21a, Mey20, Mic21, Ass23, Boo24]. Simone [Lun22]. Simonsen [DeC22]. Sims [Kre21]. Simulmatics [Mar21b]. since [AH20, Cam23, Die22, Jon22a, Kai23]. Sing [Lin24a]. Singh [Pet24a]. singular [Bel21]. singulière [Bel21]. Siry [Cha21a]. Sites [Hag21]. Sittler [MF20]. Situating [Ale22]. Six [HG22]. Sixteenth [Dav20, Hal21b, Wal21b]. Sixteenth-Century [Hal21b]. Sixty [Edi20a]. Skidmore [Ang23]. Skies [Ste20, vdL22]. Skill [And21, Gun24]. Skin [Gar20b, Kan23b]. Skins [PW24]. Sky [Lee24]. Slavery [PW24, Sta20]. Slotten [Lau23]. Slovakia [Šte22b]. Slovensku [Šte22b]. Small [Nai22, TS20, Mat24]. Small-scale [TS20]. Smil [Cow21]. Smith [Fra20, Fut22, Ger21c, Hei23]. Smritikumar [Cha20]. Snapshots [Kru20]. Snatcher [Hai21]. sobre [Boo24, Sol24b]. Social [Abn23, Die21, Dru20, Hus23, Luc20a, Mac24, Mil20a, Mil23b, Per21b, Sch21a, Sch21c, Sin23, Tri20, Ven24, Zan22, Mic21]. sociales [Die21]. Socialism [Bek24, Fox21]. Socialist [Bek24, Gat20, Ign22, Zou22, Eng20]. Socializing [Ott22]. Society [Ano21b, Ano22b, Ano23a, Cho22, Cig23, Cor23b, Dim23, Dio24, GMR24, Dis20, Eng20, Sch20b, Sch23a, Zac23]. Socio [Wit22]. Socio-Technological [Wit22]. Socioeconomic [Ram22c]. Sociology [Mit23]. Sociotechnical [Har23b]. Soffer [Sin23]. Software [Bur23, McC22, SM21, Bur23]. Soho [Fer23a]. sol [Lóp24]. solar [Lóp24]. Sometimes [KdS24]. Sonic [Boh21, Dac20]. Sonnic [Mar24]. Soon [Lue21]. Sophie [Ach22, Duh22, RR21b]. Soul [Les24, Mol22]. Sound [Bit21, Kan23b, Kvi23b, Kvi23a, Yam23b, Yam23a, Hor23, Mei22a]. Soundings [Moo20]. Sounds [Mon22]. Soundtracked [Lin24a]. Sources [SC20a]. South [Har21, Kif24a, Moo20, Mor24, Ram22a, Ram21b, Ung22, Arn23b, Car21, Kan23b, San20, Yoo20, Hur20]. Southwest [Gie22, Mun20]. Southwestern [Cat21]. Sovereign [vdL22]. Soviet [Coh23, Gli23, Lip22, Smi20, Vee22, Wal22, Chu22b, Hil23a, KJ23]. Soy [Het23]. Soybean [Het23, KdS24]. soziale [MF20]. Sozialstaat [Kra22]. Space [Aeb23, Bim22, Chu21, Cub23, Ell22, Han23, Lau21, Lau23, RO23, Ras22, SG20, Vol22, ZP24, Sca21, Fin22]. Spaceflight [Ros22a]. Spaces [Wen20]. Spain [Ign21, Lab21, Lóp22, Mai23, RR21b, VÁ21, Hof21]. Spalding [Per23a]. Spanish [Bea20, Cas24]. Spanish-Language [Cas24]. Sparkassen [Kra23]. Sparked [Brü21]. Sparks [Sch23c]. Spatial [Gli23]. Speaker [Kvi23b]. Spectacle [Kri24]. Spectacular [Coo20]. Spectatorship [Kri21]. Specula [Nov23b]. Speculations [Nov23b]. Speech [Bre23]. Sphere [Hes20a]. Spinner [Ber21e]. Spinning [Ram23, Yi22]. Spires [Han23]. Spirited [d’A20]. Splendor [Mea21]. Springs [Mor24]. Sputnik [Lau23]. Squares [Hof23]. St [Lin24a]. Staatssicherheit [Mac22b]. Stafford [Jam23]. Stage [Aeb23]. Staging [Wor24]. Stamping [Hil23a]. Standard [Jon22a]. Standards [Mil24, Wil22b]. Starosielski [Arc23a]. Stasi [Aug21]. State [Als22, Aum21, Gat20, Gid24, Hei23, Ign22, Kui22, Lif23, Loz22, Mar22a, Ros20, Ryd20, Sch21a, Yu21, Zul24, dCC22, Kra22, Mac22b, Kno22a]. State-Related [Ryd20]. State-Socialist [Gat20, Ign22]. Statecraft [Spi20]. States [Cat21, Koc23, VÁ21, Jam22, RW23, Sla22, Tat21, Wal21a, Aru23, Ess21, Lie23]. Statgesellschaft [Eng20]. Statistics [HG24, Bou21]. Statues [Wac23]. Staudammbau [Huh24]. Steadman [Wac22]. Stealth [Zia23]. Steam [Cro23, Jam23, Mil20b, Chu21]. Steams [Mil20b]. Steamship [Bec23]. Steel [Kob22a, Kob22c, Sar22]. Steelmaking [JR23]. Steenson [Hal20]. Stefan [Bor23, Dah23]. Stefano [Mor22]. Stein [Lan23b]. Steiner [De 23a]. Stelu [Dah23]. Stepansky [Mul21]. Stéphane [Cha21c]. Stephanie [Sch23a]. Stephen [Gro23, Kli23, Wil20, Wri21]. Steppe [Sei24].

Steps [Mod21]. Stereographs [Coa24]. Sterilization [KM23]. Steven [De 23a, Luc20a, Per23a]. Stewart [Sla23]. Stiller [Mal24]. Stoffgeschichte [Soe21, Ver22a]. Stomach [Fut22]. Stone [Sco23]. Stordeur [Zyl20]. storia [How22, San21, Sca21, Ber22b]. Stories [His23a, Kag23, San20]. Story [Bou22, Jon22b, Ste22a, Vir22]. Storytelling [Ano21c]. Straeten [Kal23]. Strahlen [Aug21]. Strands [Dav23b]. Straße [MF20]. Straßenverkehr [MF20]. Strasser [Ber21c]. Strategic [Bla20b, Eam21]. Strathman [Ott21a]. Streamliner [LM20]. Streams [Mei22a]. Street [Bot20]. Streets [Muk22]. String [Mar23b]. Striphas [Bin23]. Strom [Aug21]. Struggle [Boy21, Dis21b, Wij22]. students [Aud22]. Studies [Low20, Luc20a, Yam23b, Boo24, Bit21]. Studio [Sha22a]. Study [Bea21, Kal23, Map23]. Stuff [Man20]. Stupid [Hes20b]. Suárez [Lóp22]. Subaltern [GL23]. Submarine [Hun20]. Submarines [Bin24]. Subterranean [Duq24, d’A20]. Suburban [Mit24]. Suburbs [Rib23]. Subway [Gal22]. Suez [Far24]. Sufi [Gün21]. suisse [Pou24]. Sulphuric [Bol21]. Sumner [dZ21]. Sun [Nis23, Lóp24]. Super [Ham20, Low23, Hai21]. Super-Explosive [Ham20]. Super-Express [Low23]. Supermarket [Wil22a]. Superrealistas. [Zap20]. Supp [Sto22]. Supp-Montgomerie [Sto22]. Supplement [Ano20f]. Supplying [Eng20]. Support [Bim22, Les20]. Supremacy [Mor24]. surco [Zul24]. Surgical [Mur20, Zou22]. Surveillance [Aum21, Sch21c, Jan23a]. Survival [DiM22, Har23b, Lan23b]. Survive [MK23]. Susan [Arc23a, Hes20b, Isr23, Kan20, Tho23a]. Suspending [Sch20c]. Sustainability [Tri20, Lin21]. Sustainable [Boy21, Die22, Jør21, SEO20]. Sustaining [Zak23]. Suvobrata [Muk24, vdS22]. Svelch [vM20]. Svik [Bed22]. Swalwell [McD22]. Swansea [Ber22a]. Swartz [Arc23b]. Sweden [Sör23, Uts21]. Swedish [BE20, JR23]. Sweetman [War23]. Swirl [Vai23]. Switched [Kar23]. Switzerland [Aug21]. Swords [Hai23, Lip22]. Sylvest [Har23b]. Sylvia [NG20]. Symbolic [Dac20]. Symposium [BBC^{+}21]. Synthesizer [Kar23]. System [Bau20, RW23, Sla23, GMR24, Sha22a]. système [Pou24]. Systemic [BHJI^{+}22]. Systems [Bab22, Bim22, Dav20, Kaa23, McC22, Tom21]. Szmytka [Bar22].

T [Bea20, Bla21, Ell22, Rey22]. Table [Ree22, Mac22a]. Tables [Mor23]. Tact [vE22]. Tage [Bee21]. Tailors [Val21]. Taiwan [Wan21]. Take [Edi20b, Gul20]. Takes [Mel23]. Takeshi [Zai23]. Taking [Moo23, Pre21, See20b]. Tale [Dav20, Wit22]. Talking [Vir22]. Tamed [Ten22]. Tampon [Røs22b]. Tan [Zen22]. Tank [Han20, Hub24]. Tanzania [Kif24a, Edw23]. Tapio [Ert21]. targets [Mac22b]. Tarr [vdL21]. Tarter [Sim20]. Taste [Gol24, Kag20]. Tastes [His23a]. Tata [Ram22b]. Taylor [Fru21, Cri22]. Tea [Lou21]. Teaching [Cai22]. Teams [Sub22]. Teasley [Abe23]. Tech [Lad22, NG21, San20]. Technical [Cha22b, Fox21, Gün21, Hen21, Lin24b, Mil24, dCC22, Mor23, Whi23]. Technici [Lin24b]. Technicians [Lin24b, Sha22a]. technicien [Fon23b].

Technicity [Kri21]. Technics [Mar21a]. Technik [Hår20a, Rie22]. technikgeschichte [Cam21b]. Technique [Dav23a, Has22, Cha22b, Nèg22a]. Techniques [Die21, HP21b, Bas21, Bau23, Duh22, dCC22]. Technisch [Lin24b]. Technische [Fox21]. Techno [Eri21, Nel22, San20, Rag24]. Techno-Diplomacy [Nel22]. Techno-moral [San20]. Techno-Utopian [Rag24]. Technocracy [Pre21]. Technocratic [Dim23]. Technocrats [McC20]. Technogenic [Van20a]. Technological [Ale22, Cha21b, El 21, KMÅ20, Kus23a, Lev20, Mch24, Mon23, Per23b, Pra20, Rey22, Roy21, Rus22, Ves23a, Wel21, Wit22, Pea22]. technologie [Sou24a]. Technologies [Abe21, Aeb23, Arn23b, Ban21, Coo20, Dac20, Fra21, Ful24, GC24, Gro21, Gué22a, Ign22, Kri21, Kru20, Mik20, Nai22, Nov23a, Rof20, SG20, Sch22, ZJ22, Mei21]. technologists [And22]. Technology [Ano21b, Ano22b, Ano23a, AL20, Bas21, Bec23, Bel23, BBC^{+}21, Ber21b, Ber21e, BHJI^{+}22, Bor23, Cam21b, Cam23, Car21, Cat21, Che22, Cho22, CC21, Cop24, Cos20, Cru21, Dan21, Dav23b, De 23a, Den23, Dim23, Eam21, Edi20b, Fil23, Fro23, Gar24, Ger21b, Gil20, GL23, Hag21, HP20, Has22, Hes20a, Hes20b, HHK^{+}20, Hil23b, HG22, Isr23, Kaa23, Kag23, Kan23a, KB22a, Kes21, Kra21a, Kri24, Kuo23, Lev21, Li22, Li24, Lóp22, ŁPO^{+}20, Low20, Luc20a, Map23, McC20, Mil22, Mol22, Muk24, Neg22b, Nor20, Oso22, Per21b, RO23, Ram23, Rei21a, Ren22, RCCAM24, Rie22, Rob22a, Rog23, Røs22b, Sca23, SV21, Sch23c, Sch20d, See20b, Şer24, Sha22a, Sta23b, SD23, Swa21, Tat21, The22, TT22, Twa20a, Twa20b, Van20b, Vet23, Wen20, Wen24]. Technology [WAB^{+}23, Yoo20, Zac23, ZK22, How22, Mat21, Nèg22a, RR21a, Sol24b, Sou24a, Kru22, Nye20a, Opp22, Pat24, Pol21, SS21, Wån23, Yan22, vE22, Ale23, Bea20, Cha20, Eva21, Gat20, Kal23, McN24, Nye20b, Pin21, Pop24]. Technology-in-Operation [Hil23b]. Technopolitics [Bin24]. Technoscience [Rib23, Rin20, The22]. Technosciences [AL20]. Technosphere [Ott22]. Técnica [Lóp22, How22, Sol24b]. tecnología [Mat21]. Ted [Bin23, Loz22]. Teenagers [vM20]. Telecommunication [Nel22]. Telecommunications [Col21, Kay23, War20]. telefónico [Mat24]. Telegraph [Hes20b, HH21, Kir23, Nel22, Rus20, RW23, Sto22]. Telegraphs [Rus22]. Telegraphy [Giu22, Hun20]. Telehealth [Sta23b]. Telephone [AH20, Ber23, Bon21b, Yan20, Mat24]. Telescope [Rav20]. Television [Koe23, Mur20, Rog21, Kan20, Atw23]. Tell [Hol21]. Temples [Ram21b]. Temporality [Kal23]. temptation [Lin24b]. Tepa [Tel20]. Teresa [Dal22, Lab21]. Terms [El 21]. Terraferma [Bia22]. Territories [Mun20]. Terrors [Tho23a]. Terry [Lie23]. Testing [Her23, Goo21]. text [Dua23]. Textile [Che22, Ven24]. Textual [Aka22]. Thaggert [BF24]. Thatcherist [War20]. Theater [Kri24]. Theatre [Aeb23, Kri21, Wor24]. Their [Cha22a, Tap23]. Theory [Le 23, Bla20b, Le 23]. Therapy [Ren22]. There [Sta23b, Ves23a, vdS22, Tho20]. Thermodynamics [Ger21a]. Therrien [Hai21]. Things [Che23, Ove22, Rob20, Twa20b, Vee22]. Think [Han20]. Thinking [Kli23, Pet24a, Sol24b, Tuc22, Wen22, Nèg22a, Kvi23a]. Third [Zan22, Zel22]. Third-Front [Zan22]. Thirsty [Dav20]. Thomas [Cie21, Gou21, Hår20a, Has22, Heß22, Ion21, Kra22, Lie23, Lin22, Mit22, Nor23, Sca23, Tat22, Col21, Lan23a]. Thor [Dac20]. Thought [Les20]. Thousand [Ale23, Bar21b, Bla21]. Thousand-Year [Ale23]. Threatens [Cam24]. Threats [Wel21, Wil21b]. Three [Aug20, Dav20, How22]. Tiago [Sou24b]. Tianjin [Cie23]. Tiina [Sme20]. Tilman [Mac24]. Tilmann [Ell22]. Tim [Fru21]. Timber [Loo22, Zak23]. Time [Bra20, Ful24, Gro22, McC21, SG20, Wit21]. Timekeepers [Gün21]. Timekeeping [McC21]. Timelines [Sme20]. Times [Jac21, Jon23, Wal22]. Timo [Sme20]. Timothy [SD23]. Tirthankar [MO22]. Title [Ano20c, Ano21d]. Tkaczyk [Goo21, Kvi23a]. today [Jud22]. Todt [Zel22]. Toister [Con22]. Tokugawa [Wit21]. Tokyo [NOK22]. Toledo [Dav20]. Tom [Boy21, Cor20, Fut22, Jon23, Tom21]. Tomás [Ďur23]. Tomato [O’S20]. Tombs [GX24]. Tomorrow [Den20b]. Ton [Ves20]. Tool [Eng22, Ott21b, Gab22]. Topography [Hal21b]. Toribio [Di 21]. Torpedo [Jam22]. Torsten [Cat21]. Toshihiro [Her23]. Total [Tho23a]. totalitaire [Lin24b]. totalitarian [Lin24b]. Touch [Plo20]. Tourism [Gro23]. Tousignant [AG24]. Tovar [Sol24b]. Town [Sar22]. Toxic [Cam24, Dru20]. Toxicology [AG24]. Tóxicos [Bar23]. Tracing [SD24]. Track [Sla22]. Tracking [Coh23]. Tracks [Sha22b, Wan21]. Trade [Hun20]. Tradition [Bea20, Zem21]. Traditions [Nai22]. Tragedy [Amm22]. Train [Gal22, Rus20, Low23]. Training [Kif24b, Yam23c]. Trains [Per23a]. Trajectories [Chu22b]. Trans [Ras22, Mei21]. Trans-Iranian [Ras22]. Transatlantic [DeC22]. Transcontinental [Cli20, Sha22b]. Transfer [Ber21e, TT22]. Transferred [Mik20]. Transformation [Aug20, Dah23, Ram22a, Roy21, Tan21a, Wen24]. Transformations [Bae23]. Transformed [Ash21, Bur23, Cor23b, Dae21, Ham24, Nor21, Zem21]. Transforming [Abe21, Hor23]. Transfusion [Ber21a]. Transhumanism [Ram21b]. transición [Cam22]. Transit [Pos23]. transition [Cam22, Sta22]. Transitions [Mil20a]. Translated [Gro21]. Transmissions [Mil22]. Transnational [Cha22c, Nel22, Nis23, PÑ22, SS21, Sto23, Per23a, Ryd20]. Transpacific [Tsu23]. transparence [Pre23a]. Transparency [Fal24]. Transparent [Fer23b]. Transport [Bek24, Gie22, Nor20, Boo24]. Transportation [Han22, See20a, vdL21, Bae23]. transporte [Boo24]. Travail [Die21, Gui22]. Travelers [Zyl20]. Traveling [Jon22b]. Treasure [Ove22]. Treatises [Mor23]. Trecento [How22]. Trial [Rob20]. Triangle [Rad23]. tribute [Nèg22a]. Trip [Miy22]. Tropicalizing [Rua24]. Troubled [Kou24, Ove22, Ber24]. trust [Mac22b]. Trusting [SC20b]. Truth [Pry23, Sch20c]. Tsarist [Hol22]. Tschernobyl [Gut23]. Tsuchiya [RO23]. Tubs [Koc23]. Tucumán [Dim24]. turbines [Gra22]. turbo [Hai21]. Turchetti [Lun22]. Turin [Con22]. Turing [Con22]. Türkei [Aud22]. Turkey [Aud22, Har24, TT22]. Turkish [Har24]. Turmoil [SD24]. Turn [Die22, Ren22, Sör23]. Turner [Hir23]. Turret [Plo23]. TV [Mac24, Sch20c]. TVs [Den20b]. Twagira [Gol24]. Twenties [Zac23]. Twentieth [Bae23, Bas21, Bru23, Cha21b, Cub23, Dae21, Gol24, Ham21, Hof21, Lal21a, Low20, Sta23a, Boo24, Huh24]. Twentieth-Century [Bae23, Bas21, Bru23, Cha21b, Gol24, Hof21]. Twenty [HG22]. Twenty-Six [HG22]. Twitter [Hes20b]. Two [Cor23a, Fut22, Wit22]. Tworek [Gum24]. Typewriter [Ing22]. Tyranny [Ove22].

U [Die22]. U-Turn [Die22]. U.S [Gil24b, PW24]. U.S. [Bar21b, Cal21, Cor23a, Rua24, Rus22]. Überwachung [Aum21]. Uhl [Pop24]. Ukrainian [Wen20]. Ulf [Kri24]. Uncertainty [Ren22]. Uncomputable [Gab23]. Underground [Wil20, Ask22]. Understanding [MS21]. Undone [Mil20a]. Unearthing [Fro23]. Unending [Zha21]. Ungendering [Hes20a]. Unidos [VÁ21]. Unilever [Røs22b]. Unintended [Ung22]. Union [Lip22, Nel22, Coh23, Gli23]. United [Aru23, Cat21, Ess21, Koc23, Lie23, VÁ21, Jam22, RW23, Sla22, Tat21, Wal21a]. unity [Mic21]. Universe [Sol24a]. University [Fox21, ŁPO^{+}20]. Unmanning [Sch21c]. Unnatural [Str22a]. unsere [Bee21]. unserer [Dis20]. Untold [Bou22, Vir22]. Upcycled [Jør21]. Upscaling [Kra21b]. Urban [Chu21, Die22, Gre21, Har20b, Nor20, Pry23, SEO20, Zyl22, Eng20, Sin23]. Urdu [Gre21, Tho24]. USA [Wil22a, Mey20]. Use [Ber23, Bon23, Cha22a, Sch20b]. Useful [Ach22, Bro22, Roy21, SV21, Sch21a, Tan21b]. USENET [Zea22]. User [Dae24, Hum24, Mar22b, Fer23b]. User-Friendliness [Fer23b]. User-Inventor [Dae24]. Users [Kaa23]. Using [Has21, Tay23]. Utility [BE20, Rei21a]. utopia [Aum21]. Utopian [Rag24]. Utopias [Bol21, Ger21b]. Utopie [Aum21]. Uttar [Arn23a].

V [Dis21b, Šte22b]. Vânia [Mei23]. Vaccination [Yu21]. Vaclav [Cow21]. Vaginal [Nov23b]. Valencian [GF21]. Valentine [Fla23]. Valeria [Pet24a]. Valérie [RR21a]. Values [Gil24b]. Valuing [Tho23b]. vanguardias [Zap20]. Vanhaelen [Wac23]. Vanished [ZJ22]. Vargha [Bro21b]. Variation [Ima22]. Variety [O’S20]. Vaupel [Jan23b, Rom22]. Veins [Gue22b]. Velayutham [Ram21a]. Velocipede [Pel24]. Velocipedomania [Pel24]. Venezia [How22]. Venice [How22, Bon23]. Venkat [Pra23]. vent [Gra22]. Vergara [Wen24]. Vérin [Nèg22a]. verleiding [Lin24b]. Vernacular [Don22, McD22, Gro21]. vers [Pre23a]. Versagtes [Mac22b]. Versailles [Bos20]. Versorgung [Eng20]. versus [Plo23]. Vertesi [Sub22]. Vertical [Wir23]. Vertrauen [Mac22b]. Very [Rin22]. Veterinarians [Den23]. Vichy [Die21]. Victor [Kan24, Mat24]. Victoria [But23]. Victorian [Ada20, Cro23, Giu22, Har23a]. Video [Koc23, O’D22, Šve21]. Videocassettes [Ask22]. VIII [Lóp22]. Viktor [Gat20]. Viktoria [Goo21, Kvi23a]. Village [Ber21d, Har24]. Vincanne [Vai23]. Vincent [Wal22]. Vinsel [Ess21]. Violations [Ess21]. Virtual [Gro22]. Virtuale [Mac24]. Virtually [Nol23]. Viruses [Erd21]. Viscera [Gar20b]. Viscose [Ald20]. Visibility [Buc23]. Visible [Mei22b, Wei20]. Visier [Mac22b].

Vision [Tuc22]. visionarie [San21]. visionary [San21]. Visions [Aks23, Cor23b, Dim23, Fru21]. Visual [Don23, Gre22, Jac23b, KB22b, Per23a, Rob22b, Yam23b]. Visualizing [Kag20, Miy22]. Vitale [Rib23]. Vogel [Ryd20]. Voices [Yan20]. Vol [Ano22a, Ano22c, Low20]. ’volkse [Lin24b]. Volume [Ano20a, Ano20c, Ano20e, Ano20d, Ano20g, Ano20h, Ano20f, Ano21a, Ano21d, Ano21f, Ano21e, Ano21g, Ano21h, Ano22e, Ano22d, Ano22f, Ano22g, Ano23c, Ano23b, Ano23d, Ano23e, Ano24b, Ano24a, Cam23, Mic21]. Volumes [HG22]. Vom [NG21, Aum21]. Vormoderne [Rie22]. Vostral [Dru20]. Vote [Erd21]. Vox [Wel21]. Voyager [Kes21]. Vrignon [HP21a]. vs [BE20]. Vulnerable [Sla23].

W [Gag21, Han22, Kri21, Las23, Rup24, SD23, Tys24]. Wait [Gal22]. Waiting [Kal20]. Wake [Kno20]. Walking [Mor24]. Wall [LM20]. Walls [Cue22]. Walt [BMSS24]. Waltenberger [Soe21]. Walton [Luc20a]. Wang [Bec23]. Waqar [Pea22]. War [Aug21, Bab22, Bin24, Bro21b, Cri22, Ger21b, Gil24b, Ham20, Han20, Har23b, Jam22, Kan24, Kra21b, Kui22, Lau21, Lou21, Mit24, Mul21, Pou23b, PW24, RO23, Rav20, Rey22, Rib23, Ros20, Røs22b, Rua24, Tys24, VÁ21, Wil22a, Zai20, Zen22, dZ21, Abe21, Ell22, Bed22, Fla23, Gär20a, Hai23, See20b, SD23, Tho23a, Zyl20]. Ware [Sta22]. Warfare [Sch21c, Roy24]. Waring [Wri21]. Warm [Sta22]. Warmes [Sta22]. warming [VÁ21]. Wars [Cri22, dCC22]. Was [Kif24a, Sto22]. Wasn’t [Sta23b]. Wasser [Sta22]. Waste [Hay22, Hol21, Kra21b]. watch [Jud22]. Watches [Kag23, Jud22]. Watchful [But23]. Watching [Edi20b]. Water [Aka22, Dav20, Di 21, Dim24, Ert21, Gue22b, Hög22, Ism23, Kam22, RR21b, Wei20, Bel21, Sta22, Dis21b, Mal24, Ram21a]. Waterfall [Jac23a]. Waters [Bia22, Dis21b, Ber24]. Waterworks [Cro23, BMSS24]. Watery [Tay23]. Watson [Kay23]. Watt [Fer23a]. Watters [Cai22]. Wax [Rob22c]. Waxworks [Wac23]. Wayne [Lue21]. Ways [Bla21, Nèg22a]. Wealth [Dim24, Ign21]. Weaponization [Las23]. Weapons [Gus24, Her23, KJ23, Roy24]. Weather [Sim24]. Weaving [MO22, Ram22a]. Weber [Bor23, Hår20a, Cam21b]. Wege [Sch20b]. Wei [Coa22]. Weiße [Sta22]. welfare [Kra22]. Well [Zha20, Tri20]. Well-Being [Tri20]. Wellock [Ion21]. Wells [Zha20]. Welt [Sch21b]. Weltkriege [Jan23b]. Wendt [Pin21]. Went [Erd21, Kra22]. West [Duq24, Fer23a, Mal24, Sch20b, See22, Tho21, d’A20, Sch23b, Pou23b]. Western [Coh23, Har23b, McC21, Mul21, PZ23, Uts21, Bec23]. Westwick [Zia23]. Wharton [Tel20]. Wheeler [Cor20]. Where [Fer23a, Sch20a]. Whidden [Pel24]. While [Mor24]. White [Mor24, Pag23, Pos23, PW24, Sta22, Erd21]. Whitman [BMSS24]. Who [McG24, Smi23, Sta23b, Twa20b, San21]. Widespread [Ima22]. Wie [Kra22, Sch21b]. Wielenga [Ram22a]. Wiggins [Mcg21]. Wilder [Col20]. Wiley [HL22]. William [Pom20, Rey22]. Williamson [GB22, Zdr21]. Willoughby [PW24]. Wilson [Cas21, Bix24, Emm21]. Win [Ham20]. Wind [Kap20, Gra22]. windmills [Gra22]. Winfried [Fox21]. Wings [Nis23]. Winick [Che23]. Wintersteen [Bac22]. Wire [Cal21]. Wired [Kay23]. Wireless [Sco23, Cam21a]. Wires [Cat21]. Wiretapping [Aru23]. Wirklichkeit [Sch21b]. Wisdom [Dim24]. Wissen [Huh24]. Wissenschaftler [Mac22b]. Wissensgeschichte [Huh24]. without [Nye21]. Witnessing [Bon21a]. Wittje [Hor23]. Wittner [Dav23b]. Wolf [Rob22a]. Wolfgang [Fox21]. Wolfram [Mea21]. Womack [Fou23, Ren22]. Women [Ang23, BF24, Ger21b, Hes20a, KJ23, Lau21, Rup24, Sch23b, Smi23, Tap23, Tur20, Twa20b, Wån23, Lab21, Lee24, Tur20]. Womenspace [Zea22]. Wong [Mel23]. Words [Tap23, Gil24b]. Work [Atw23, Bon21a, Cer20, Nag21, Naq22, Sar22, Sch21a, Tys24, Die21, Gui22, Nèg22a, Ger21a, Rus24a]. Workarounds [Kob22a]. Worker [Poo23]. Workers [Kay23, Sha22b, Smi23]. Working [Cie21, GF21, KB22b, Mak23]. Working-Class [Cie21]. Works [BMSS24]. workshop [Die21]. World [Ale23, BMSS24, Cha22c, Cie23, Col20, Cri22, Fla23, Gum24, Ham20, Hol21, Jac23a, Kay23, Low20, Low23, Mei21, Mil23b, Nye21, Pea22, See20b, Spi21, Tay23, dZ21, Sch21b, VÁ21, Kra21b, Bab22, Kal20, Koc21, Sin23, Wen22]. Worlds [She20]. Worthen [Kri21]. Wretched [Kru22]. Wright [Hal20]. Writers [Wån23]. Writing [Dac20, Ing22, Kra21a, SS21, Tuc22]. Wrought [Hal21a]. wurde [Kra22].

X [Zap20]. Xi [ZP24]. Xiao [Yao21]. XII [Gui22]. Ximo [Bar23]. XIX [Boo24, Ver22b, Wil21b]. XIXe [Gab22]. XVe [Cha21c]. XVI [Gui22, Nèg22a]. XVIIe [Gou21]. XVIII [Bel21, RR21b]. XVIIIe [Cha21c]. XX [Bel21, Boo24]. XXe [Bas21, Cha22b, Gab22, Gou21]. XXI [Nèg22a, Ver22b, Wil21b].

Yagou [McN24]. Yaman [Lag24]. Yanai [Con22]. Yann [Bia22, Wis22]. Yasar [Yan20]. Yates [Jon22a]. Year [Ale23]. Years [Edi20a, Fut22, Luc20a, How22]. Yemen [Aka22]. Ying [Zen22]. York [Gal22, Wil20]. Yost [Cer20]. Yugoslavia [Bek24]. Yuji [Aeb23]. Yuka [RO23]. Yulia [Wit21]. Yuriko [Tsu23]. Yves [Jud22].

Z [O’D22]. Zac [Bre23]. Zacatecas [Gue22b]. Zaidi [Pea22]. Zajácz [Wij22]. Zakharova [Bas21]. Zalc [Gul20]. Zallen [Bri20]. Zarenreich [Hol22]. Zaretsky [Aug20]. Zedong [ZP24]. Zeide [His23a]. Zeitalter [Jan23b]. Zeller [Nor23]. Zengjian [Mei21]. Zeynep [Mar21a]. Zhang [Zak23]. Zieger [Arc23a]. Zimmerman [Sin23]. Zimring [Jør21]. Zinc [Hom22]. Zoglauer [Hår20a]. Zorina [Str22b]. Zukunft [Kub22, See22]. Zum [Cam21b, NG21]. Zunino [Pet24a]. Zur [Sch21b]. Zwierlein [Ten22].

References

Abel:2021:TCW

[Abe21] Jessamyn R. Abel. Technologies of Cold War diplomacy: Transforming Postwar Japan. Technology and Culture, 62(1):128–155, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
783908.

Abel:2023:DMJ

[Abe23] Jessamyn R. Abel. Designing Modern Japan by Sarah Teasley (review). Technology and Culture, 64(3):1014–1016, July 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/904025.

Abnet:2023:ABS

[Abn23] Dustin Abnet. Arcade Britannia: A Social History of the British Amusement Arcade by Alan Meades (review). Technology and Culture, 64(3):1002–1003, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/904018.

Achim:2022:SUN

[Ach22] Miruna Achim. The Science of Useful Nature in Central America: Landscapes, Networks and Practical Enlightenment, 1784– 1838 by Sophie Brockmann (review). Technology and Culture, 63 (2):570–572, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/854031.

Arribet-Deroin:2022:CAM

[AD22] Danielle Arribet-Deroin. Le Canon au Moyen Age et à La Renaissance, 1338–1559 [The cannon from the Middle Ages to the Renaissance, 1338–1559] by Emmanuel de Crouy-Chanel (review). Technology and Culture, 63(2):565–566, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/854028.

Adams:2020:CIB

[Ada20] Annmarie Adams. Contagion, Isolation, and Biopolitics in Victorian London by Matthew L. Newsom Kerr (review). Technology and Culture, 61(1):347–349, January 2020. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/752947.

Adebayo:2021:AJN

[Ade21] Damilola Adebayo. Atomic Junction: Nuclear Power in Africa after Independence by Abena Dove Osseo-Asare (review). Technology and Culture, 62(2):586–587, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/794409.

Aebischer:2023:PSS

[Aeb23] Pascale Aebischer. Performance Space and Stage Technologies: A Comparative Perspective on Theatre History ed. by Nawata Yuji and Hans Joachim Dethlefs (review). Technology and Culture, 64 (2):588–589, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893054.

Amoako-Gyampah:2024:EET

[AG24] Akwasi Kwarteng Amoako-Gyampah. Edges of Exposure: Toxicology and the Problem of Capacity in Postcolonial Senegal by Noémi Tousignant (review). Technology and Culture, 65(2):716–718, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
926341.

Abildgaard:2020:LNT

[AH20] Mette Simonsen Abildgaard and Lee Humphreys. Landline natives: Telephone practices since the 1950s as innovation. Technology and Culture, 61(3):923–947, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/763594.

Akallah:2022:HWE

[Aka22] Jethron Ayumbah Akallah. A History of Water Engineering and Management in Yemen: Material Remains and Textual Foundations by Ingrid Hehmeyer (review). Technology and Culture, 63 (4):1191–1192, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/868092.

Akselrad:2023:VCH

[Aks23] Daniel Akselrad. Visions of control: The head-up display, perceptual labor, and a lesson for augmented reality. Technology and Culture, 64(3):761–789, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/903972.

Arapostathis:2020:GTA

[AL20] Stathis Arapostathis and Léonard Laborie. Governing technosciences in the Age of Grand Challenges: A European historical perspective on the entanglement of science, technology, diplomacy, and democracy. Technology and Culture, 61(1):318–332, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
752933.

Aldrich:2020:FSL

[Ald20] Mark Aldrich. Fake Silk: The Lethal History of Viscose Rayon by Paul David Blanc (review). Technology and Culture, 61(3):980– 981, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
763598.

Alekna:2022:NNN

[Ale22] John Alekna. Neither nation nor empire: Situating Shanghai radio in a global technological moment, 1922–25. Technology and Culture, 63(4):1078–1105, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/868054.

Alexander:2023:TWC

[Ale23] Jennifer Karns Alexander. Technology in World Civilization: A Thousand-Year History by Arnold Pacey and Francesca Bray (review). Technology and Culture, 64(2):578–580, April 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893049.

Alsina:2020:GDE

[Als20] Marc J. Alsina. The Grand Designers: The Evolution of the Airplane in the 20th Century by John D. Anderson, Jr. (review). Technology and Culture, 61(3):976–978, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/763596.

Alsina:2022:APC

[Als22] Marc J. Alsina. Aviation for the people: Class and state aviation in Perón’s “New Argentina,” 1946–55. Technology and Culture, 63(1):153–181, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/844167.

Ammermann:2022:FNJ

[Amm22] Friedrich N. Ammermann. In the Forest of No Joy: The Congo– Océan Railroad and the Tragedy of French Colonialism by J. P. Daughton (review). Technology and Culture, 63(4):1187–1188, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
868070.

Andersen:2021:SSN

[And21] Håkon With Andersen. Sailing School: Navigating Science and Skill, 1550–1800 by Margaret E. Schotte (review). Technology and Culture, 62(4):1245–1246, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/819813.

Andersen:2022:EEA

[And22] Casper Andersen. Experten der Erschließung: Akteure der deutschen Kolonialtechnik in Afrika und Europa 1890–1943 [Experts in development: German colonial technologists in Africa and Europe 1890–1943] by Sebastian Beese (review). Technology and Culture, 63(3):889–890, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/859746.

Angus:2023:RMW

[Ang23] Siobhan Angus. Rare Merit: Women in Photography in Canada, 1840–1940 by Colleen Skidmore (review). Technology and Culture, 64(4):1347–1348, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/911036.

Anonymous:2020:AIV

[Ano20a] Anonymous. Author index: Volume 61 (2020). Technology and Culture, 61(4):1271–1280, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/777657.

Anonymous:2020:A

[Ano20b] Anonymous. Awards. Technology and Culture, 61(2):662–677, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
761580.

Anonymous:2020:TIV

[Ano20c] Anonymous. Title index: Volume 61 (2020). Technology and Culture, 61(4):1281–1287, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/777658.

Anonymous:2020:VNJa

[Ano20d] Anonymous. Volume 61, number 1, January 2020. Technology and Culture, 61(1):??, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2020:VNA

[Ano20e] Anonymous. Volume 61, number 2, April 2020. Technology and Culture, 61(2):??, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2020:VNS

[Ano20f] Anonymous. Volume 61, number 2 supplement, April 2020. Technology and Culture, 61(2S):??, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/

194.

Anonymous:2020:VNJb

[Ano20g] Anonymous. Volume 61, number 3, July 2020. Technology and Culture, 61(3):??, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2020:VNO

[Ano20h] Anonymous. Volume 61, number 4, October 2020. Technology and Culture, 61(4):??, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2021:AIV

[Ano21a] Anonymous. Author index to volume 62 (2021). Technology and Culture, 62(4):1297–1307, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/819814.

Anonymous:2021:SHT

[Ano21b] Anonymous. Society for the history of technology awards and fellowships 2020. Technology and Culture, 62(3):861–872, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/803108.

Anonymous:2021:SH

[Ano21c] Anonymous. Storytelling historians. Technology and Culture, 62 (4):1217–1218, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/819777.

Anonymous:2021:TIV

[Ano21d] Anonymous. Title index to volume 62 (2021). Technology and Culture, 62(4):1308–1313, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/819815.

Anonymous:2021:VNJa

[Ano21e] Anonymous. Volume 62, number 1, January 2021. Technology and Culture, 62(1):??, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2021:VNA

[Ano21f] Anonymous. Volume 62, number 2, April 2021. Technology and Culture, 62(2):??, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2021:VNJb

[Ano21g] Anonymous. Volume 62, number 3, July 2021. Technology and Culture, 62(3):??, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2021:VNO

[Ano21h] Anonymous. Volume 62, number 4, October 2021. Technology and Culture, 62(4):??, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2022:AIV

[Ano22a] Anonymous. Author index vol. 63 (2002). Technology and Culture, 63(4):1261–1277, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/868101.

Anonymous:2022:SHT

[Ano22b] Anonymous. Society for the history of technology awards and fellowships 2021. Technology and Culture, 63(2):507–522, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/854012.

Anonymous:2022:V

[Ano22c] Anonymous. Vol 63 (2022). Technology and Culture, 63(4): 1272–1277, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/868101#sub01.

Anonymous:2022:VNJa

[Ano22d] Anonymous. Volume 63, number 1, January 2022. Technology and Culture, 63(1):??, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2022:VNA

[Ano22e] Anonymous. Volume 63, number 2, April 2022. Technology and Culture, 63(2):??, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2022:VNJb

[Ano22f] Anonymous. Volume 63, number 3, July 2022. Technology and Culture, 63(3):??, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2022:VNO

[Ano22g] Anonymous. Volume 63, number 4, October 2022. Technology and Culture, 63(4):??, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2023:SHT

[Ano23a] Anonymous. Society for the history of technology awards and fellowships 2022. Technology and Culture, 64(2):550–565, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/893047.

Anonymous:2023:VNJa

[Ano23b] Anonymous. Volume 64, number 1, January 2023. Technology and Culture, 64(1):??, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2023:VNA

[Ano23c] Anonymous. Volume 64, number 2, April 2023. Technology and Culture, 64(2):??, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2023:VNJb

[Ano23d] Anonymous. Volume 64, number 3, July 2023. Technology and Culture, 64(3):??, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL /journal/194.

Anonymous:2023:VNO

[Ano23e] Anonymous. Volume 64, number 4, October 2023. Technology and Culture, 64(4):??, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/journal/194.

Anonymous:2024:VNJ

[Ano24a] Anonymous. Volume 65, number 1, January 2024. Technology and Culture, 65(1):??, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/journal/194.

Anonymous:2024:VNA

[Ano24b] Anonymous. Volume 65, number 2, April 2024. Technology and Culture, 65(2):??, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/journal/194.

Arceneaux:2020:SCG

[Arc20] Noah Arceneaux. Screen Culture: A Global History by Richard Butsch (review). Technology and Culture, 61(4):1217–1219, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
777623.

Arceneaux:2023:ACL

[Arc23a] Noah Arceneaux. Assembly Codes: The Logistics of Media ed. by Matthew Hockenberry, Nicole Starosielski, and Susan Zieger (review). Technology and Culture, 64(2):604–606, April 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893063.

Arceneaux:2023:CEB

[Arc23b] Noah Arceneaux. Cellular: An Economic and Business History of the International Mobile-Phone Industry by Daniel D. Garcia-Swartz and Martin Campbell-Kelly (review). Technology and Culture, 64(3):1005–1007, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/904020.

Arnold:2023:REI

[Arn23a] David Arnold. Railways’ Economic Impact on Uttar Pradesh and Colonial North India (1860–1914): The Iron Raj by Ian D. Derbyshire (review). Technology and Culture, 64(2):643–645, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893084.

Arnold:2023:CCB

[Arn23b] David Arnold. The camera and the cattle: Bovine photography and technologies of “Improvement” in colonial South India. Technology and Culture, 64(4):1045–1070, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/910994.

Arun:2023:LHW

[Aru23] P. Arun. The Listeners: A History of Wiretapping in the United States by Brian Hochman (review). Technology and Culture, 64 (1):229–230, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/879611.

Ash:2021:ETR

[Ash21] Eric H. Ash. An Empire Transformed: Remolding Bodies and Landscapes in the Restoration Atlantic by Kate Luce Mulry (review). Technology and Culture, 62(4):1243–1244, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/819812.

Askari:2022:USL

[Ask22] Kaveh Askari. Underground: The Secret Life of Videocassettes in Iran by Blake Atwood (review). Technology and Culture, 63 (2):572–573, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/854032.

Asseraf:2023:MMC

[Ass23] Arthur Asseraf. Media of the Masses: Cassette Culture in Modern Egypt by Andrew Simon (review). Technology and Culture, 64 (4):1349–1350, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/911037.

Atwood:2023:TWI

[Atw23] Blake Atwood. Television at Work: Industrial Media and American Labor by Kit Hughes (review). Technology and Culture, 64 (1):240–241, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/879613.

Auderset:2022:EML

[Aud22] Juri Auderset. Eigensinnige Musterschüler: Ländliche Entwicklung und internationales Expertenwissen in der Türkei (1947– 1980) [Headstrong model students: Rural development and international expert knowledge in Turkey (1947–1980)] by Heinrich Hartmann (review). Technology and Culture, 63(1):268–270, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
844189.

Augustine:2020:RNT

[Aug20] Dolores L. Augustine. Radiation Nation: Three Mile Island and the Political Transformation of the 1970s by Natasha Zaretsky (review). Technology and Culture, 61(3):981–983, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/763621.

Augustine:2021:SIK

[Aug21] Dolores L. Augustine. Strahlen im Kalten Krieg. Nuklearer Alltag und atomarer Notfall in der Schweiz [Radiation in the Cold War: The nuclear everyday and atomic emergency in Switzerland] by Sibylle Marti, and: Strom für die Republik. Die Stasi und das Kernkraftwerk Greifswald [Electricity for the Republic: The Stasi and the Greifswald nuclear power plant] by Sibylle Marti (review). Technology and Culture, 62(3):918–921, July 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/803114.

Aumann:2021:SKU

[Aum21] Philipp Aumann. Vom Scheitern der kybernetischen Utopie. Die Entwicklung von Überwachung und Informationsverarbeitung im MfS [About the collapse of the cybernetic utopia: Surveillance and information processing in the State Security Service] by Christian Booß (review). Technology and Culture, 62(3):923–925, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/803116.

Babintseva:2022:PSH

[Bab22] Ekaterina Babintseva. The Power of Systems: How Policy Sciences Opened Up the Cold War World by Eglė Rindzevičiūtė (review). Technology and Culture, 63(3):886–888, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/859745.

Backer:2022:FRI

[Bac22] Kellen Backer. The Fishmeal Revolution: The Industrialization of the Humboldt Current by Kristin A. Wintersteen (review). Technology and Culture, 63(2):589–590, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/854042.

Bae:2023:CMT

[Bae23] Keung Yoon Bae. Cine-Mobility: Twentieth-Century Transformations in Korea’s Film and Transportation by Han Sang Kim (review). Technology and Culture, 64(4):1345–1346, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/911035.

Bantjes:2021:OMA

[Ban21] Rod Bantjes. The optical machine’s asynchronic progress: Perceptual paradigms and 3-D enhancement technologies, 1700–1925. Technology and Culture, 62(4):1119–1142, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/819773.

Barron:2021:THN

[Bar21a] Géraldine Barron. Toward a history of nautical archaeology: Boats and history by Éric Rieth (review). Technology and Culture, 62 (3):931–933, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/803120.

Bartos:2021:AKT

[Bar21b] Jeffrey Bartos. The accumulated knowledge of a thousand generations: U.S. mining engineers as public intellectuals, 1885–1920. Technology and Culture, 62(4):1172–1198, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/819775.

Barnett:2022:KEA

[Bar22] William C. Barnett. Krakow: An Ecobiography by Adam Izdebski and Rafał Szmytka (review). Technology and Culture, 63 (2):538–540, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/854052.

Barbier:2023:TIC

[Bar23] Laura Barbier. Tóxicos invisibles: La construcción de la ignorancia ambiental ed. by Ximo Guillem-Llobat and Agustí Nieto-Galan (review). Technology and Culture, 64(2):599–600, April 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893060.

Bassett:2021:TGA

[Bas21] Ross Bassett. Les techniques et la globalization au XXe siècle [Technology and globalization in the Twentieth-Century] ed. by Liliane Hilaire–Pérez and Larissa Zakharova (review). Technology and Culture, 62(1):257–258, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/783915.

Baudry:2020:PIP

[Bau20] Jérôme Baudry. A politics of intellectual property: Creating a patent system in revolutionary France. Technology and Culture, 61(4):1017–1044, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/777601.

Baudry:2023:RLC

[Bau23] Jérôme Baudry. Les Réparations dans l’Histoire: Cultures techniques et savoir-faire dans la longue durée ed. by Gianenrico Bernasconi et al. (review). Technology and Culture, 64(3):934–935, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
903982.

Bergman:2021:SIC

[BBC^{+}21] Yoel Bergman, Florian Bettel, Maria Elvira Callapez, Donna J. Drucker, Francesco Gerali, Dick van Lente, Tiina Männistö-Funk, Stefan Poser, and Thomas Schuetz. The 47th symposium of the International Committee for the History of Technology, ICOHTEC’s first digital meeting, July 15–17, 2020. Technology and Culture, 62(3):873–893, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/803109.

Blomkvist:2020:RIU

[BE20] Pär Blomkvist and Martin Emanuel. Regulating innovation: Utility vs. leisure in Swedish moped history, 1952–1961. Technology and Culture, 61(3):815–842, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/763584.

Beatty:2020:TTM

[Bea20] Edward (Ted) Beatty. Technology and Tradition in Mesoamerica after the Spanish Invasion: Archeological Perspectives ed. by Rani T. Alexander (review). Technology and Culture, 61(3):967–968, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
763595.

Beardsley:2021:FAC

[Bea21] Amanda Beardsley. The female absorption coefficient: The miniskirt study, gender, and latter-day Saint architectural acoustics. Technology and Culture, 62(3):659–684, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/803101.

Becker:2023:WTC

[Bec23] Bert Becker. Western Technology and China’s Industrial Development: Steamship Building in Nineteenth-Century China, 1828– 1895 by Hsien-ch’un Wang (review). Technology and Culture, 64 (2):645–646, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893085.

Bednarek:2022:CAG

[Bed22] Janet R. Daly Bednarek. Civil Aviation and the Globalization of the Cold War by Peter Svik (review). Technology and Culture, 63(2):583–585, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/854039.

Beerbuhl:2021:BKG

[Bee21] Margrit Schulte Beerbühl. Braunschweig und der Kaffee. Die Geschichte des Röstkaffeemarktes von den Anfängen bis in unsere Tage [Braunschweig and coffee: The history of the coffee roasting market from the beginning to the present] by Peter Albrecht (review). Technology and Culture, 62(2):597–598, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794381.

Beeckmans:2023:ACH

[Bee23] Luce Beeckmans. Age of Concrete: Housing and the Shape of Aspiration in the Capital of Mozambique by David Morton, and: The Politics of Housing in (Post-)Colonial Africa ed. by Kirsten Rüther et al. (review). Technology and Culture, 64(3):939–942, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
903985.

Bekus:2024:RST

[Bek24] Nelly Bekus. The Road to Socialism: Transport Infrastructure in Socialist Bulgaria and Yugoslavia (1945–1989) by Lyubomir Pozharliev (review). Technology and Culture, 65(2):727–729, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
926347.

Beltran:2021:DFF

[Bel21] Alain Beltran. D’eau et de feu: forges et énergie hydraulique, XVIII^{e}–XX^{e} siècle. Une histoire singulière de l’industrialisation française [Of water and fire: Forges and hydraulic power, 18th– 20th centuries. A singular history of French industrialization] by Serge Benoît (review). Technology and Culture, 62(3):966–967, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
803140.

Bellido:2023:PME

[Bel23] Jose Bellido. Patents in miniature: The effects of microfilm as an information technology, 1938–68. Technology and Culture, 64 (2):407–433, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893041.

Benedict:2022:KYR

[Ben22] Carol A. Benedict. Know Your Remedies: Pharmacy and Culture in Early Modern China by He Bian (review). Technology and Culture, 63(1):244–246, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/844176.

Berner:2021:BRT

[Ber21a] Boel Berner. Blood Relations: Transfusion and the Making of Human Genetics by Jenny Bangham (review). Technology and Culture, 62(3):908–909, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/803147.

Berrones:2021:MMR

[Ber21b] Jethro Hernández Berrones. Mestizo Modernity: Race, Technology, and the Body in Post-revolutionary Mexico by David S. Dalton (review). Technology and Culture, 62(4):1268–1269, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/819792.

Berry:2021:CEM

[Ber21c] Dominic J. Berry. Collecting Experiments: Making Big Data Biology by Bruno J. Strasser (review). Technology and Culture, 62 (1):283–285, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/783930.

Berth:2021:CHM

[Ber21d] Christiane Berth. Connected: How a Mexican Village Built Its Own Cell Phone Network by Roberto J. González (review). Technology and Culture, 62(4):1278–1279, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/819797.

Bertucci:2021:SHI

[Ber21e] Paola Bertucci. Spinners’ hands, Imperial minds: Migrant labor, embodied expertise, and the failed transfer of silk technology across the Atlantic. Technology and Culture, 62(4):1003–1031, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
819770.

Bernasconi:2022:SCG

[Ber22a] Gianenrico Bernasconi. Swansea Copper: A Global History by Chris Evans and Louise Miskell (review). Technology and Culture, 63(1):255–257, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/844182.

Bertocco:2022:SCM

[Ber22b] Mattia Bertocco. Storia Culturale del Made in Italy [The cultural history of “Made in Italy”] by Carlo Marco Belfanti (review). Technology and Culture, 63(2):537–538, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/854051.

Berth:2023:ICG

[Ber23] Christiane Berth. Interrupted conversations: Gender and telephone use in Mexico, 1930s–70s. Technology and Culture, 64(1): 124–148, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/879584.

Bertilorenzi:2024:AAD

[Ber24] Marco Bertilorenzi. Le acque agitate della patria: L’industrializzazione del Piave (1882–1966) [The troubled waters of the homeland: The industrialization of the Piave River (1882–1966)] by Giacomo Bonan (review). Technology and Culture, 65(2):724–725, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/926345.

Bess:2022:PRP

[Bes22] Michael K. Bess. The Panama Railroad by Peter Pyne (review). Technology and Culture, 63(4):1189–1190, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/868081.

Barrett-Fox:2024:RJC

[BF24] Jason Barrett-Fox. Riding Jane Crow: African American Women on the American Railroad by Miriam Thaggert (review). Technology and Culture, 65(1):367–368, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/920534.

Bhimull:2022:SER

[BHJI^{+}22] Chandra Bhimull, Gabrielle Hecht, Edward Jones-Imhotep, Chakanetsa Mavhunga, Lisa Nakamura, and Asif Siddiqi. Systemic and epistemic racism in the history of technology. Technology and Culture, 63(4):935–952, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/868048.

Biasillo:2022:ATP

[Bia22] Roberta Biasillo. Acque di Terraferma: il Padovano [Inland Waters: The Padua Province] ed. by Pedro Garcia Duarte and Yann Giraud (review). Technology and Culture, 63(2):540–542, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
854014.

Bimm:2022:FBM

[Bim22] Jordan Bimm. Far Beyond the Moon: A History of Life Support Systems in the Space Age by David P. D. Munns and Kärin Nickelsen (review). Technology and Culture, 63(1):302–304, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
844207.

Binder:2023:ACB

[Bin23] Jeffrey M. Binder. Algorithmic Culture Before the Internet by Ted Striphas (review). Technology and Culture, 64(4):1362–1363, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
911043.

Bini:2024:AAU

[Bin24] Elisabetta Bini. The Atomic Archipelago: US Nuclear Submarines and Technopolitics of Risk in Cold War Italy by Davide Orsini (review). Technology and Culture, 65(1):384–386, January 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/920543.

Bitter:2021:DSS

[Bit21] Joella Bitter. Digital Sound Studies ed. by Mary Caton Lingold et al. (review). Technology and Culture, 62(1):255–257, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/783914.

Bix:2024:MCM

[Bix24] Amy Sue Bix. Mid-Century Modernism and the American Body: Race, Gender, and the Politics of Power in Design by Kristina Wilson (review). Technology and Culture, 65(1):369–370, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920535.

Black:2020:BGH

[Bla20a] Brian C. Black. Burning Up: A Global History of Fossil Fuel Consumption by Simon Pirani (review). Technology and Culture, 61 (2):700–702, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/761565.

Black:2020:OAS

[Bla20b] Jeremy Black. The Origins of American Strategic Bombing Theory by Craig Morris (review). Technology and Culture, 61(3):970–971, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
763613.

Black:2021:TWD

[Bla21] Brian C. Black. A Thousand Ways Denied: The Environmental Legacy of Oil in Louisiana by John T. Arnold (review). Technology and Culture, 62(3):935–937, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/803123.

Blaszczyk:2023:CNH

[Bla23] Regina Lee Blaszczyk. The Culture of Nature in the History of Design ed. by Kjetil Fallan (review). Technology and Culture, 64 (3):1016–1017, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/904026.

Blalock:2024:OGW

[BMSS24] Stephanie M. Blalock, Kevin McMullen, Stefan Schöberlein, and Jason Stacy. “One of the grand works of the world”: Walt Whitman’s advocacy for the Brooklyn Waterworks, 1856–59. Technology and Culture, 65(1):237–263, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/920522.

Beech:2023:SPP

[BN23] Lucy Beech and Tamar Novick. Sex panic and the productive infertility of the freemartin. Technology and Culture, 64(4):1071–1092, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
910995.

Bohlman:2021:HMS

[Boh21] Andrea F. Bohlman. Hush: Media and Sonic Self-Control by Mack Hagood (review). Technology and Culture, 62(4):1280–1281, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
819798.

Bolman:2021:SUH

[Bol21] Brad Bolman. Sulphuric Utopias: A History of Maritime Fumigation by Lukas Engelmann and Christos Lynteris (review). Technology and Culture, 62(2):616–617, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/794392.

Bonea:2021:MLW

[Bon21a] Amelia Bonea. The Magic Lantern at Work: Witnessing, Persuading, Experiencing, and Connecting ed. by Martyn Jolly and Elisa deCourcy (review). Technology and Culture, 62(1):302–304, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
783941.

Bonea:2021:CTP

[Bon21b] Amelia Bonea. “Contagion by telephone”: Print media and knowledge about infectious diseases in Britain, 1880s–1914. Technology and Culture, 62(4):1063–1086, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/819772.

Bonan:2023:HEA

[Bon23] Giacomo Bonan. Hydraulic engineers and antiquarians: Political use of the past in Nineteenth-Century Venice. Technology and Culture, 64(3):845–874, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/903975.

Booth:2022:LBG

[Boo22] Douglas Booth. The Lure of the Beach: A Global History by Robert C. Ritchie (review). Technology and Culture, 63(4): 1232–1234, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/868079.

Booth:2024:CME

[Boo24] Rodrigo Booth. La ciudad en movimiento: Estudios históricos sobre transporte colectivo y movilidad en Santiago de Chile, siglos XIX y XX [The city on the move: Historical studies on public transport and mobility in Santiago de Chile, nineteenth and twentieth centuries] ed. by Simón Castillo and Marcelo Mardones (review). Technology and Culture, 65(2):675–677, April 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/926321.

Borg:2023:PTM

[Bor23] Kevin L. Borg. The Persistence of Technology: From Maintenance and Repair to Reuse and Disposal ed. by Heike Weber and Stefan Krebs (review). Technology and Culture, 64(3):936–937, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
903983.

Boschiero:2020:MMM

[Bos20] Luciano Boschiero. Machines, motion, mechanics: Philosophers engineering the fountains of Versailles. Technology and Culture, 61(4):1108–1128, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/777605.

Bothereau:2020:IPR

[Bot20] Benjamin Bothereau. Illuminated publics: Representations of street lamps in revolutionary France. Technology and Culture, 61(4):1045–1075, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/777602.

Bothereau:2021:IPE

[Bot21] Benjamin Bothereau. Iluminated Paris: Essays on Art and Lighting in the Belle Époque by Hollis Clayson (review). Technology and Culture, 62(1):304–306, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/783942.

Bouk:2021:ANQ

[Bou21] Dan Bouk. America by the Numbers: Quantification, Democracy, and the Birth of National Statistics by Emmanuel Didier (review). Technology and Culture, 62(2):608–610, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794390.

Bousquet:2022:CCU

[Bou22] Antoine Bousquet. Camp Century: The Untold Story of America’s Secret Arctic Military Base under the Greenland Ice by Kristian H. Nielsen and Henry Nielsen (review). Technology and Culture, 63(4):1234–1235, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/868080.

Bouvier:2023:MEM

[Bou23] Yves Bouvier. Making Energy Markets: The Origins of Electricity Liberalisation in Europe by Ronan Bolton (review). Technology and Culture, 64(2):597–598, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893059.

Boyd:2021:PPP

[Boy21] William Boyd. Pollution, Politics, and Power: The Struggle for Sustainable Electricity by Tom McGarity (review). Technology and Culture, 62(1):265–267, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/783920.

Bray:2020:CMM

[Bra20] Francesca Bray. The craft of mud-making: Cropscapes, time, and history. Technology and Culture, 61(2):645–661, April 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/761579.

Brescia:2023:PDF

[Bre23] Ray Brescia. The Paradox of Democracy: Free Speech, Open Media, and Perilous Persuasion by Zac Gershberg and Sean Illing (review). Technology and Culture, 64(2):601–602, April 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893061.

Bridges:2020:ALD

[Bri20] R. Shaw Bridges. American Lucifers: The Dark History of Artificial Light, 1750–1865 by Jeremy Zallen (review). Technology and Culture, 61(3):990–992, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/763620.

Brock:2021:LAM

[Bro21a] David C. Brock. The Long Arm of Moore’s Law: Microelectronics and American Science by Cyrus Mody (review). Technology and Culture, 62(4):1260–1262, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/819787.

Brown:2021:PAI

[Bro21b] Hayley Brown. Polio across the Iron Curtain: Hungary’s Cold War with an Epidemic by Dóra Vargha (review). Technology and Culture, 62(1):278–280, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/783927.

Brown:2022:LLC

[Bro22] Logan Brown. Learning to love computers: Useful cinema and the mediation of American computing, 1958–62. Technology and Culture, 63(3):665–688, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/859720.

Bruggemeier:2021:PEH

[Brü21] Franz-Josef Brüggemeier. Powering Empire: How Coal Made the Middle East and Sparked Global Carbonization by On Barak (review). Technology and Culture, 62(1):264–265, January 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/783919.

Brule:2023:SMS

[Bru23] Émeline Brulé. From Savage Minds to Savage Machines: Racial Science and Twentieth-Century Design by Ginger Nolan (review). Technology and Culture, 64(2):586–587, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893053.

Buckley:2023:BFI

[Buc23] Eve Buckley. Before the Flood: The Itaipu Dam and the Visibility of Rural Brazil by Jacob Blanc (review). Technology and Culture, 64(1):290–292, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/879592.

Buns:2023:NBG

[Bun23] Melina Antonia Buns. Nuclear Bodies: The Global Hibakusha by Robert A. Jacobs (review). Technology and Culture, 64(3):983– 985, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
904008.

Burk:2023:SRH

[Bur23] Dan L. Burk. Software Rights: How Patent Law Transformed Software Development in America by Gerardo Con Diaz (review). Technology and Culture, 64(2):617–618, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893070.

Butsch:2023:SSW

[But23] Richard Butsch. Schools and Screens: A Watchful History by Victoria Cain (review). Technology and Culture, 64(1):234–236, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
879594.

Cain:2022:TMH

[Cai22] Victoria Cain. Teaching Machines: The History of Personalized Learning by Audrey Watters (review). Technology and Culture, 63(3):853–855, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/859727.

Calhoun:2021:WWM

[Cal21] Gloria Calhoun. Why wire mattered: Building U.S. networked infrastructures, 1845–1910. Technology and Culture, 62(1):156–184, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
783909.

Calascibetta:2022:AMC

[Cal22] Franco Calascibetta. Autarchia e multinazionali: Casi di imprese estere del settore chimico in Italia durante il Fascismo [Autarky and multinationals: Cases of foreign companies in Italy’s chemical sector during Fascism] ed. by Marco Bertilorenzi (review). Technology and Culture, 63(4):1248–1249, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/868089.

Campbell:2021:WID

[Cam21a] Bruce B. Campbell. Wireless Internationalism and Distant Listening: Britain, Propaganda, and the Invention of Global Radio, 1920–1939 by Simon J. Potter (review). Technology and Culture, 62(3):947–948, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/803129.

Camprubi:2021:PTR

[Cam21b] Lino Camprubí. Provokationen der technikgeschichte. Zum reflexionszwang historischer Forschung [Provocations in the History of Technology: On the imperative to reflect in historical research] ed. by Martina Heßler und Heike Weber (review). Technology and Culture, 62(1):253–255, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/783913.

Camprubi:2022:CTL

[Cam22] Lino Camprubí. Ciencia en transición. El lastre franquista ante el reto de la modernización [Science in transition. The Francoist burden and modernization’s challenge] ed. by Lorenzo Delgado and Santiago M. López (review). Technology and Culture, 63 (2):554–555, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/854021.

Camprubi:2023:PHS

[Cam23] Lino Camprubí. Portuguese history since the Late Middle Ages through the history of technology: A four-volume overview. Technology and Culture, 64(2):574–577, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893048.

Camp:2024:OBH

[Cam24] Michael Camp. Oil Beach: How Toxic Infrastructure Threatens Life in the Ports of Los Angeles and Beyond by Christina Dunbar-Hester (review). Technology and Culture, 65(1):371–372, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920536.

Cardon:2021:CCL

[Car21] Nathan Cardon. Cycling on the color line: Race, technology, and bicycle mobilities in the Early Jim Crow South, 1887–1905. Technology and Culture, 62(4):973–1002, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/819769.

Castro:2021:FMQ

[Cas21] J. Justin Castro. Ferrocarril y modernización en Quito: Un cambio dramático entre 1905 y 1922 [The railway and modernization in Quito: A dramatic change between 1905 and 1922] by Wilson Miño Grijalva (review). Technology and Culture, 62(2):628–629, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
794398.

Castro:2024:EHL

[Cas24] J. Justin Castro. Engineering history in Latin America: a review of Spanish-language books. Technology and Culture, 65(2):651–666, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
926319.

Cater:2021:WBN

[Cat21] Casey P. Cater. Wires That Bind: Nation, Region, and Technology in the Southwestern United States, 1854–1920 by Torsten Kathke (review). Technology and Culture, 62(4):1270–1271, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
819793.

Cirac-Claveras:2021:PEE

[CC21] Gemma Cirac-Claveras. Program Earth: Environmental Sensing Technology and the Making of a Computational Planet by Jennifer Gabrys (review). Technology and Culture, 62(1):297–299, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
783938.

Ceruzzi:2020:MIW

[Cer20] Paul Ceruzzi. Making IT Work: A History of the Computer Services Industry by Jeffrey R. Yost (review). Technology and Culture, 61(1):377–378, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/752958.

Chatterjee:2020:TRC

[Cha20] Animesh Chatterjee. Technology and Rural Change in Eastern India: 1830–1980 by Smritikumar Sarkar (review). Technology and Culture, 61(4):1219–1220, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/777624.

Chang:2021:ACM

[Cha21a] Jiat-Hwee Chang. Air-Conditioning in Modern American Architecture, 1890–1970 by Joseph M. Siry (review). Technology and Culture, 62(4):1254–1256, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/819784.

Charnley:2021:EMO

[Cha21b] Berris Charnley. Evolution Made to Order: Plant Breeding and Technological Innovation in Twentieth-Century America by Helen Anne Curry (review). Technology and Culture, 62(1):285–286, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
783931.

Chatzis:2021:MIE

[Cha21c] Konstantinos Chatzis. Mobilités d’ingénieurs en Europe, XVe– XVIIIe siècle [Mobility of engineers in Europe, 15th–18th century] ed. by Stéphane Blond et al. (review). Technology and Culture, 62 (2):636–637, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/794384.

Channell:2022:PMT

[Cha22a] David F. Channell. Physical Models: Their Historical and Current Use in Civil and Building Engineering Design ed. by Bill Addis (review). Technology and Culture, 63(4):1250–1251, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
868090.

Chapron:2022:LTA

[Cha22b] Emmanuelle Chapron. Le livre technique avant le XXe siècle: À l’échelle du monde [Technical books before the 1900s: On a global scale] ed. by Liliane Hilaire–Pérez et al. (review). Technology and Culture, 63(3):904–906, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/859754.

Chatzis:2022:BCT

[Cha22c] Konstantinos Chatzis. Building capabilities for a transnational world: Foreign languages at the École des Ponts et Chaussées, 1790s–1850s. Technology and Culture, 63(3):775–807, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/859724.

Chen:2022:CCC

[Che22] BuYun Chen. The craft of color and the chemistry of dyes: Textile technology in the Ryukyu Kingdom, 1700–1900. Technology and Culture, 63(1):87–117, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/844165.

Cheung:2023:DMT

[Che23] S. Y. Cheung. Designing Motherhood: Things that Make and Break Our Births by Michelle Millar Fisher and Amber Winick (review). Technology and Culture, 64(4):1307–1308, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/911015.

Choi:2022:NSD

[Cho22] Hyungsub Choi. Nanoscale: Society’s Deep Impact on Science, Technology and Innovation in India by Pankaj Sekhsaria, and: Instrumental Lives: An Intimate Biography of an Indian Laboratory by Pankaj Sekhsaria (review). Technology and Culture, 63(3):859– 861, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
859730.

Churella:2021:SCR

[Chu21] Albert J. Churella. Steam City: Railroads, Urban Space, and Corporate Capitalism in Nineteenth-Century Baltimore by David Schley (review). Technology and Culture, 62(3):960–961, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/803137.

Chunikhin:2022:EEC

[Chu22a] Kirill Chunikhin. Establishing eye contact with one historical photograph. Technology and Culture, 63(3):599–602, July 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/859717.

Chunikhin:2022:RRH

[Chu22b] Kirill Chunikhin. Risk and respirators: The hazardous trajectories of Soviet occupational safety, 1940s–80s. Technology and Culture, 63(3):603–633, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/859718.

Churella:2023:CRN

[Chu23] Albert Churella. The Coming of the Railway: a New Global History, 1750–1850 by David Gwyn (review). Technology and Culture, 64(4):1331–1333, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/911028.

Cieraad:2021:HWC

[Cie21] Irene Cieraad. How the Working-Class Home Became Modern, 1900–1940 by Thomas C. Hubka (review). Technology and Culture, 62(3):953–954, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/803133.

Cieraad:2023:DWF

[Cie23] Irene Cieraad. Dwelling in the World: Family, House, and Home in Tianjin, China, 1860–1960 by Elizabeth LaCouture (review). Technology and Culture, 64(3):938–939, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/903984.

Ciglioni:2023:EAH

[Cig23] Laura Ciglioni. Engaging the Atom: The History of Nuclear Energy and Society in Europe from the 1950s to the Present ed. by Arne Kaijser et al. (review). Technology and Culture, 64(3):987– 988, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
904010.

Clark:2021:MCI

[Cla21] Kim Clark. A Modern Contagion: Imperialism and Public Health in Iran’s Age of Cholera by Amir A. Afkhami (review). Technology and Culture, 62(2):610–611, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/794380.

Cliver:2020:CIR

[Cli20] Robert Cliver. The Chinese and the Iron Road: Building the Transcontinental Railroad ed. by Gordon H. Chang and Shelley Fisher Fishkin (review). Technology and Culture, 61(4):1258–1259, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
777650.

Coates:2022:PRC

[Coa22] Jennifer Coates. Powers of the Real: Cinema, Gender, and Emotion in Interwar Japan by Diane Wei Lewis (review). Technology and Culture, 63(3):910–911, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/859757.

Coates:2024:NRM

[Coa24] Jamie Coates. The New Real: Media and Mimesis in Japan from Stereographs to Emoji by Jonathan E. Abel (review). Technology and Culture, 65(1):412–414, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/920558.

Cohn:2021:PLE

[Coh21] Julie Cohn. Power Lines: Electricity in American Life and Letters, 1882–1952 by Jennifer L. Lieberman (review). Technology and Culture, 62(1):267–269, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/783921.

Cohle:2023:URT

[Coh23] Ana Cohle. Under the Radar: Tracking Western Radio Listeners in the Soviet Union by R. Eugene Parta (review). Technology and Culture, 64(4):1350–1352, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/911038.

Collopy:2020:DWF

[Col20] Peter Sachs Collopy. Documenting the World: Film, Photography, and the Scientific Record ed. by Gregg Mitman and Kelley Wilder (review). Technology and Culture, 61(2):697–698, April 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/761563.

Coleman:2021:EPE

[Col21] Leo Coleman. Empire and Post-Empire Telecommunications in India: A History by Pradip Ninan Thomas (review). Technology and Culture, 62(1):316–318, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/783949.

Conner:2022:PTS

[Con22] Thomas H. Conner. Photography from the Turin Shroud to the Turing Machine by Yanai Toister (review). Technology and Culture, 63(3):912–913, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/859758.

Coopersmith:2020:SFG

[Coo20] Jonathan Coopersmith. Spectacular Flops: Game-Changing Technologies That Failed by Michael Brian Schiffer (review). Technology and Culture, 61(4):1235–1236, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/777637.

Cooper:2023:IPH

[Coo23] Austin R. Cooper. Inspectors for Peace: A History of the International Atomic Energy Agency by Elisabeth Roehrlich (review). Technology and Culture, 64(1):254–256, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/879626.

Coppola:2024:BLT

[Cop24] Al Coppola. British Literature and Technology, 1600–1830 ed. by Kristin M. Girten and Aaron R. Hanlon (review). Technology and Culture, 65(2):709–710, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/926337.

Cortada:2020:GGH

[Cor20] James W. Cortada. From Gutenberg to Google: The History of Our Future by Tom Wheeler (review). Technology and Culture, 61(1):380–381, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/752957.

Cornillie:2023:MAC

[Cor23a] Thomas C. Cornillie. The Machine Age: Curation and memory in two U.S. exhibits. Technology and Culture, 64(2):566–573, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893090.

Cortada:2023:PCV

[Cor23b] James W. Cortada. Prophets of Computing: Visions of Society Transformed by Computing ed. by Dick van Lente (review). Technology and Culture, 64(4):1360–1362, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/911042.

Cossart:2020:GCA

[Cos20] Brice Cossart. The Gun in Central Africa: A History of Technology and Politics by Giacomo Macola (review). Technology and Culture, 61(3):968–970, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/763610.

Cowan:2021:GMM

[Cow21] Ruth Schwartz Cowan. Growth: From Microorganisms to Megacities by Vaclav Smil (review). Technology and Culture, 62(3):916– 917, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
803113.

Cox:2022:CBM

[Cox22] Peter Cox. Cycling and the British: A Modern History by Neil Carter (review). Technology and Culture, 63(1):282–283, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
844196.

Crangle:2021:RPO

[Cra21] Richard Crangle. Robert Paul and the Origins of British Cinema by Ian Christie (review). Technology and Culture, 62(4): 1282–1283, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/819799.

Crim:2022:CWI

[Cri22] Brian E. Crim. Cork Wars: Intrigue and Industry in World War II by David A. Taylor (review). Technology and Culture, 63(2):567–568, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/854029.

Crook:2023:ASW

[Cro23] Tom Crook. The Architecture of Steam: Waterworks and the Victorian Sanitary Crisis by James Douet (review). Technology and Culture, 64(4):1310–1312, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/911017.

Cruz:2021:HAI

[Cru21] Felipe Fernandes Cruz. Hacking airspace: The insurgent technology of Brazil’s hot air balloons, 1970–present. Technology and Culture, 62(1):27–56, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/783904.

Cubasch:2023:LSN

[Cub23] Alwin J. Cubasch. Life in Space: NASA Life Sciences Research during the Late Twentieth Century by Maura Phillips Mackowski (review). Technology and Culture, 64(2):627–629, April 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893075.

Cuenca:2022:BMW

[Cue22] Catherine Cuenca. Beyond the museum walls: The pandemic and the prototypes exhibit at the Conservatoire National des Arts et Métiers. Technology and Culture, 63(2):477–482, April 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/854009.

Curtis:2023:GGP

[Cur23] Stephan Curtis. Germs and Governance: The Past, Present and Future of Hospital Infection, Prevention and Control ed. by Ann Marie Rafferty, Marguerite Dupree, and Fay Bound Albertini (review). Technology and Culture, 64(3):930–931, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/903980.

Cai:2024:BCB

[CW24] Cong Cai and Di Wu. Building Chinese Braille with Chinese Characteristics by Huang Nai (review). Technology and Culture, 65(1):414–416, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/920559.

dAvignon:2020:SGP

[d’A20] Robyn d’Avignon. Spirited geobodies: Producing subterranean property in Nineteenth-Century Bambuk, West Africa. Technology and Culture, 61(2S):S20–S48, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/761592.

Dack:2020:SWT

[Dac20] John Dack. Sonic Writing: Technologies of Material, Symbolic and Signal Inscriptions by Thor Magnusson (review). Technology and Culture, 61(1):373–375, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/752949.

Daemmrich:2021:IIE

[Dae21] Arthur Daemmrich. From Insight to Innovation: Engineering Ideas That Transformed America in the Twentieth Century by David P. Billington, Jr. (review). Technology and Culture, 62 (4):1262–1264, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/819788.

Daen:2024:CUI

[Dae24] Laurel Daen. Creating a user-inventor community: How disabled people innovated and marketed disability in early nineteenthcentury America. Technology and Culture, 65(1):117–141, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920518.

Daheur:2023:NEO

[Dah23] Jawad Daheur. A New Ecological Order: Development and the Transformation of Nature in Eastern Europe ed. by Ştefan Dorondel and Stelu Şerban (review). Technology and Culture, 64(1): 286–288, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/879601.

Dalton:2022:HCF

[Dal22] David S. Dalton. Historia de la ciencia ficción latinoamericana I: Desde los orígenes hasta la modernidad [A history of Latin American science fiction: From its origins to modernity] ed. by Teresa López–Pellisa and Silvia G. Kurlat Ares (review). Technology and Culture, 63(4):1227–1228, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/868076.

Daniels:2021:BEG

[Dan21] Mario Daniels. British Exploitation of German Science and Technology, 1943–1949 by Charlie Hall (review). Technology and Culture, 62(4):1264–1265, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/819789.

Das:2024:ADC

[Das24] Debojyoti Das. Agrarian Development in Colonial India: The British and Bihar by Peter Robb (review). Technology and Culture, 65(1):374–376, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/920538.

Davids:2020:TTT

[Dav20] Karel Davids. A Tale of Three Thirsty Cities: The Innovative Water Systems of Toledo, London and Paris in the Second Half of the Sixteenth Century by Jaime-Chaim Shulman (review). Technology and Culture, 61(3):953–955, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/763617.

Davids:2023:IMM

[Dav23a] Karel Davids. Ingenuity in the Making: Matter and Technique in Early Modern Europe ed. by Richard J. Oosterhoff, José Ramón Marcaida, and Alexander Marr (review). Technology and Culture, 64(1):259–260, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/879623.

Davids:2023:SMC

[Dav23b] Mila Davids. Strands of Modernization: The Circulation of Technology and Business Practices in East Asia, 1850–1920 ed. by David B. Sicilia and David G. Wittner (review). Technology and Culture, 64(2):641–643, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893083.

Dazhi:2023:EPA

[Daz23] Yao Dazhi. Edmond Pâris et l’art naval: Des pirogues aux cuirassés by Géraldine Barron (review). Technology and Culture, 64(3):960–961, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/903995.

deCrouy-Chanel:2022:AMH

[dCC22] Emmanuel de Crouy-Chanel. Les artilleurs et la monarchie hispanique (1560–1610): Guerres, savoirs techniques, État [Artillerymen and the Hispanic monarchy (1560–1610): Wars, technical knowledge, and the State] by Brice Cossart (review). Technology and Culture, 63(4):1251–1253, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/868091.

DeSmet:2023:HPA

[De 23a] Elsa De Smet. Hybrid Practices: Art in Collaboration with Science and Technology in the Long 1960s ed. by David Cateforis, Steven Duval, and Shepherd Steiner (review). Technology and Culture, 64(1):264–266, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/879595.

Wildt:2023:AAD

[De 23b] Annemarie De Wildt. Artefacts, Archives, and Documentation in the Relational Museum by Mike Jones (review). Technology and Culture, 64(4):1299–1301, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/911011.

DeCourcy:2022:PMT

[DeC22] Elisa DeCourcy. Photography’s Materialities: Transatlantic Photographic Practices over the Long Nineteenth Century ed. by Geoff Bender and Rasmus R. Simonsen (review). Technology and Culture, 63(4):1223–1225, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/868074.

Degani:2022:GGH

[Deg22] Michael Degani. Gone to Ground: A History of Environment and Infrastructure in Dar es Salaam by Emily Brownell (review). Technology and Culture, 63(2):542–544, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/854015.

Delis:2021:BSR

[Del21] Apostolos Delis. Bridging the Seas: The Rise of Naval Architecture in the Industrial Age, 1800–2000 by Larrie D. Ferreiro (review). Technology and Culture, 62(2):588–589, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794393.

Denning:2020:MEC

[Den20a] Andrew Denning. Mobilizing empire: The Citroën central Africa expedition and the interwar civilizing mission. Technology and Culture, 61(1):42–70, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/752960.

Dennis:2020:TTH

[Den20b] Michael Aaron Dennis. The TVs of Tomorrow: How RCA’s Flatscreen Dreams Led to the First LCDs by Benjamin Gross (review). Technology and Culture, 61(4):1240–1241, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/777640.

Denton:2023:CPG

[Den23] Chad B. Denton. Are carcasses political? German veterinarians and the modernization of rendering technology, 1864–1940. Technology and Culture, 64(1):90–123, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/879583.

DiTullio:2021:HWM

[Di 21] Matteo Di Tullio. The History of Water Management in the Iberian Peninsula: Between the 16th and 19th Centuries ed. by Ana Duarte Rodrigues and Carmen Toribio Marín (review). Technology and Culture, 62(4):1247–1248, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/819780.

Diebolt:2021:SSF

[Die21] Claude Diebolt. Les Sciences sociales face à Vichy. Le colloque ^{≪} Travail et Techniques ^{≫} de 1941 [Social sciences against Vichy: The “work and techniques” workshop in 1941] ed. by Isabelle Gouarné (review). Technology and Culture, 62(1):287–288, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
783932.

Dienel:2022:UTF

[Die22] Hans-Liudger Dienel. A U-Turn to the Future: Sustainable Urban Mobility since 1850 ed. by Martin Emanuel et al. (review). Technology and Culture, 63(3):872–874, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/859737.

DiMoia:2021:DHS

[DiM21] John P. DiMoia. “Difficult heritage” & selective elision: The Seoul Power Plant. Technology and Culture, 62(2):561–572, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/794378.

DiMoia:2022:MSC

[DiM22] John P. DiMoia. Manual for Survival: A Chernobyl Guide to the Future by Kate Brown (review). Technology and Culture, 63(1):246–248, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/844177.

Dimas:2023:TVE

[Dim23] Carlos S. Dimas. Technocratic Visions: Engineers, Technology, and Society in Mexico ed. by Justin Castro and James A. Garza (review). Technology and Culture, 64(3):972–974, July 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/904002.

Dimas:2024:WWE

[Dim24] Carlos Salvador Dimas. Water, wealth, and engineering wisdom: Shaping Tucumán’s agricultural future, 1890–1910. Technology and Culture, 65(2):623–650, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/926318.

Diogo:2024:LOS

[Dio24] Maria Paula Diogo. Lifelines of Our Society: a Global History of Infrastructures by Dirk van Laak (review). Technology and Culture, 65(2):699–701, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/926333.

Disco:2020:AIF

[Dis20] Cornelis Disco. Alles im Fluss: Die Lebensadern unserer Gesellschaft [Everything flows: The lifeblood of our society] by Dirk van Laak (review). Technology and Culture, 61(4):1216–1217, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
777621.

Disco:2021:KHA

[Dis21a] Cornelis Disco. Katrina: A History, 1915–2015 by Andy Horowitz (review). Technology and Culture, 62(2):584–585, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794400.

Disco:2021:WBN

[Dis21b] Nil Disco. Water Brings No Harm: Management Knowledge and the Struggle for the Waters of Kilimanjaro by Matthew V. Bender (review). Technology and Culture, 62(4):1248–1250, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/819781.

Dong:2022:RBI

[Don22] Yuting Dong. Red brick imperialism: How vernacular knowledge shaped Japanese colonial expertise in Northeast China, 1905–45. Technology and Culture, 63(1):118–152, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/844166.

Donev:2023:ECM

[Don23] Alexander Donev. Early Cinema, Modernity and Visual Culture: The Imaginary of the Balkans by Ana Grgić (review). Technology and Culture, 64(1):236–238, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/879606.

Donges:2024:AII

[Don24] Alexander Donges. American Independent Inventors in an Era of Corporate R&D by Eric S. Hintz (review). Technology and Culture, 65(1):403–405, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/920553.

Drucker:2020:TSS

[Dru20] Donna J. Drucker. Toxic Shock: A Social History by Sharra L. Vostral (review). Technology and Culture, 61(4):1261–1263, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
777652.

Duan:2023:CTR

[Dua23] Shaoming Duan. [Chinese text] [Research on the modernization of Chinese handicrafts (1840–1911)] by Chen Cong (review). Technology and Culture, 64(3):967–968, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/903999.

Duhem:2022:LTA

[Duh22] Ludovic Duhem. L’analogie dans les techniques [Analogy in techniques] ed. by Sophie de Beaune et al. (review). Technology and Culture, 63(1):234–235, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/844170.

Dupre:2021:AES

[Dup21] Sven Dupré. Artisanal Enlightenment: Science and the Mechanical Arts in Old Regime France by Paola Bertucci (review). Technology and Culture, 62(3):956–958, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/803135.

Duque:2024:RGG

[Duq24] Lorena Campuzano Duque. A Ritual Geology: Gold and Subterranean Knowledge in Savanna West Africa by Robyn D’Avignon (review). Technology and Culture, 65(2):718–720, April 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/926342.

Durco:2023:RDC

[Ďur23] Michal Ďurčo. The Rise and Decline of Communist Czechoslovakia’s Railway Sector by Tomáš Nigrin (review). Technology and Culture, 64(3):956–958, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/903993.

deWildt:2022:COC

[dWB22] Annemarie de Wildt and Errol Boom. Creating an online community: Corona in the city. Technology and Culture, 63(2):461–470, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
854007.

deZwarte:2021:PGW

[dZ21] Ingrid de Zwarte. Plants Go to War: A Botanical History of World War II by Judith Sumner (review). Technology and Culture, 62 (4):1295–1296, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/819807.

Eames:2021:CBS

[Eam21] Anthony Eames. A “Corruption of British science?”: The strategic defense initiative and British technology policy. Technology and Culture, 62(3):812–838, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/803106.

Editors:2020:SYS

[Edi20a] The Editors. Sixty years of scholarship. Technology and Culture, 61(4):1196, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/777614.

Editors:2020:THT

[Edi20b] The Editors. Take 1: Historians of technology watching Chernobyl. Technology and Culture, 61(4):1149–1153, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/777609.

Editors:2022:HCA

[Edi22] The Editors. How curators and archivists responded to the pandemic. Technology and Culture, 63(2):458–460, April 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/854006.

Edward:2023:CEI

[Edw23] Frank Edward. The City Electric: Infrastructure and Ingenuity in Postcolonial Tanzania by Michael Degani (review). Technology and Culture, 64(4):1329–1331, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/911027.

ElHariry:2021:OOT

[El 21] Shorouk El Hariry. On our own terms: Towards a history of Arab technological landscapes and cultures. Technology and Culture, 62(1):241–252, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/783912.

Ellenbogen:2020:FEB

[EL20] Josh Ellenbogen and Alison Langmead. Forms of equivalence: Bertillonnage and the history of information management. Technology and Culture, 61(1):207–238, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/752965.

Ellis:2022:LOS

[Ell22] Thomas Ellis. Limiting Outer Space: Astroculture After Apollo ed. by Alexander C. T. Geppert, and: Militarizing Outer Space: Astroculture, Dystopia and the Cold War ed. by Alexander C. T. Geppert, Daniel Brandau, and Tilmann Siebeneichner (review). Technology and Culture, 63(4):1253–1256, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/868093.

Emmerson:2021:RRR

[Emm21] Allison L. C. Emmerson. Recycling and Reuse in the Roman Economy ed. by Chloë N. Duckworth and Andrew Wilson (review). Technology and Culture, 62(2):606–608, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794391.

Engels:2020:VHB

[Eng20] Jens Ivo Engels. Versorgung der Haupstadt der Bewegung: Infrastrukturen und Statgesellschaft im Nationalsozialistischen München [Supplying the capital of the movement: Infrastructures and urban society in national socialist Munich] by Mathias Irlinger (review). Technology and Culture, 61(4):1243–1245, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/777642.

Engels:2022:RAH

[Eng22] Jens Ivo Engels. Rhythm analysis: A heuristic tool for historical infrastructure research. Technology and Culture, 63(3):830–852, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
859726.

Erdogan:2021:CDC

[Erd21] Julia Gül Erdogan. Crime Dot Com: From Viruses to Vote Rigging, How Hacking Went Global by Geoff White (review). Technology and Culture, 62(4):1272–1273, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/819794.

Ericson:2021:JPJ

[Eri21] Kjell D. Ericson. Judging the Perle Japonaise: The techno–legal separation of culture from nature in 1920s Paris. Technology and Culture, 62(4):1032–1062, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/819771.

Ertsen:2021:FWS

[Ert21] Maurits W. Ertsen. Finnish Water Services: Experiences in Global Perspective by Tapio S. Katko (review). Technology and Culture, 62(1):295–297, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/783937.

Esposito:2023:OSC

[Esp23] James Esposito. Oxygen sense: Creating embodied knowledge to promote health innovation in the Royal air force, 1939–45. Technology and Culture, 64(1):34–62, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/879581.

Esselborn:2021:MVA

[Ess21] Stefan Esselborn. Moving Violations: Automobiles, Experts, and Regulations in the United States by Lee Vinsel (review). Technology and Culture, 62(1):290–292, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/783934.

Evans:2021:TIR

[Eva21] Chris Evans. Technology in the Industrial Revolution by Barbara Hahn (review). Technology and Culture, 62(3):964–965, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/803139.

Fallan:2024:TMH

[Fal24] Kjetil Fallan. Transparency: The Material History of an Idea by Daniel Jütte (review). Technology and Culture, 65(2):711–712, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
926338.

Fari:2024:CSL

[Far24] Simone Fari. Il Canale di Suez e l’Italia (1856–1869) [The Suez Canal and Italy (1856–1869)] by Andrea Giuntini (review). Technology and Culture, 65(2):729–731, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/926348.

Ferguson:2023:SMM

[Fer23a] Christopher Ferguson. The Soho Manufactory, Mint and Foundry, West Midlands: Where Boulton, Watt and Murdoch Made History by George Demidowicz (review). Technology and Culture, 64(4):1314–1316, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/911019.

Fernandez:2023:TDP

[Fer23b] Luke Fernandez. Transparent Designs: Personal Computing and the Politics of User-Friendliness by Michael Black (review). Technology and Culture, 64(2):618–620, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893071.

Fernandes:2024:CSH

[Fer24] Gilberto Fernandes. As ceaseless as the sea: How modern construction machines disrupted Canadian senses and sensibilities, 1870s– 1940s. Technology and Culture, 65(1):211–236, January 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/920521.

Filippello:2023:AMT

[Fil23] Marcus Filippello. African Motors: Technology, Gender, and the History of Development by Joshua Grace (review). Technology and Culture, 64(3):944–946, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/903987.

Finkelstein:2022:ECA

[Fin22] Clemens Finkelstein. Echo’s Chambers: Architecture and the Idea of Acoustic Space by Joseph L. Clarke (review). Technology and Culture, 63(1):295–296, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/844203.

Flade:2023:CCC

[Fla23] Falk Flade. Conflict in Cooperation: Crossborder Infrastructures in Europe Facing the Second World War ed. by Valentine Aldebert et al. (review). Technology and Culture, 64(3):948–950, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/903989.

Fontaine:2023:DLD

[Fon23a] Marion Fontaine. La désindustrialisation de la Lorraine du fer by Pascal Raggi (review). Technology and Culture, 64(1):274–276, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
879624.

Fontanon:2023:CPM

[Fon23b] Claudine Fontanon. Claude-Pierre Molard (1759–1837): Un technicien dans la cité ed. by Liliane Hilaire–Pérez and François Jarrige (review). Technology and Culture, 64(3):965–967, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/903998.

Fouche:2023:MBL

[Fou23] Rayvon Fouché. The Matter of Black Living: The Aesthetic Experiment of Racial Data, 1880–1930 by Autumn Womack (review). Technology and Culture, 64(2):593–594, April 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893057.

Fox:2021:THM

[Fox21] Robert Fox. Die Technische Hochschule München im Nationalsozialismus [The Technical University of Munich under National Socialism] ed. by Wolfgang A. Herrmann and Winfried Nerdinger (review). Technology and Culture, 62(3):910–911, July 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/803148.

Frederick-Rothwell:2021:MAC

[FR21] Betsy Frederick-Rothwell. Modern Architecture and Climate: Design before Air Conditioning by Daniel Barber (review). Technology and Culture, 62(2):632–633, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/794383.

Francks:2020:MMC

[Fra20] Penelope Francks. Mass Media, Consumerism and National Identity in Postwar Japan by Martyn David Smith (review). Technology and Culture, 61(3):958–960, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/763619.

Fraser:2021:SIB

[Fra21] Jennifer Fraser. Seeing infrared: Breast cancer, Inuit, and the extractive colonality of disease distributions and diagnostic imaging technologies. Technology and Culture, 62(3):709–740, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/803103.

Frehner:2020:RCF

[Fre20] Brian Frehner. A River in the City of Fountains: An Environmental History of Kansas City and the Missouri River by Amahia K. Mallea (review). Technology and Culture, 61(4):1245–1247, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
777643.

Freidenfelds:2022:IAR

[Fre22] Lara Freidenfelds. IVF and Assisted Reproduction: A Global History by Sarah Ferber et al. (review). Technology and Culture, 63(1):248–249, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/844178.

Friedel:2022:BBL

[Fri22a] Robert Friedel. Beyond Bakelite: Leo Baekeland and the Business of Science and Invention by Joris Mercelis (review). Technology and Culture, 63(1):254–255, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/844181.

Friss:2022:CPP

[Fri22b] Evan Friss. Cycling Pathways: The Politics and Governance of Dutch Cycling Infrastructure, 1920–2020 by Henk-Jan Dekker (review). Technology and Culture, 63(3):880–881, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/859741.

Frohlich:2023:IUT

[Fro23] Xaq Frohlich. Introduction: Unearthing technology in public histories of food. Technology and Culture, 64(3):903–908, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/903977.

Frumer:2020:IPG

[Fru20] Yulia Frumer. Imperialism and profit in the guise of benevolence: Mizuno, Moore, and DiMoia, eds., Engineering Asia. Technology and Culture, 61(1):333–336, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/752934.

Frumer:2021:RSR

[Fru21] Yulia Frumer. Rise of the Self-Replicators: Early Visions of Machines, AI and Robots That Can Reproduce and Evolve by Tim Taylor and Alan Dorin (review). Technology and Culture, 62(4): 1274–1275, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/819795.

Fullilove:2024:DSD

[Ful24] Courtney Fullilove. Debate: Seeds as deep time technologies. Technology and Culture, 65(1):7–38, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/920514.

Futselaar:2022:EST

[Fut22] Ralf Futselaar. On an Empty Stomach: Two Hundred Years of Hunger Relief by Tom Scott-Smith (review). Technology and Culture, 63(4):1239–1241, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/868084.

Fava-Verde:2020:MPC

[FV20] Jean-François Fava-Verde. Managing privacy: Cryptography or private networks of communication in the Nineteenth Century. Technology and Culture, 61(3):798–814, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/763583.

Gallardo-Albarran:2021:LDM

[GA21] Daniel Gallardo-Albarrán. Landscapes of Disease: Malaria in Modern Greece by Katerina Gardikas (review). Technology and Culture, 62(1):280–281, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/783928.

Gabriac:2022:PPO

[Gab22] Amandine Gabriac. La plaque photographique – Un outil pour la fabrication et la diffusion des savoirs (XIXe–XXe siècle) [The photographic plate: A tool for the production and diffusion of knowledge (19th–20th centuries)] ed. by Denise Borlée and Hervé Doucet (review). Technology and Culture, 63(1):288–289, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/844199.

Gaboury:2023:UPP

[Gab23] Jacob Gaboury. Uncomputable: Play and Politics in the Long Digital Age by Alexander R. Galloway (review). Technology and Culture, 64(3):1007–1009, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/904021.

Gagnon:2021:PCH

[Gag21] Justine Gagnon. Powering Up Canada: A History of Power, Fuel, and Energy from 1600 ed. by R. W. Sandwell (review). Technology and Culture, 62(1):269–271, January 2021. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/783922.

Galviz:2022:LSL

[Gal22] Carlos López Galviz. Last Subway: The Long Wait for the Next Train in New York City by Philip Mark Plotch (review). Technology and Culture, 63(3):874–876, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/859738.

Gardebo:2020:EEP

[Gär20a] Johan Gärdebo. Engineering the Environment: Phytotrons and the Quest for Climate Control in the Cold War by David P. D. Munns (review). Technology and Culture, 61(3):985–986, July 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/763614.

Garner:2020:CEC

[Gar20b] Kenneth Garner. Corporeality in Early Cinema: Viscera, Skin, and Physical Form ed. by Marina Dahlquist, et al. (review). Technology and Culture, 61(2):695–696, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/761562.

Garda:2024:HDP

[Gar24] Maria B. Garda. Hacking Diversity: The Politics of Inclusion in Open Technology Cultures by Christina Dunbar-Hester (review). Technology and Culture, 65(2):740–742, April 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/926354.

Gatejel:2020:TES

[Gat20] Luminita Gatejel. Technology and the Environment in State-Socialist Hungary: An Economic History by Viktor Pál (review). Technology and Culture, 61(3):960–962, July 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/763615.

Goncu-Berk:2022:MDM

[GB22] Gozde Goncu-Berk. Making Disability Modern: Design Histories ed. by Bess Williamson and Elizabeth Guffey (review). Technology and Culture, 63(1):241–243, January 2022. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/844174.

Gomez-Cruz:2024:DEM

[GC24] Edgar Gómez-Cruz. Different Engines: Media Technologies From Latin America by Andrés Burbano (review). Technology and Culture, 65(2):697–699, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/926332.

Gerali:2021:BEF

[Ger21a] Francesco Gerali. The Birth of Energy: Fossil Fuels, Thermodynamics, and the Politics of Work by Cara New Daggett (review). Technology and Culture, 62(4):1287–1289, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/819803.

Gere:2021:STU

[Ger21b] Charlie Gere. Science, Technology, and Utopias: Women Artists and Cold War America by Christine Filippone (review). Technology and Culture, 62(2):590–591, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/794394.

Gerritsen:2021:EIM

[Ger21c] Anne Gerritsen. Entangled Itineraries: Materials, Practices, and Knowledges across Eurasia ed. by Pamela H. Smith (review). Technology and Culture, 62(1):260–262, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/783917.

Gil-Farrero:2021:CNC

[GF21] Judit Gil-Farrero. Cultivating Nature: The Conservation of a Valencian Working Landscape by Sarah R. Hamilton (review). Technology and Culture, 62(3):939–941, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/803125.

Gideon:2024:CSP

[Gid24] Jasmine Gideon. Conjuring the State: Public Health Encounters in Highland Ecuador, 1908–1945 by A. Kim Clark (review). Technology and Culture, 65(2):681–682, April 2024. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/926324.

Giersch:2022:MRM

[Gie22] C. Patterson Giersch. Mountain Rivers, Mountain Roads: Transport in Southwest China, 1700–1850 by Nanny Kim (review). Technology and Culture, 63(1):284–285, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/844197.

Gilligan:2020:CCP

[Gil20] Ian Gilligan. Ceramics in Circumpolar Prehistory: Technology, Lifeways, and Cuisine ed. by Peter Jordan and Kevin Gibbs (review). Technology and Culture, 61(4):1221–1222, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/777626.

Gil:2024:TAE

[Gil24a] Magdalena Gil. ! Alerta! Engineering on Shaky Ground by Elizabeth Reddy (review). Technology and Culture, 65(2):688–689, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
926327.

Giles:2024:WPC

[Gil24b] Kendall Giles. Words and Power: Computers, Language, and U.S. Cold War Values by Bernadette Longo (review). Technology and Culture, 65(2):742–744, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/926355.

Giuntini:2022:ISC

[Giu22] Andrea Giuntini. Imperial Science: Cable Telegraphy and Electrical Physics in the Victorian British Empire by Bruce J. Hunt (review). Technology and Culture, 63(2):576–577, April 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/854034.

Guillem-Llobat:2023:TSK

[GL23] Ximo Guillem-Llobat. Technology and subaltern knowledge at the Catalan Olive Oil Museum. Technology and Culture, 64(3):909– 917, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
903978.

Glisic:2023:SRA

[Gli23] Iva Glisic. Spatial Revolution: Architecture and Planning in the Early Soviet Union by Christina E. Crawford (review). Technology and Culture, 64(2):589–591, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893055.

Gross:2022:GDR

[GM22] Benjamin Gross and Allison Marsh. Going digital: The research library and the pandemic. Technology and Culture, 63 (4):1137–1139, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/868056.

Galaz-Mandakovic:2024:RRD

[GMR24] Damir Galaz-Mandakovic and Francisco Rivera. Revolution and resistance in the desert: The Guggenheim System’s impact on nitrate mining and society in Atacama, 1926–31. Technology and Culture, 65(2):603–622, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/926317.

Golaszewski:2024:EEG

[Gol24] Devon Golaszewski. Embodied Engineering: Gendered Labor, Food Security, and Taste in Twentieth-Century Mali by Laura Ann Twagira (review). Technology and Culture, 65(1):365–367, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920533.

Gooday:2021:THM

[Goo21] Graeme Gooday. Testing Hearing: The Making of Modern Aurality ed. by Viktoria Tkaczyk et al. (review). Technology and Culture, 62(2):618–619, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/794415.

Gottschang:2020:RMP

[Got20] Suzanne Z. Gottschang. Reproductive modernities in policy: Maternal mortality, midwives, and cesarean sections in China, 1900s– 2000s. Technology and Culture, 61(2):617–644, April 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/761578.

Goumri:2021:RIS

[Gou21] Maël Goumri. Risques Industriels: Savoirs, régulations, politique d’assistance fin XVIIe – début XXe siècle [Industrial risks: Knowledge, regulations and assistance policies, late 17th-early 20th century] ed. by Thomas Le Roux (review). Technology and Culture, 62(1):288–290, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/783933.

Grandi:2022:PDV

[Gra22] Elisa Grandi. La Puissance du vent: Des moulins à vent aux éoliennes modernes [The power of the wind: From windmills to modern wind turbines] by Philippe Bruyerre (review). Technology and Culture, 63(4):1236–1237, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/868082.

Green:2021:PUP

[Gre21] Nile Green. Print and the Urdu Public: Muslims, Newspapers, and Urban Life in Colonial India by Megan Eaton Robb (review). Technology and Culture, 62(3):914–915, July 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/803112.

Greenlees:2022:MAV

[Gre22] Janet Greenlees. Mapping AIDS: Visual Histories of an Enduring Epidemic by Lukas Engelmann (review). Technology and Culture, 63(4):1229–1230, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/868077.

Grove:2021:VIC

[Gro21] Linda Grove. Vernacular Industrialism in China: Local Innovation and Translated Technologies in the Making of a Cosmetics Empire, 1900–1940 by Eugenia Lean (review). Technology and Culture, 62(3):958–959, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/803136.

Gross:2022:RTC

[Gro22] Benjamin Gross. Research in the time of COVID: Virtual fellowships at the Linda Hall Library. Technology and Culture, 63 (4):1140–1156, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/868057.

Gross:2023:REE

[Gro23] Robert Groß. The Riviera, Exposed: An Ecohistory of Postwar Tourism and North African Labor by Stephen L. Harp (review). Technology and Culture, 64(3):1009–1010, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/904022.

Guedon:2022:MTA

[Gué22a] Stéphanie Guédon. Mobile Technologies in the Ancient Sahara and Beyond ed. by Chloë N. Duckworth et al. (review). Technology and Culture, 63(3):855–856, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/859728.

Guerrero:2022:SVD

[Gue22b] Saul Guerrero. Silver Veins, Dusty Lungs: Mining, Water, and Public Health in Zacatecas, 1835–1946 by Rocio Gomez (review). Technology and Culture, 63(4):1197–1199, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/868099.

Guizard:2022:CPA

[Gui22] Fabrice Guizard. Chevaux, paysans et artisans: Le travail attelé entre Meuse et Loire (XII^{e}–XVI^{e} siècle) [Horses, peasants and craftsmen: Harnessed work horses between Meuse and Loire (12– 16th centuries)] by Floriana Bardoneschi (review). Technology and Culture, 63(3):871–872, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/859736.

Guldi:2020:ATQ

[Gul20] Jo Guldi. Another take on quantitative methods: Claire Lemercier and Claire Zalc, Quantitative Methods in the Humanities. Technology and Culture, 61(1):341–343, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/752936.

Gumbert:2024:NGC

[Gum24] Heather Gumbert. News from Germany: The Competition to Control World Communications, 1900–1945 by Heidi J. S. Tworek (review). Technology and Culture, 65(1):416–418, January 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/920560.

Gunergun:2021:TSM

[Gün21] Feza Günergun. Timekeepers and sufi mystics: Technical knowledge bearers of the Ottoman Empire. Technology and Culture, 62 (2):348–372, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/794368.

Gunning:2024:RSA

[Gun24] Oliver Gunning. The Republic of Skill: Artisan Mobility, Innovation, and the Circulation of Knowledge in Premodern Europe ed. by David Garrioch (review). Technology and Culture, 65(1): 399–401, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/920551.

Gusterson:2024:LNE

[Gus24] Hugh Gusterson. Land of Nuclear Enchantment: A New Mexican History of the Nuclear Weapons Industry by Lucie Genay (review). Technology and Culture, 65(1):386–388, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/920544.

Gutting:2023:AMD

[Gut23] Alicia Gutting. Ausgestrahlt: Die mediale Debatte um “Tschernobyl” in der Bundesrepublik und in Frankreich 1986/87 by Katrin Jordan (review). Technology and Culture, 64(3):989–990, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
904011.

Ge:2024:LPS

[GX24] Yejing Ge and Weimin Xiong. The launch of a political satellite: Constructing the Ming Tombs Reservoir and “Mass engineering” in Communist China. Technology and Culture, 65(1):143–175, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920519.

Hagmann:2021:CHE

[Hag21] Johannes-Geert Hagmann. Contested heritage in East Asia: Colonial memory & technology sites. Technology and Culture, 62 (2):547–550, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/794376.

Hahn:2023:EMI

[Hah23] Barbara Hahn. Endangered Maize: Industrial Agriculture and the Crisis of Extinction by Helen Anne Curry (review). Technology and Culture, 64(3):1011–1013, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/904023.

Haigh:2021:MSP

[Hai21] Thomas Haigh. The Media Snatcher: pc/core/turbo/engine/grafx/16/cdrom2/super/duo/arcade/rx by Carl Therrien (review). Technology and Culture, 62(1):325–326, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
783954.

Haines:2023:PSI

[Hai23] Monamie Bhadra Haines. Ploughshares and Swords: India’s Nuclear Program in the Global Cold War by Jayita Sarkar (review). Technology and Culture, 64(3):981–983, July 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/904007.

Hajduk:2021:ADH

[Haj21] John Hajduk. Amplified: A Design History of the Electric Guitar by Paul Atkinson (review). Technology and Culture, 62(3):968– 969, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
803141.

Halpern:2020:AIH

[Hal20] Orit Halpern. Architectural Intelligence: How Designers and Architects Created the Digital Landscape by Molly Wright Steenson (review). Technology and Culture, 61(4):1265–1267, October 2020.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/777654.

Halevi:2021:WHA

[Hal21a] Leor Halevi. What hath allah wrought? The global invention of prescriptive machines for the Islamic consumer, 1975–2010. Technology and Culture, 62(3):741–779, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/803104.

Hall:2021:EEC

[Hal21b] Bert Hall. Engineering the Eternal City: Infrastructure, Topography, and the Culture of Knowledge in Late Sixteenth-Century Rome by Pamela O. Long (review). Technology and Culture, 62 (2):634–635, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/794404.

Hamerla:2020:SHR

[Ham20] Rich Hamerla. The Secret History of RDX: The Super-Explosive That Helped Win World War II by Colin F. Baxter (review). Technology and Culture, 61(1):366–367, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/752939.

Hammers:2021:MAT

[Ham21] Roslyn Lee Hammers. Machine Art in the Twentieth Century by Andreas Broeckmann (review). Technology and Culture, 62 (4):1258–1259, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/819786.

Hamilton:2024:SRH

[Ham24] Vivien Hamilton. Seduced by Radium: How Industry Transformed Science in the American Marketplace by Maria Rentetzi (review). Technology and Culture, 65(1):405–407, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/920554.

Hanly:2020:TTA

[Han20] B. Jack Hanly. Think Tank Aesthetics: Midcentury Modernism, the Cold War, and the Neoliberal Present by Pamela M. Lee (review). Technology and Culture, 61(4):1267–1269, October 2020.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/777655.

Hanna:2022:IMT

[Han22] Stephen R. Hanna. Imperial Military Transportation in British Asia: Burma 1941–1942 by Michael W. Charney (review). Technology and Culture, 63(3):883–884, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/859743.

Hankins:2023:AAH

[Han23] Michael W. Hankins. Assured Access: A History of the US Air Force Space Launch Enterprise, 1945–2020 by David N. Spires (review). Technology and Culture, 64(2):629–630, April 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893076.

Haard:2020:TMM

[Hår20a] Mikael Hård. Technik als Motor der Modernisierung ed. by Thomas Zoglauer, Karsten Weber and Hans Fries (review). Technology and Culture, 61(3):948–949, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/763622.

Hart:2020:PDH

[Har20b] Jennifer Hart. Of pirate drivers and honking horns: Mobility, authority, and urban planning in Late-Colonial Accra. Technology and Culture, 61(2S):S49–S76, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/761593.

Hargreaves:2021:IDM

[Har21] Gregory A. Hargreaves. Industrial Development and Manufacturing in the Antebellum Gulf South: A Reevaluation by Michael S. Frawley (review). Technology and Culture, 62(4):1289–1290, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
819804.

Hardy:2023:DIV

[Har23a] Penelope K. Hardy. Discovery, Innovation, and the Victorian Admiralty: Paper Navigators by Erika Behrisch (review). Technology and Culture, 64(3):961–963, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/903996.

Harms:2023:CWC

[Har23b] Victoria Harms. Cold War Civil Defence in Western Europe: Sociotechnical Imaginaries of Survival and Preparedness ed. by Marie Cronqvist, Rosanna Farbøl, and Casper Sylvest (review). Technology and Culture, 64(1):256–258, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/879599.

Hartmann:2024:FMT

[Har24] Heinrich Hartmann. Forming the Modern Turkish Village: Nation Building and Modernization in Rural Turkey during the Early Republic by Özge Sezer (review). Technology and Culture, 65(1): 395–397, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/920549.

Hashimoto:2021:MUS

[Has21] Takehiko Hashimoto. Making and using scientific instruments in Japan: How scholars and craftsmen cooperated, 1781–1853. Technology and Culture, 62(2):401–422, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/794370.

Hasselgren:2022:HPD

[Has22] Björn Hasselgren. Une histoire politique du canal de Göta: Technique, infrastructure et pouvoirs en Europe du Nord (années 1790– 1832) [A political history of Göta Canal: Technology, infrastructure and power in northern Europe (1790–1832)] by Thomas Gauchet (review). Technology and Culture, 63(4):1192–1194, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
868096.

Hay:2022:EPI

[Hay22] Amy M. Hay. Economic Poisoning: Industrial Waste and the Chemicalization of American Agriculture by Adam M. Romero (review). Technology and Culture, 63(3):893–894, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/859748.

Heide:2020:IRF

[Hei20] Lars Heide. IBM: The Rise and Fall and Reinvention of a Global Icon by James Cortada (review). Technology and Culture, 61 (3):978–979, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/763601.

Heinze:2023:EHO

[Hei23] Robert Heinze. Every Household Its Own Government: Improvised Infrastructure, Entrepreneurial Citizens, and the State in Nigeria by Daniel Jordan Smith (review). Technology and Culture, 64(3):943–944, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/903986.

Henckes:2021:TDE

[Hen21] Nicolas Henckes. The Technical Delusion: Electronics, Power, Insanity by Jeffrey Sconce (review). Technology and Culture, 62 (1):309–311, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/783945.

Hersey:2020:HHF

[Her20] Mark D. Hersey. Harvest of Hazards: Family Farming, Accidents, and Expertise in the Corn Belt, 1940–1975 by Derek S. Oden (review). Technology and Culture, 61(4):1260–1261, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/777651.

Herran:2023:PFN

[Her23] Néstor Herran. Political Fallout: Nuclear Weapons Testing and the Making of a Global Environmental Crisis by Toshihiro Higuchi (review). Technology and Culture, 64(3):985–986, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/904009.

Hesselbein:2020:UTW

[Hes20a] Chris Hesselbein. Ungendering Technology: Women Retooling the Masculine Sphere by Carol J. Haddad (review). Technology and Culture, 61(2):706–708, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/761568.

Hessler:2020:BLA

[Hes20b] Martina Hessler. Bored, Lonely, Angry, Stupid: Changing Feelings About Technology, From the Telegraph to Twitter by Luke Fernandez and Susan J. Matt (review). Technology and Culture, 61(4):1222–1224, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/777627.

Hessler:2022:NHM

[Heß22] Martina Heßler. A New History of Modern Computing by Thomas Haigh and Paul E. Ceruzzi (review). Technology and Culture, 63 (2):580–581, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/854037.

Hetherington:2023:ASE

[Het23] Kregg Hetherington. The Age of the Soybean: an Environmental History of Soy During the Great Acceleration ed. by Claiton Marcio da Silva and Claudio de Majo (review). Technology and Culture, 64(4):1326–1327, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/911025.

Henrich-Franke:2021:RHP

[HF21] Christian Henrich-Franke. The Radio Hobby, Private Associations, and the Challenge of Modernity in Germany by Bruce B. Campbell (review). Technology and Culture, 62(2):648–649, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794386.

Huang:2022:RTS

[HG22] Xing Huang and Xiaowu Guan. Review of twenty-six volumes of history of science and technology in China. Technology and Culture, 63(4):1168–1184, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/868059.

Humeres:2024:DBF

[HG24] Mónica Humeres and Magdalena Gil. Dreaming of a bright future: Statistics, disaster, and the birth of energopolitics in 1930s Chile. Technology and Culture, 65(2):571–602, April 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/926316.

Hua-Henning:2021:ORB

[HH21] Jan Hua-Henning. Opening the red Box: The fire alarm telegraph and politics of risk response in Imperial Germany, 1873– 1900. Technology and Culture, 62(3):685–708, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/803102.

Heymann:2020:CET

[HHK^{+}20] Matthias Heymann, Per Högselius, Elena Kochetkova, John Martin, Ole Sparenberg, Frank Veraart, and Anna Åberg. Challenging Europe: Technology, environment, and the quest for resource security. Technology and Culture, 61(1):282–294, January 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/752930.

Hill:2023:CES

[Hil23a] Rachel Hill. Cover essay: Stamping Soviet cosmonauts, craft, and Cosmos. Technology and Culture, 64(3):651–664, July 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/903967.

Hill:2023:IAL

[Hil23b] Rachel Hill. Introducing the act of looking at technologyin-operation. Technology and Culture, 64(4):1227–1233, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
911002.

Hirsh:2023:CHB

[Hir23] Richard F. Hirsh. Charged: A History of Batteries and Lessons for a Clean Energy Future by James Morton Turner (review). Technology and Culture, 64(2):595–596, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893058.

Hisano:2023:ATS

[His23a] Ai Hisano. Acquired Tastes: Stories about the Origins of Modern Food ed. by Benjamin R. Cohen, Michael S. Kideckel, and Anna Zeide (review). Technology and Culture, 64(1):297–299, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
879598.

Hisano:2023:EAB

[His23b] Ai Hisano. Eye appeal is buy appeal. Technology and Culture, 64(4):1235–1247, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/911003.

Hachez-Leroy:2022:CFH

[HL22] Florence Hachez-Leroy. The Chemistry of Fear: Harvey Wiley’s Fight for Pure Food by Jonathan Rees (review). Technology and Culture, 63(4):1241–1243, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/868085.

Hasenohrl:2020:ECH

[HM20] Ute Hasenöhrl and Jan-Henrik Meyer. The energy challenge in historical perspective. Technology and Culture, 61(1):295–306, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
752931.

Hockenberry:2023:PHC

[Hoc23] Matthew Hockenberry. The Prison House of the Circuit: Politics of Control from Analog to Digital by Jeremy Packer et al. (review). Technology and Culture, 64(4):1358–1360, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/911041.

Hofmann:2021:PCS

[Hof21] Anna Catharina Hofmann. The Politics of Chemistry: Science and Power in Twentieth-Century Spain by Agustí Nieto-Galan (review). Technology and Culture, 62(1):323–324, January 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/783953.

Hof:2023:SUP

[Hof23] Barbara Hof. The Squares: US Physical and Engineering Scientists in the Long 1970s by Cyrus C. M. Mody (review). Technology and Culture, 64(1):266–267, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/879621.

Hogselius:2022:ASO

[Hög22] Per Högselius. Atomic shocks of the old: Putting water at the center of nuclear energy history. Technology and Culture, 63(1):1–30, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
844162.

Holtorf:2021:FNW

[Hol21] Cornelius Holtorf. The Future of Nuclear Waste: What Art and Archaeology Can Tell Us about Securing the World’s Most Hazardous Material by Rosemary A. Joyce (review). Technology and Culture, 62(3):921–923, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/803115.

Holzberger:2022:BIF

[Hol22] Helena Holzberger. Bildfabriken. Industrie und Fotografie im Zarenreich (1860–1917) [Image factories: Industry and photography in Tsarist Russia (1860–1917)] by Lenka Fehrenbach (review). Technology and Culture, 63(1):286–287, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/844198.

Homburg:2022:ZCB

[Hom22] Ernst Homburg. Zinc for Coin and Brass: Bureaucrats, Merchants, Artisans, and Mining Laborers in Qing China, ca. 1680– 1830s by Hailian Chen (review). Technology and Culture, 63(1): 257–259, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/844183.

Hornsey:2022:BEL

[Hor22] Richard Hornsey. Becoming “Escalator-Legged” in interwar London: Mechanization, habit, and the mobile body. Technology and Culture, 63(4):1005–1032, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/868051.

Horning:2023:AET

[Hor23] Susan Schmidt Horning. The Age of Electroacoustics: Transforming Science and Sound by Roland Wittje (review). Technology and Culture, 64(1):245–247, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/879632.

Howard:2022:IAS

[How22] Deborah Howard. “Ingegnosi artificij.” Serenissima Repubblica di Venezia: Trecento anni di storia della scienza, della tecnica e dell’innovazione (1474–1788). [ “Ingenious artifices.” The most serene Republic of Venice: Three hundred years of history of science, technology and innovation (1474–1788)] by Roberto Berveglieri (review). Technology and Culture, 63(1):250–251, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
844179.

Howell:2023:BMS

[How23] Joel Howell. Building Medical Schools, Making Doctors: Architecture and the Modern American Physician by Katherine L. Carroll (review). Technology and Culture, 64(4):1309–1310, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/911016.

Haigh:2020:CCC

[HP20] Thomas Haigh and Mark Priestley. Contextualizing Colossus: Codebreaking technology and institutional capabilities. Technology and Culture, 61(3):871–900, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/763592.

Hatton-Proulx:2021:FPH

[HP21a] Clarence Hatton-Proulx. Face à la puissance. Une histoire des énergies alternatives à l’âge industriel [Facing power: A history of alternative energies in the industrial age] ed. by François Jarrige and Alexis Vrignon (review). Technology and Culture, 62(1): 271–272, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/783923.

Hilaire-Perez:2021:CTB

[HP21b] Liliane Hilaire-Pérez. The codification of techniques: Between bureaucracy and the markets in Early Modern Europe from a global perspective. Technology and Culture, 62(2):442–466, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/794372.

Hubbard:2024:OUG

[Hub24] Jennifer Hubbard. Oceans Under Glass: Tank Craft & the Sciences of the Sea by Samantha Muka (review). Technology and Culture, 65(1):407–408, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/920555.

Huhle:2024:WIF

[Huh24] Teresa Huhle. Wissen im Fluss: Der lateinamerikanische Staudammbau im 20. Jahrhundert als globale Wissensgeschichte [Knowledge in flux: Latin American dam construction in the twentieth century as a global history of knowledge] by Frederik Schulze (review). Technology and Culture, 65(2):695–697, April 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/926331.

Humeres:2024:LAA

[Hum24] Mónica Humeres. Living with Algorithms: Agency and User Culture in Costa Rica by Ignacio Siles (review). Technology and Culture, 65(2):683–684, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/926325.

Hunt:2020:STH

[Hun20] Bruce J. Hunt. Submarine Telegraphy and the Hunt for Gutta Percha: Challenge and Opportunity in a Global Trade by Helen Godfrey (review). Technology and Culture, 61(4):1247–1248, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
777644.

Hurt:2020:RDE

[Hur20] R. Douglas Hurt. Regenerating Dixie: Electric Energy in the Modern South by Casey P. Cater (review). Technology and Culture, 61(3):988–990, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/763600.

Husain:2023:DDS

[Hus23] Faisal H. Husain. To dam or not to dam: The social construction of an Ottoman hydraulic project, 1701–02. Technology and Culture, 64(2):456–484, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893043.

Hvattum:2024:BET

[Hva24] Mari Hvattum. The Built Environment through the Prism of the Colonial Periodical Press ed. by Alice Santiago Faria, Anne Shelley, and Sandra Ataíde Lobo (review). Technology and Culture, 65(1):429–431, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/920567.

Ignaciuk:2021:CPS

[Ign21] Agata Ignaciuk. The Circulation of Penicillin in Spain: Health, Wealth and Authority by María Jesús Santesmases (review). Technology and Culture, 62(2):622–623, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/794411.

Ignaciuk:2022:IMC

[Ign22] Agata Ignaciuk. Innovation and maladjustment: Contraceptive technologies in state-socialist Poland, 1950s–1970s. Technology and Culture, 63(1):182–208, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/844168.

Imaizumi:2022:WEG

[Ima22] Asuka Imaizumi. Widespread enthusiasm: Grassroots participation and regional variation in Early Japanese patenting, 1885– 99. Technology and Culture, 63(2):377–400, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/854002.

Inglis:2022:TCC

[Ing22] James Inglis. The Typewriter Century: A Cultural History of Writing Practices by Martyn Lyons (review). Technology and Culture, 63(4):1216–1217, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/868069.

Ionescu:2021:SEH

[Ion21] Tudor B. Ionescu. Safe Enough? A History of Nuclear Power and Accident Risk by Thomas R. Wellock (review). Technology and Culture, 62(4):1285–1287, October 2021. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/819802.

Ismail:2023:EES

[Ism23] Shehab Ismail. The engineer as economist: Sewers and the making of the water consumer in colonial Cairo, 1890. Technology and Culture, 64(2):434–455, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893042.

Israel:2023:SAI

[Isr23] Paul Israel. Scientific Americans: Invention, Technology, and National Identity by Susan Branson (review). Technology and Culture, 64(4):1316–1318, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/911020.

Jackson:2021:EAL

[Jac21] Donald C. Jackson. Engineering America: The Life and Times of John A. Roebling by Richard Haw (review). Technology and Culture, 62(4):1266–1267, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/819791.

Jackson:2023:FNF

[Jac23a] Donald C. Jackson. Fixing Niagara Falls: Environment, Energy, and Engineers at the World’s Most Famous Waterfall by Daniel Macfarlane (review). Technology and Culture, 64(1):288–290, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
879618.

Jacobs:2023:DMA

[Jac23b] Robert (Bo) Jacobs. Discordant Memories: Atomic Age Narratives and Visual Culture by Alison Fields (review). Technology and Culture, 64(2):625–627, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893074.

Jamison:2022:MTB

[Jam22] Tommy Jamison. Manning the torpedo boats: How gendered insecurities shaped naval war in the United States and Britain, 1860– 1900. Technology and Culture, 63(4):1106–1136, October 2022.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/868055.

Jamison:2023:ISM

[Jam23] Tommy Jamison. Imperial Steam: Modernity on the Sea Route to India, 1837–74 by Jonathan Stafford (review). Technology and Culture, 64(4):1312–1314, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/911018.

Janiewski:2023:SCA

[Jan23a] Dolores E. Janiewski. Surveillance Capitalism in America ed. by Josh Lauer and Kenneth Lipartito (review). Technology and Culture, 64(1):230–232, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/879615.

Janneck:2023:EIZ

[Jan23b] Rouven Janneck. Ersatzstoffe im Zeitalter der Weltkriege: Geschichte, Bedeutung, Perspektiven ed. by Elisabeth Vaupel (review). Technology and Culture, 64(1):276–278, January 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/879629.

Johns:2020:OSM

[Joh20] Adrian Johns. Orders of service: Markus Krajewski, The Server. Technology and Culture, 61(2):682–685, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/761582.

John:2023:DNM

[Joh23] Richard R. John. Debating new media: Rewriting communications history. Technology and Culture, 64(2):308–358, April 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893040.

Jones:2022:ERG

[Jon22a] Geoffrey Jones. Engineering Rules: Global Standard Setting since 1880 by JoAnne Yates and Craig N. Murphy (review). Technology and Culture, 63(3):865–866, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/859733.

Jones:2022:TBS

[Jon22b] Maggie E. C. Jones. Traveling Black: A Story of Race and Resistance by Mia Bay (review). Technology and Culture, 63(3):869– 870, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
859735.

Jones:2023:RRP

[Jon23] Alisha Lola Jones. Race and Radio: Pioneering Black Broadcasters in New Orleans by Bala James Baptiste, and: Black Radio/ Black Resistance: The Life & Times of the Tom Joyner Morning Show by Micaela di Leonardo (review). Technology and Culture, 64(4):1352–1356, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/911039.

Jorgensen:2021:AUS

[Jør21] Finn Arne Jørgensen. Aluminum Upcycled: Sustainable Design in Historical Perspective by Carl A. Zimring (review). Technology and Culture, 62(4):1291–1292, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/819805.

Jansson:2023:ISS

[JR23] Måns Jansson and Göran Rydén. Improving Swedish steelmaking: Circulation and localized knowledge-making in early modernity. Technology and Culture, 64(2):515–542, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893045.

Judet:2022:NFM

[Jud22] Pierre Judet. Des nations, des firmes et des montres. Histoire globale de l’industrie horlogére de 1850 à nos jours [Countries, companies, and watches: A global history of the watch industry from 1850 to today] by Pierre-Yves Donzé (review). Technology and Culture, 63(2):591–592, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/854043.

Kaarlenkaski:2023:CHT

[Kaa23] Taija Kaarlenkaski. Cows and humans as technology users: Multispecies agency and gender in automated milking systems in Finland. Technology and Culture, 64(4):1093–1119, October 2023.

CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/910996.

Kagliwal:2020:VTH

[Kag20] Barkha Kagliwal. Visualizing Taste: How Business Changed the Look of What You Eat by Ai Hisano (review). Technology and Culture, 61(4):1224–1226, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/777628.

Kagliwal:2023:HTW

[Kag23] Barkha Kagliwal. A historian of technology watches “The King, his kitchen, and other stories”. Technology and Culture, 64(3):919– 929, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
903979.

Kaiserfeld:2021:HKP

[Kai21] Thomas Kaiserfeld. Histories of Knowledge in Postwar Scandinavia: Actors, Arenas, and Aspirations ed. by Johan Östling et al. (review). Technology and Culture, 62(3):912–913, July 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/803111.

Kaiserfeld:2023:BLF

[Kai23] Thomas Kaiserfeld. Beyond the Lab and the Field: Infrastructures as Places of Knowledge Production since the Late Nineteenth Century ed. by Eike-Christian Heine and Martin Meiske (review). Technology and Culture, 64(1):268–269, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/879609.

Kalman:2020:DRA

[Kal20] David Zvi Kalman. Delayed Response: The Art of Waiting from the Ancient to the Instant World by Jason Farman (review). Technology and Culture, 61(4):1226–1227, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/777630.

Kalinovsky:2023:TTS

[Kal23] Artemy Kalinovsky. Technology, Temporality and the Study of Central Asia ed. by Jonas van der Straeten and Julia Obertreis (review). Technology and Culture, 64(3):950–952, July 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/903990.

Kamp:2022:PDW

[Kam22] Marianne Kamp. Pipe Dreams: Water and Empire in Central Asia’s Aral Sea Basin by Maya K. Peterson (review). Technology and Culture, 63(4):1185–1186, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/868060.

Kane:2020:BSH

[Kan20] Carolyn Kane. Bright Signals: A History of Color Television by Susan Murray (review). Technology and Culture, 61(1):375–376, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
752956.

Kannan:2023:CCP

[Kan23a] Thamarai Selvan Kannan. Crafts and Craftsmen in Pre-colonial Eastern India: Technology and Culture by Asha Shukla Choubey (review). Technology and Culture, 64(1):263–264, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/879597.

Kannan:2023:SSC

[Kan23b] Thamarai Selvan Kannan. Skin and sound: Caring for and crafting bovine hide in South India. Technology and Culture, 64(4): 1121–1139, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/910997.

Kandilarov:2024:BCC

[Kan24] Evgeniy Kandilarov. Balkan Cyberia: Cold War Computing, Bulgarian Modernization, and the Information Age behind the Iron Curtain by Victor Petrov (review). Technology and Culture, 65 (2):739–740, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/926353.

Kapoor:2020:WPA

[Kap20] Nathan Kapoor. Wind and power in the Anthropocene: Cymene Howe, Ecologics and Dominic Boyer, Energopolitics. Technology and Culture, 61(2):686–689, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/761583.

Kardos:2023:SBM

[Kar23] Leah Kardos. Switched On: Bob Moog and the Synthesizer Revolution by Albert Glinsky (review). Technology and Culture, 64 (3):1003–1005, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/904019.

Kar:2024:HJB

[Kar24] Souvik Kar. Homi J. Bhabha: A Life by Bakhtiar K. Dadabhoy (review). Technology and Culture, 65(1):410–412, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/920557.

Kathke:2023:KRN

[Kat23] Torsten Kathke. Knowledge Regulation and National Security in Postwar America by Mario Daniels and John Krige (review). Technology and Culture, 64(2):636–637, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893080.

Kay:2020:WTH

[Kay20] Michael Kay. When They Hid the Fire: A History of Electricity and Invisible Energy in America by Daniel French (review). Technology and Culture, 61(4):1252–1254, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/777647.

Kay:2023:SWW

[Kay23] Michael Kay. Serving a Wired World: London’s Telecommunications Workers and the Making of an Information Capital by Katie Hindmarch-Watson (review). Technology and Culture, 64(3):994– 996, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
904014.

Kaufmann-Buhler:2022:DMP

[KB22a] Jennifer Kaufmann-Buhler. Designing for maintenance: Plant care technology in the Office. Technology and Culture, 63(4):1033–1056, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
868052.

Kaufmann-Buhler:2022:WEG

[KB22b] Jennifer Kaufmann-Buhler. Working in the electronic garden: The visual rhetoric of Herman Miller’s Action Office. Technology and Culture, 63(4):927–934, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/868047.

Kreysel:2024:GSF

[KdS24] Caroline Kreysel and Nathaly Yumi da Silva. Globalizing the Soybean: Fat, Feed, and Sometimes Food, c. 1900–1950 by Ines Prodöhl (review). Technology and Culture, 65(1):376–378, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920539.

Kessler:2021:TPV

[Kes21] Elizabeth A. Kessler. Technology’s palette: Voyager’s eyes and the hyperchromatic enhancement of Jupiter and Saturn. Technology and Culture, 62(4):1087–1118, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/836032.

Kifyasi:2024:HEW

[Kif24a] Andrea Azizi Kifyasi. How effective was the global south knowledge exchange? The Chinese-funded medical projects in Tanzania, 1968–1990s. Technology and Culture, 65(1):39–61, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/920515.

Kifyasi:2024:CSC

[Kif24b] Andrea Azizi Kifyasi. On the cover: Showcasing China’s on-the-job training in rural Africa. Technology and Culture, 65(1):1–5, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920513.

Kinkela:2023:DAA

[Kin23] David Kinkela. The Defoliation of America: Agent Orange Chemicals, Citizens, and Protests by Amy M. Hay (review). Technology and Culture, 64(1):282–283, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/879608.

Kirchhof:2022:ANB

[Kir22] Astrid Mignon Kirchhof. Atomgeschäfte. Die Nuklearexportpolitik der Bundesrepublik Deutschland 1970–1979 [Nuclear deals: The nuclear export policy of the Federal Republic of Germany 1970– 1979] by Daniel Romberg (review). Technology and Culture, 63 (2):546–547, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/854017.

Kirwan:2023:SCD

[Kir23] Adrian Kirwan. Shaping communications: The development of the National Telegraph Network in Ireland, 1850–70. Technology and Culture, 64(4):1185–1209, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/911000.

Kasperski:2023:WRN

[KJ23] Tatiana Kasperski and Paul Josephson. Women, reactors, and nuclear weapons: From revolutionary liberation to the “Miss atom” pageant in (post-)Soviet Russia. Technology and Culture, 64(3):791–822, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/903973.

Kline:2023:ANH

[Kli23] Ronald R. Kline. AI Narratives: A History of Imaginative Thinking about Intelligent Machines ed. by Stephen Cave, Kanta Dihal, and Sarah Dillon (review). Technology and Culture, 64(1):223–225, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
879596.

Kuzma-Markowska:2023:LDS

[KM23] Sylwia Kuźma-Markowska. Laboratory of Deficiency: Sterilization and Confinement in California, 1900–1950s by Natalie Lira (review). Technology and Culture, 64(1):295–297, January 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/879616.

Kalmbach:2020:CTF

[KMÅ20] Karena Kalmbach, Andreas Marklund, and Anna Åberg. Crises and technological futures: Experiences, emotion, and action. Technology and Culture, 61(1):272–281, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/752929.

Knowles:2020:DAI

[Kno20] Scott Gabriel Knowles. Documenting Aftermath: Information Infrastructures in the Wake of Disasters by Megan Finn (review). Technology and Culture, 61(2):704–706, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/761567.

Knoblauch:2022:AAC

[Kno22a] William M. Knoblauch. Atomic Americans: Citizens in a Nuclear State by Sarah E. Robey (review). Technology and Culture, 63 (4):1243–1244, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/868086.

Knowles:2022:MDC

[Kno22b] Scott Gabriel Knowles. The Meanings of a Disaster: Chernobyl and Its Afterlives in Britain and France by Karena Kalmbach (review). Technology and Culture, 63(3):906–907, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/859755.

Kobiljski:2022:EWD

[Kob22a] Aleksandra Kobiljski. Energy workarounds: Designing coals for the Japanese steel industry, 1895–1911. Technology and Culture, 63(2):326–348, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/854000.

Kobiljski:2022:GIN

[Kob22b] Aleksandra Kobiljski. The global industrial now. Technology and Culture, 63(2):309–325, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/853999.

Kobori:2022:WEE

[Kob22c] Satoru Kobori. When energy efficiency begets air pollution: Fuel conservation in Japan’s steel industry, 1945–60. Technology and Culture, 63(2):401–426, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/854003.

Kochetkova:2021:PPH

[Koc21] Elena Kochetkova. The Power of the Periphery: How Norway Became an Environmental Pioneer for the World by Peder Anker (review). Technology and Culture, 62(3):941–942, July 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/803126.

Kocurek:2023:HTP

[Koc23] Carly A. Kocurek. Hot Tubs and Pac-Man: Gender and the Early Video Game Industry in the United States (1950s–1980s) by Anne Ladyem McDivitt (review). Technology and Culture, 64 (4):1365–1367, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/911045.

Koerner:2023:WDW

[Koe23] Emily Rees Koerner. Why don’t we look at television? Technology and Culture, 64(4):1249–1259, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/911004.

Konagaya:2021:HIM

[Kon21] Daisuke Konagaya. How an Imperial military laboratory became a museum for peace. Technology and Culture, 62(2):551–560, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
794377.

Kosciejew:2021:IHC

[Kos21] Marc Kosciejew. Information: A Historical Companion ed. by Ann Blair et al. (review). Technology and Culture, 62(4):1276–1277, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
819796.

Kouli:2024:EOT

[Kou24] Yaman Kouli. European Objects: The Troubled Dreams of Harmonization by Brice Laurent (review). Technology and Culture, 65(2):746–747, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/926357.

Kranakis:2021:WTH

[Kra21a] Eda Kranakis. Writing technology into history. Technology and Culture, 62(1):212–240, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/783911.

Krasnodebski:2021:UFW

[Kra21b] Marcin Krasnodȩbski. Upscaling forest waste: The French quest for fuel autarky after World War I. Technology and Culture, 62 (1):105–127, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/783907.

Krapp:2022:WSD

[Kra22] Peter Krapp. Wie der Sozialstaat digital wurde: Die Computerisierung der Rentenversicherung im geteilten Deutschland [How the welfare state went digital: The computerization of pension insurance in divided Germany] by Thomas Kasper (review). Technology and Culture, 63(4):1209–1210, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/868065.

Krause:2023:DKC

[Kra23] Detlef Krause. Die Digitalisierung der Kreditwirtschaft: Computereinsatz in den Sparkassen der Bundesrepublik und der DDR 1957–1991 by Martin Schmitt (review). Technology and Culture, 64(1):225–227, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/879627.

Krebs:2021:RIM

[Kre21] Stefan Krebs. Repairing Infrastructures: The Maintenance of Materiality and Power by Christopher R. Henke and Benjamin Sims (review). Technology and Culture, 62(4):1256–1258, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/819785.

Kriger:2021:STT

[Kri21] Sarah Kriger. Shakespeare, Technicity, Theatre by W. B. Worthen, and: Shakespeare, Spectatorship, and the Technologies of Performance by Pascale Aebischer (review). Technology and Culture, 62 (2):652–654, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/794416.

Kriger:2024:TET

[Kri24] Sarah Kriger. The Theater of Electricity: Technology and Spectacle in the Late 19th Century by Ulf Otto (review). Technology and Culture, 65(2):737–738, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/926352.

Krupar:2020:DNL

[Kru20] Jason Krupar. The disappearing nuclear landscape: Snapshots of lost atomic technologies. Technology and Culture, 61(2):512–548, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
761574.

Krupar:2022:WAA

[Kru22] Jason Krupar. The Wretched Atom: America’s Global Gamble with Peaceful Nuclear Technology by Jacob Darwin Hamblin (review). Technology and Culture, 63(2):548–549, April 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/854018.

Kubot:2022:PZB

[Kub22] Tina Kubot. Mit der Post in die Zukunft: Der Bildschirmtext in der Bundesrepublik Deutschland von 1977 bis 2001 [Postal service into the future: Bildschirmtext in the Federal Republic of Germany, 1977–2001] by Hagen Schönrich (review). Technology and Culture, 63(4):1211–1212, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/868066.

Kucher:2021:IRM

[Kuc21] Michael Kucher. The Italian Renaissance of Machines by Paolo Galluzzi (review). Technology and Culture, 62(2):640–642, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
794396.

Kuitenbrouwer:2022:PFR

[Kui22] Vincent Kuitenbrouwer. Powerful Frequencies: Radio, State Power, and the Cold War in Angola, 1931–2002 by Marissa J. Moorman (review). Technology and Culture, 63(4):1217–1219, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
868071.

Kuo:2023:SIP

[Kuo23] Laureen Kuo. The search for independence: Planning for France’s information technology autonomy, 1958–66. Technology and Culture, 64(2):485–514, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893044.

Kushner:2023:CCT

[Kus23a] Scott Kushner. Controlling crowds: On the technological management of entertainment audiences. Technology and Culture, 64 (1):7–33, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/879580.

Kushner:2023:CER

[Kus23b] Scott Kushner. Cover essay: Reading ordinary photographs. Technology and Culture, 64(1):1–6, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/879633.

Kvicalova:2023:TSN

[Kvi23a] Anna Kvicalova. Thinking with Sound: a New Program in the Sciences and Humanities around 1900 by Viktoria Tkaczyk (review). Technology and Culture, 64(4):1341–1343, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/911033.

Kvicalova:2023:SQS

[Kvi23b] Anna Kvicalova. Sound on the quiet: Speaker identification and auditory objectivity in Czechoslovak fonoscopy, 1975–90. Technology and Culture, 64(2):379–406, April 2023. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893089.

Labbato:2021:AMG

[Lab21] Maria Labbato. Anticoncepción, mujeres y género: La ‘píldora’ en España y Polonia (1960–1980) [Contraception, women, and gender: The ‘pill’ in Spain and Poland (1960–1980)] by Agata Ignaciuk and Teresa Ortiz Gómez (review). Technology and Culture, 62(3):906–908, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/803146.

Ladd:2022:AIP

[Lad22] Brian Ladd. Autonorama: The Illusory Promise of High-Tech Driving by Peter Norton (review). Technology and Culture, 63 (4):1196–1197, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/868098.

Lagendijk:2024:PPE

[Lag24] Vincent Lagendijk. The Politics and Policies of European Economic Integration, 1850–1914 by Yaman Kouli and Léonard Laborie (review). Technology and Culture, 65(1):392–393, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920547.

Lallani:2021:FAT

[Lal21a] Shayan S. Lallani. Food and Aviation in the Twentieth Century: The Pan American Ideal by Bryce Evans (review). Technology and Culture, 62(4):1293–1294, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/819806.

Laloux:2021:CCL

[Lal21b] Ludovic Laloux. Caves et celliers dans l’Europe médiévale et moderne [Caverns and cellars in medieval and modern Europe] ed. by Clément Alix et al. (review). Technology and Culture, 62(4): 1252–1254, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/819783.

Lang:2023:SRI

[Lan23a] Claudia Lang. Science and Religion in India: Beyond Disenchantment by Renny Thomas (review). Technology and Culture, 64(3):976–978, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/904004.

Lanzillo:2023:ICA

[Lan23b] Amanda Lanzillo. Industrial Craft in Australia: Oral Histories of Creativity and Survival by Jesse Adams Stein (review). Technology and Culture, 64(3):968–970, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/904000.

Lanzillo:2024:PPC

[Lan24] Amanda M. Lanzillo. Prison papermaking: Colonial ideals of industrial experimentation in India. Technology and Culture, 65 (1):63–87, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/920516.

Lassman:2023:FCF

[Las23] Thomas C. Lassman. Flying Camelot: The F-15, the F-16, and the Weaponization of Fighter Pilot Nostalgia by Michael W. Hankins (review). Technology and Culture, 64(1):250–252, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/879607.

Launius:2021:LAL

[Lau21] Roger D. Launius. Lady Astronauts, Lady Engineers, and Naked Ladies: Women and the American Space Community During the Cold War, 1960s–1980s by Karin Hilck (review). Technology and Culture, 62(1):299–300, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/783939.

Launius:2023:BSS

[Lau23] Roger D. Launius. Beyond Sputnik and the Space Race: The Origins of Global Satellite Communications by Hugh R. Slotten (review). Technology and Culture, 64(2):630–632, April 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893077.

Lavelle:2021:REC

[Lav21] Peter Lavelle. Remains of the Everyday: A Century of Recycling in Beijing by Joshua Goldstein (review). Technology and Culture, 62(4):1233–1234, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/819801.

Roux:2023:CIT

[Le 23] Ronan Le Roux. Code: From Information Theory to French Theory by Bernard D. Geoghegan (review). Technology and Culture, 64(4):1364–1365, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/911044.

Lee:2024:WSG

[Lee24] Youngju Lee. Women in the Sky: Gender and Labor in the Making of Modern Korea by Hwasook Nam (review). Technology and Culture, 65(1):363–365, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/920532.

Leslie:2020:WCT

[Les20] Christopher Leslie. As we could have thought: Deploying historical narratives of the Memex in support of innovation. Technology and Culture, 61(2):480–511, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/761573.

Leslie:2024:MSD

[Les24] Stuart W. Leslie. “Modernism with a soul”: Designing and building communities for corporate and civic life. Technology and Culture, 65(1):343–357, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/920528.

Levitt:2020:WLB

[Lev20] Theresa Levitt. When lighthouses became public goods: The role of technological change. Technology and Culture, 61(1):144–172, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
752963.

Levings:2021:FTC

[Lev21] Colin Levings. The fishery technology complex: From mapping to depletion of Pacific Ocean perch, 1880s–1970s. Technology and Culture, 62(1):185–211, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/783910.

Li:2022:EPC

[Li22] Xiaoyue Li. Electrical Palestine: Capital and Technology from Empire to Nation by Fredrik Meiton (review). Technology and Culture, 63(2):563–565, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/854027.

Li:2024:FHR

[Li24] Yujie Li. Fixing the Huai River: Technology of labor formation in Maoist China, 1950–53. Technology and Culture, 65(1):177–209, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920520.

Liebhold:2023:CME

[Lie23] Peter Liebhold. Chronicles of Mechanical Engineering in the United States ed. by Thomas H. Fehring and Terry S. Reynolds (review). Technology and Culture, 64(2):649–650, April 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893087.

Lifset:2023:ESH

[Lif23] Robert Lifset. The Extraction State: A History of Natural Gas in America by Charles Blanchard (review). Technology and Culture, 64(3):991–992, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/904012.

Lintsen:2021:AHC

[Lin21] Harry Lintsen. The Anthropocene and the Humanities: From Climate Change to a New Age of Sustainability by Carolyn Merchant (review). Technology and Culture, 62(2):604–606, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794405.

Lintsen:2022:CEH

[Lin22] Harry Lintsen. The Contamination of the Earth. A History of Pollutions in the Industrial Age by François Jarrige and Thomas le Roux (review). Technology and Culture, 63(2):544–545, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/854016.

Linkis:2024:SBA

[Lin24a] Sara Tanderup Linkis. Soundtracked Books from the Acoustic Era to the Digital Age: A Century of “Books That Sing” by Justin St. Clair (review). Technology and Culture, 65(1):422–423, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920563.

Lintsen:2024:TTV

[Lin24b] Harry Lintsen. Technici en de totalitaire verleiding: Het Technisch Gilde en de ’volkse gemeenschap’ [Technicians and the totalitarian temptation: The Technical Guild and the “people’s community”] by Hans Schippers (review). Technology and Culture, 65(1):397–399, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/920550.

Lipton:2022:SUA

[Lip22] Miriam F. Lipton. The Soviet Union’s Agricultural Biowarfare Programme: Ploughshares to Swords by Anthony Rimmington (review). Technology and Culture, 63(2):568–570, April 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/854030.

Lees-Maffei:2020:SRL

[LM20] Grace Lees-Maffei. Streamliner: Raymond Loewy and Image-Making in the Age of American Industrial Design by John Wall (review). Technology and Culture, 61(2):699–700, April 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/761564.

Loeb:2024:AIA

[Loe24] Zachary Loeb. The Apple II Age: How the Computer Became Personal by Laine Nooney (review). Technology and Culture, 65 (1):418–420, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/920561.

Loomis:2022:TSR

[Loo22] Erik Loomis. Timber, Sail, and Rail: An Archaeology of Industry, Immigration, and the Loma Prieta Mill by Marco G. Meniketti (review). Technology and Culture, 63(4):1199–1201, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/868100.

Lopez:2022:NAD

[Lóp22] Santiago M. López. Del noventayochismo al desarrollismo. Técnica e ingeniería en España VIII [From 1898 to Franco’s developmentalism: Technology and engineering in Spain VIII] ed. by Manuel Silva Suárez (review). Technology and Culture, 63 (2):556–557, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/854022.

Lopez:2024:SAS

[Lóp24] Diego Arango López. El sol al servicio de la humanidad: Historia de la energía solar en Chile [The sun in the service of humanity: History of solar energy in Chile] ed. by Mauricio Osses, Cecilia Ibarra, and Bárbara Silva (review). Technology and Culture, 65(2):677–679, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/926322.

Lourdusamy:2021:TWH

[Lou21] John B. Lourdusamy. Tea War: A History of Capitalism in China and India by Andrew B. Liu (review). Technology and Culture, 62(4):1231–1232, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/819790.

Lovheim:2023:CML

[Löv23] Daniel Lövheim. The Charisma Machine: The Life, Death, and Legacy of One Laptop per Child by Morgan G. Ames (review). Technology and Culture, 64(2):608–609, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893065.

Low:2020:BED

[Low20] Morris Low. Behind the Exhibit: Displaying Science and Technology at World’s Fairs and Museums in the Twentieth Century. Artefacts: Studies in the History of Science and Technology, vol. 12 ed. by Elena Canadelli (review). Technology and Culture, 61 (4):1269–1270, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/777656.

Low:2023:DSE

[Low23] Morris Low. Dream Super-Express: A Cultural History of the World’s First Bullet Train by Jessamyn R. Abel (review). Technology and Culture, 64(1):269–271, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/879588.

Lozano:2022:CDR

[Loz22] Jennifer M. Lozano. Cuba’s Digital Revolution: Citizen Innovation and State Policy ed. by Ted A. Henken and Sara Garcia Santamaria (review). Technology and Culture, 63(4):1205–1206, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
868063.

Lotysz:2020:TPI

[ŁPO^{+}20] Sławomir Łotysz, Ciro Paoletti, Glen O’Sullivan, Kamna Tiwary, and Magdalena Zdrodowska. “Technology and Power”: The International Committee for the History of Technology’s (ICOHTEC), University of Silesia, Katowice, Poland (July 22–27, 2019). Technology and Culture, 61(4):1188–1195, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/777613.

Lucas:2020:FYM

[Luc20a] Adam Lucas. Fifty Years of Medieval Technology and Social Change: AVISTA Studies in the History of Medieval Technology, Science, and Art ed. by Steven A. Walton (review). Technology and Culture, 61(4):1214–1215, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/777619.

Lucas:2020:MMI

[Luc20b] Adam Lucas. Manufacturing modernity: Innovations in early modern Europe — an introduction. Technology and Culture, 61 (4):995–1016, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/777600.

Ludwig:2023:DPR

[Lud23] Jason Ludwig. Digitize and Punish: Racial Criminalization in the Digital Age by Brian Jefferson (review). Technology and Culture, 64(4):1356–1358, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/911040.

Luesink:2021:GMC

[Lue21] David Luesink. Global Medicine in China: A Diasporic History by Wayne Soon (review). Technology and Culture, 62(4):1229–1230, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
819779.

Lundin:2022:GAD

[Lun22] Per Lundin. Greening the Alliance: The Diplomacy of NATO’s Science and Environmental Initiatives by Simone Turchetti (review). Technology and Culture, 63(1):276–278, January 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/844193.

Lundberg:2023:PBP

[Lun23] Christa Lundberg. Papal Bull: Print, Politics, and Propaganda in Renaissance Rome by Margaret Meserve (review). Technology and Culture, 64(2):606–607, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893064.

Lutz:2020:PPH

[Lut20] Raechel Lutz. Petroleum’s park: How oil shaped the Palisades Interstate Park, 1900–1960. Technology and Culture, 61(3):713– 738, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
763578.

Macfarlane:2020:NEH

[Mac20] Daniel Macfarlane. Nature empowered: Hydraulic models and the engineering of Niagara falls. Technology and Culture, 61(1): 109–143, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/752962.

Macedo:2022:RJG

[Mac22a] Marta Macedo. Rio de Janeiro in the Global Meat Market, c. 1850 to c. 1930: How Fresh and Salted Meat Arrived at the Carioca Table by Maria-Aparecida Lopes (review). Technology and Culture, 63(3):897–898, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/859750.

Macrakis:2022:VVW

[Mac22b] Kristie Macrakis. Versagtes Vertrauen: Wissenschaftler der DDR im Visier der Staatssicherheit [Betrayed trust: GDR scientists as targets of state security] by Reinhard Buthmann (review). Technology and Culture, 63(2):558–559, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/854023.

Macmillan:2023:SPR

[Mac23] Thomas Macmillan. Shredding Paper: The Rise and Fall of Maine’s Mighty Paper Industry by Michael G. Hillard (review). Technology and Culture, 64(1):278–279, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/879610.

Mackinnon:2024:VGT

[Mac24] Katie Mackinnon. Van Gogh TV’s “Piazza Virtuale”: The Invention of Social Media at documenta IX in 1992 by Tilman Baumgärtel (review). Technology and Culture, 65(1):420–422, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920562.

Maines:2023:GCS

[Mai23] Rachel Maines. Gendered Capitalism: Sewing Machines and Multinational Business in Spain and Mexico, 1850–1940 by Paula A. de la Cruz–Fernández (review). Technology and Culture, 64 (4):1303–1305, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/911013.

Maines:2024:FCU

[Mai24] Rachel Maines. Feminism and capitalism under the nuclear cloud & Barbie. Technology and Culture, 65(1):333–342, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/920527.

Mak:2023:WPG

[Mak23] Bonnie Mak. Working with Paper: Gendered Practices in the History of Knowledge ed. by Carla Bittel, Elaine Leong, and Christine von Oertzen (review). Technology and Culture, 64(1):261–262, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
879591.

Mallea:2024:WAC

[Mal24] Amahia Mallea. Water and Agriculture in Colorado and the American West: First in Line for the Rio Grande by David Stiller (review). Technology and Culture, 65(2):722–723, April 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/926344.

Manoff:2020:SBE

[Man20] Marlene Manoff. The Stuff of Bits: An Essay on the Materialities of Information by Paul Dourish (review). Technology and Culture, 61(3):955–956, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/763602.

Mapunda:2023:DWS

[Map23] Bertram Mapunda. Debate: Why study precolonial African technology and material culture? Technology and Culture, 64(3):665– 676, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
903968.

Marfella:2021:DTA

[Mar21a] Giorgio Marfella. Design Technics: Archaeologies of Architectural Practice ed. by Zeynep Çelik Alexander and John May (review). Technology and Culture, 62(2):630–631, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794382.

Mari:2021:ITH

[Mar21b] Will Mari. If Then: How the Simulmatics Corporation Invented the Future by Jill Lepore (review). Technology and Culture, 62 (2):655–656, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/794403.

Marshall:2021:MMM

[Mar21c] Owen Marshall. The maniac-making machine: A media history of delayed auditory feedback. Technology and Culture, 62(3):839– 860, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
803107.

Markley:2022:DFP

[Mar22a] Robert Markley. The Draining of the Fens: Projectors, Popular Politics, and State Building in Early Modern England by Eric H. Ash (review). Technology and Culture, 63(3):857–858, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/859729.

Marsh:2022:FOU

[Mar22b] Allison Marsh. Fellows online: The user experience. Technology and Culture, 63(4):1157–1167, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/868058.

Marsh:2023:ECO

[Mar23a] Allison Marsh. Extinct: A Compendium of Obsolete Objects ed. by Barbara Penner et al. (review). Technology and Culture, 64 (2):580–582, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893050.

Martland:2023:HRP

[Mar23b] Samuel J. Martland. Hungry for Revolution: The Politics of Food and the Making of Modern Chile by Joshua Frens-String (review). Technology and Culture, 64(3):1013–1014, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/904024.

Marrec:2024:LMH

[Mar24] Anaël Marrec. L’énergie des marées: Hier, aujourd’hui, demain ed. by Ewan Sonnic (review). Technology and Culture, 65(1): 382–384, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/920542.

Mas:2020:CCH

[Mas20] Catherine Mas. Contraception: A Concise History by Donna J. Drucker (review). Technology and Culture, 61(3):949–951, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
763603.

Mateos:2021:PCT

[Mat21] Gisela Mateos. Políticas de ciencia, tecnología e innovación en la Argentina de la posdictadura [Science, technology and innovation policies in post-dictatorship Argentina] ed. by Diego Aguiar et al. (review). Technology and Culture, 62(1):321–323, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/783952.

Mateos:2024:PNI

[Mat24] Gisela Mateos. El pequeño intervencionista: Ayuntamiento, empresarios y regulación del mercado telefónico en la Ciudad de México, 1881–1915 [The small interventionist: City hall, businessmen and the regulation of the telephone market in Mexico City, 1881–1915] by Víctor Cuchí Espada (review). Technology and Culture, 65(2):691–693, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/926329.

McCray:2020:TIA

[McC20] W. Patrick McCray. Technocrats of the Imagination: Art, Technology, and the Military-Industrial Avant-Garde by John Beck and Ryan Bishop (review). Technology and Culture, 61(3):986–988, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
763597.

McCrossen:2021:THW

[McC21] Alexis McCrossen. On Time: A History of Western Timekeeping by Ken Mondschein (review). Technology and Culture, 62 (2):599–600, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/794407.

McCray:2022:AOJ

[McC22] W. Patrick McCray. Art out of order: Jack Burnham, the 1970 software show, and the aesthetics of information systems. Technology and Culture, 63(3):689–717, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/859721.

McCray:2023:APF

[McC23] W. Patrick McCray. Arte Programmata: Freedom, Control, and the Computer in 1960s Italy by Lindsay Caplan (review). Technology and Culture, 64(2):609–611, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893066.

McDivitt:2022:HGB

[McD22] Anne Ladyem McDivitt. Homebrew Gaming and the Beginnings of Vernacular Digitality by Melanie Swalwell (review). Technology and Culture, 63(4):1212–1214, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/868067.

Mcguire:2020:MEH

[Mcg20] Coreen Mcguire. Managing the Experience of Hearing Loss in Britain: 1830–1930 by Graeme Gooday and Karen Sayer (review). Technology and Culture, 61(4):1227–1229, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/777632.

Mcgovern:2021:CRR

[Mcg21] Michael F. Mcgovern. Calculating Race: Racial Discrimination in Risk Assessment by Benjamin Wiggins (review). Technology and Culture, 62(3):894–896, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/803110.

McGuire:2024:WIP

[McG24] Coreen McGuire. Who invented the possum? What historians can learn from disabled innovation in Britain’s responaut communities. Technology and Culture, 65(1):89–116, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/920517.

Mchome:2024:ALE

[Mch24] Emanuel Lukio Mchome. Apartheid’s Leviathan: Electricity and the Power of Technological Ambivalence by Faeeza Ballim (review). Technology and Culture, 65(1):380–382, January 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/920541.

McNeil:2024:TNL

[McN24] Peter McNeil. Technology, Novelty, and Luxury ed. by Artemis Yagou (review). Technology and Culture, 65(1):358–359, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920529.

Meadow:2021:MMS

[Mea21] Mark A. Meadow. Making Marvels: Science and Splendor at the Courts of Europe ed. by Wolfram Koeppe (review). Technology and Culture, 62(4):1241–1242, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/819811.

Mei:2021:KGJ

[Mei21] Jianjun Mei. Kao Gong Ji: The World’s Oldest Encyclopaedia of Technologies trans. by Guan Zengjian and Konrad Herrmann (review). Technology and Culture, 62(2):591–593, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794417.

Meindl:2022:SSC

[Mei22a] Michael J. Meindl. Sound Streams: A Cultural History of Radio-Internet Convergence by Andrew J. Bottomley (review). Technology and Culture, 63(1):293–294, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/844202.

Meister:2022:MVH

[Mei22b] Anna-Maria Meister. Modernism’s Visible Hand: Architecture and Regulation in America by Michael Osman (review). Technology and Culture, 63(1):297–298, January 2022. CODEN TECUA3.

ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/844204.

Meikle:2023:PPH

[Mei23] Jeffrey L. Meikle. Plasticidade: Plasticity — a History of Plastics in Portugal ed. by Maria Elvira Callapez, Sara Marques da Cruz, and Vânia Carvalho (review). Technology and Culture, 64 (4):1305–1306, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/911014.

Melzer:2023:AMM

[Mel23] Jürgen P. Melzer. The Aeroplane and the Making of Modern India by Aashique Ahmed Iqbal, and: Hong Kong Takes Flight: Commercial Aviation and the Making of a Global Hub, 1930s–1998 by John D. Wong (review). Technology and Culture, 64(3):952–954, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
903991.

Meyer:2020:BSF

[Mey20] Jan-Henrik Meyer. Big Science Fiction: Kernfusion und Popkultur in den USA by Simon Märkl (review). Technology and Culture, 61(3):983–984, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/763612.

Mannisto-Funk:2020:SSS

[MF20] Tiina Männistö-Funk. Straßenverkehr und soziale Sichtbarkeit: Das Massenmedium Straße in Chicago 1900–1930 by David Sittler (review). Technology and Culture, 61(3):962–963, July 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/763618.

Michel:2021:EUE

[Mic21] Alain P. Michel. Engineers, unity, expansion, fragmentation (19th and 20th centuries), volume 1: The production of a social group ed. by Antoine Derouet and Simon Paye (review). Technology and Culture, 62(2):638–640, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/794389.

Mika:2020:HLR

[Mik20] Marissa Mika. The half-life of radiotherapy and other transferred technologies. Technology and Culture, 61(2S):S135–S157, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
761589.

Milan:2020:USS

[Mil20a] Stefania Milan. Undone Science: Social Movements, Mobilized Publics, and Industrial Transitions by David J. Hess (review). Technology and Culture, 61(1):371–373, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/752944.

Miller:2020:NPN

[Mil20b] David Philip Miller. A new perspective on the natural philosophy of steams and its relation to the steam engine. Technology and Culture, 61(4):1129–1148, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/777607.

Milan:2022:HTT

[Mil22] Stefania Milan. Hacked Transmissions: Technology and Connective Activism in Italy by Alessandra Renzi (review). Technology and Culture, 63(4):1207–1208, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/868064.

Miller:2023:DPM

[Mil23a] Jonson Miller. Delta of Power: The Military-Industrial Complex by Alex Roland (review). Technology and Culture, 64(2):632–634, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893078.

Milligan:2023:MWP

[Mil23b] Ian Milligan. The Modem World: A Prehistory of Social Media by Kevin Driscoll (review). Technology and Culture, 64(3):1000–1001, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
904017.

Milor:2024:ETD

[Mil24] Alice Milor. European technical democracy? How consumer associations did — or did not — shape the 1980s automotive environmental standards. Technology and Culture, 65(1):265–291, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920523.

Mitsuda:2022:CPA

[Mit22] Tatsuya Mitsuda. Communist Pigs: An Animal History of East Germany’s Rise and Fall by Thomas Fleischman (review). Technology and Culture, 63(2):525–527, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/854024.

Mitcham:2023:SAB

[Mit23] Carl Mitcham. Science and Apocalypse in Bertrand Russell: A Cultural Sociology by Javier Pérez–Jara and Lino Camprubí (review). Technology and Culture, 64(2):634–636, April 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893079.

Mitchell:2024:SEC

[Mit24] M. X. Mitchell. Suburban Empire: Cold War Militarization in the US Pacific by Lauren Hirshberg (review). Technology and Culture, 65(1):390–392, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/920546.

Miyamoto:2022:VNP

[Miy22] Yuki Miyamoto. Visualizing Nuclear Power in Japan: A Trip to the Reactor by Morris Low (review). Technology and Culture, 63(3):908–909, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/859756.

Molella:2023:AAP

[MK23] Arthur Molella and Robert Kargon. Atomville: Architects, planners, and how to survive the Bomb. Technology and Culture, 64(3):823–844, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/903974.

Mamidipudi:2022:CCH

[MO22] Annapurna Mamidipudi and Vivek S. Oak. The Crafts and Capitalism: Handloom Weaving Industry in Colonial India by Tirthankar Roy (review). Technology and Culture, 63(1):259–261, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
844184.

Mody:2021:STP

[Mod21] Cyrus C. M. Mody. Steps toward a Philosophy of Engineering: Historico-Philosophical and Critical Essays by Carl Mitcham (review). Technology and Culture, 62(2):602–604, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794406.

Molella:2022:ST

[Mol22] Arthur Molella. The soul of technology. Technology and Culture, 63(2):494–506, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/854011.

Mondelli:2022:BSB

[Mon22] Frank Mondelli. Beautiful sounds, beautiful life: Cultivating musical listening through hearing aids in 1950s Japan. Technology and Culture, 63(4):1057–1077, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/868053.

Montano:2023:MLP

[Mon23] Diana J. Montaño. Missionaries of light and progress in Mexico: Engineers and technological pilgrims craft necaxa falls, 1890s– 1914. Technology and Culture, 64(3):677–705, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/903969.

Moorman:2020:RSS

[Moo20] Marissa J. Moorman. Radio Soundings: South Africa and the Black Modern by Liz Gunner (review). Technology and Culture, 61(4):1254–1256, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/777648.

Moore:2021:ELA

[Moo21] Abigail Harrison Moore. Electric Light: An Architectural History by Sandy Isenstadt (review). Technology and Culture, 62 (1):306–307, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/783943.

Moon:2023:TNH

[Moo23] Suzanne Moon. Taking up neglected histories. Technology and Culture, 64(2):543–549, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893046.

Moraglio:2022:IRS

[Mor22] Massimo Moraglio. The Italian Railways: 1839–2019 by Stefano Maggi (review). Technology and Culture, 63(3):881–883, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/859742.

Morgan:2023:TAH

[Mor23] Daniel Patrick Morgan. Technical Arts in the Han Histories: Tables and Treatises in the Shiji and Hanshu ed. by Mark Csikszentmihalyi and Michael Nylan (review). Technology and Culture, 64 (4):1292–1293, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/911007.

Morgan:2024:ERW

[Mor24] Njogu Morgan. Everyday resistance to white supremacy: Walking and cycling while black in Springs, South Africa, 1950s–1970s. Technology and Culture, 65(2):473–495, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/926312.

Murray:2021:HDM

[MS21] John E. Murray and Javier Silvestre. How do mines explode? Understanding risk in European mining doctrine, 1803–1906. Technology and Culture, 62(3):780–811, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/803105.

Moon:2020:PB

[MSH20] Suzanne Moon, Peter Soppelsa, and Hunter Heyck. Passing the baton. Technology and Culture, 61(2):709–712, April 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/761569.

Mueller:2022:PPH

[Mue22] Lucas M. Mueller. Pyrrhic Progress: The History of Antibiotics in Anglo-American Food Production by Claas Kirchhelle (review). Technology and Culture, 63(3):900–902, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/859752.

Mukhopadhyay:2021:EEI

[Muk21] Aparajita Mukhopadhyay. The English East India Company’s Silk Enterprise in Bengal, 1750–1850: Economy, Empire and Business by Karolina Hutková (review). Technology and Culture, 62 (2):593–595, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/794401.

Mukhopadhyay:2022:IAE

[Muk22] Aparajita Mukhopadhyay. Installing Automobility: Emerging Politics of Mobility and Streets in Indian Cities by Govind Gopakumar (review). Technology and Culture, 63(3):876–877, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/859739.

Mukhopadhyay:2024:HST

[Muk24] Aparajita Mukhopadhyay. History of Science, Technology, Environment, and Medicine in India ed. by Suvobrata Sarkar (review). Technology and Culture, 65(1):409–410, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/920556.

Mullally:2021:EPW

[Mul21] Sasha Mullally. Easing Pain on the Western Front: American Nurses of the Great War and the Birth of Modern Nursing Practice by Paul E. Stepansky (review). Technology and Culture, 62 (2):620–621, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/794414.

Munday:2020:MBI

[Mun20] Pat Munday. Mining the Borderlands: Industry, Capital, and the Emergence of Engineers in the Southwest Territories by Sarah E. M. Grossman (review). Technology and Culture, 61(1):351–352, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
752943.

Mundy:2023:LFR

[Mun23] Rachel Mundy. Listening in the Field: Recording and the Science of Birdsong by Joeri Bruyninckx (review). Technology and Culture, 64(1):243–245, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/879593.

Murray:2020:NSA

[Mur20] Susan Murray. The new surgical amphitheater: Color television and medical education in postwar America. Technology and Culture, 61(3):772–797, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/763580.

Nagy:2021:PCN

[Nag21] Jeff Nagy. Pink chat: Networked sex work before the Internet. Technology and Culture, 62(1):57–81, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/783905.

Nair:2022:DTA

[Nai22] Aparna Nair. Doctoring Traditions: Ayurveda, Small Technologies, and Braided Sciences by Projit Bihari Mukharji (review). Technology and Culture, 63(3):861–863, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/859731.

Naquin:2022:MPM

[Naq22] Susan Naquin. Making the Palace Machine Work: Mobilizing People, Objects, and Nature in the Qing Empire ed. by Martina Siebert, Kai Jun Chen, and Dorothy Ko (review). Technology and Culture, 63(4):1256–1258, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/868094.

Negre:2022:PTA

[Nèg22a] Valérie Nègre. Penser la technique autrement, XVI^{e}–XXI^{e} siècle. Hommage à l’oeuvre d’Hélène Vérin [New ways of thinking about technology, 16th–21st centuries: A tribute to the work of Hélène Vérin] ed. by Robert Carvais, Anne-Françoise Garçon, and André Grelon (review). Technology and Culture, 63(2):531–532, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
854048.

Negri:2022:CFE

[Neg22b] Sabrina Negri. The Cinema in Flux: The Evolution of Motion Picture Technology from the Magic Lantern to the Digital Era by Lenny Lipton (review). Technology and Culture, 63(4):1219–1221, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
868072.

Nelson:2022:HIT

[Nel22] Sarah Nelson. The History of the International Telecommunication Union: Transnational Techno-Diplomacy from the Telegraph to the Internet ed. by Gabriele Balbi and Andreas Fickers (review). Technology and Culture, 63(2):578–579, April 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/854036.

Nieto-Galan:2020:FCM

[NG20] Agustí Nieto-Galan. Fabriano: City of Medieval and Renaissance Papermaking by Sylvia Rodgers Albro (review). Technology and Culture, 61(2):690–691, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/761584.

Nieto-Galan:2021:BBF

[NG21] Agustí Nieto-Galan. Berliner Blau: Vom frühneuzeitlichen Pigment zum modernen Hightech-Material [Prussian blue: From early modern pigment to modern high-tech material] by Alexander Kraft (review). Technology and Culture, 62(2):644–645, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794418.

Nishiyama:2023:WRS

[Nis23] Takashi Nishiyama. Wings for the Rising Sun: a Transnational History of Japanese Aviation by Jürgen P. Melzer (review). Technology and Culture, 64(4):1337–1338, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/911031.

Nishiyama:2020:MAM

[NOB^{+}20] Takashi Nishiyama, Lisa Onaga, Philip Brown, John DiMoia, Yulia Frumer, Christopher Jones, Hiromi Mizuno, Ying Jia Tan, Togo Tsukuhara, and Nobuhiro Yamane. In memoriam: Aaron S. Moore (1972–2019). Technology and Culture, 61(2):678–681, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
761581.

Nakamura:2022:LRT

[NOK22] Shinichiro Nakamura, Taikan Oki, and Shinjiro Kanae. Lost rivers: Tokyo’s sewage problem in the high-growth period, 1953–73. Technology and Culture, 63(2):427–449, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/854004.

Nolt:2023:VAP

[Nol23] Steven M. Nolt. Virtually Amish: Preserving Community at the Internet’s Margins by Lindsay Ems (review). Technology and Culture, 64(1):227–228, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/879602.

Norton:2020:UTM

[Nor20] Peter Norton. Urban transport and mobility in technology and culture. Technology and Culture, 61(4):1197–1211, October 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/777617.

Norton:2021:POR

[Nor21] Peter Norton. Policing the Open Road: How Cars Transformed American Freedom by Sarah A. Seo (review). Technology and Culture, 62(2):614–615, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/794413.

Norton:2023:CLW

[Nor23] Peter Norton. Consuming Landscapes: What We See When We Drive and Why It Matters by Thomas Zeller (review). Technology and Culture, 64(4):1333–1335, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/911029.

Novick:2023:IBR

[Nov23a] Tamar Novick. Introducing bovine regimes: When animals become technologies. Technology and Culture, 64(4):1027–1043, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
910993.

Novick:2023:CSV

[Nov23b] Tamar Novick. On the cover: Speculations with vaginal specula. Technology and Culture, 64(4):1019–1026, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/910992.

Nye:2020:NTJ

[Nye20a] David E. Nye. Narratives of Technology by J. M. van der Laan (review). Technology and Culture, 61(1):344–345, January 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/752948.

Nye:2020:TCH

[Nye20b] David E. Nye. Technology: Critical History of a Concept by Eric Schatzberg (review). Technology and Culture, 61(4):1212–1213, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
777618.

Nye:2021:HWH

[Nye21] David E. Nye. Harari’s world history: Evolution toward intelligence without consciousness? Technology and Culture, 62(4): 1219–1228, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/819778.

Osseo-Asare:2023:ADC

[OA23] Abena Dove Osseo-Asare. Atomic Doctors: Conscience and Complicity at the Dawn of the Nuclear Age by James L. Nolan, Jr.

(review). Technology and Culture, 64(2):620–622, April 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893072.

Ochigame:2023:RSJ

[Och23] Rodrigo Ochigame. Robo sapiens japanicus: Robots, Gender, Family, and the Japanese Nation by Jennifer Robertson (review). Technology and Culture, 64(2):591–593, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893056.

ODonnell:2020:ENF

[O’D20] Casey O’Donnell. I Am Error: The Nintendo Family Computer by Nathan Altice (review). Technology and Culture, 61(1):378–380, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
752937.

ODonnell:2022:AAE

[O’D22] Casey O’Donnell. Atari Age: The Emergence of Video Games in America by Michael Z. Newman (review). Technology and Culture, 63(4):1214–1215, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/868068.

Oldenziel:2020:E

[Old20] Ruth Oldenziel. Editorial. Technology and Culture, 61(3):ix–xii, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
763604.

Oldenziel:2024:PHI

[Old24] Ruth Oldenziel. Public history: Introducing Barbenheimer. Technology and Culture, 65(1):315–317, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/920525.

Oppenheim:2022:SBT

[Opp22] Lois Oppenheim. Samuel Beckett and Technology ed. by Galina Kiryushina, Einat Adar, and Mark Nixon (review). Technology and Culture, 63(4):1225–1226, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/868075.

OSullivan:2020:GVA

[O’S20] Robin O’Sullivan. Garden Variety: The American Tomato from Corporate to Heirloom by John Hoenig (review). Technology and Culture, 61(1):359–360, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/752945.

Osorio:2022:WCN

[Oso22] Alejandra Osorio. Why Chuño matters: Rethinking the history of technology in Latin America. Technology and Culture, 63(3):808– 829, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
859725.

Ott:2021:TNL

[Ott21a] Cindy Ott. Through a Native Lens: American Indian Photography by Nicole Dawn Strathman (review). Technology and Culture, 62(3):898–899, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/803132.

Ott:2021:SHD

[Ott21b] Cindy Ott. Seeing history in 2-d: A tool kit for interpreting images. Technology and Culture, 62(4):1199–1216, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/819776.

Otter:2022:ST

[Ott22] Chris Otter. Socializing the technosphere. Technology and Culture, 63(4):953–978, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/868049.

Overy:2022:ANM

[Ove22] Richard Overy. The Anarchy of Nazi Memorabilia: From Things of Tyranny to Troubled Treasure by Michael Hughes (review). Technology and Culture, 63(3):863–865, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/859732.

Oh:2023:BID

[OY23] Seunghyun Oh and Sangwoon Yoo. Beyond imagined discontinuity: Review of the book series Science and Civilization in Korea.

Technology and Culture, 64(4):1274–1291, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/911006.

Page:2023:PWB

[Pag23] Arnaud Page. Pure white bread? Bleached flour, contestations, and regulation in Great Britain, 1900–50. Technology and Culture, 64(1):149–171, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/879585.

Parisi:2020:PBH

[Par20a] David Parisi. Power Button: A History of Pleasure, Panic, and the Politics of Pushing by Rachel Plotnick (review). Technology and Culture, 61(1):357–358, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/752953.

Park:2020:PBA

[Par20b] Sun-Young Park. The Printed and the Built: Architecture, Print Culture and Public Debate in the Nineteenth Century ed. by Mari Hvattum, Anne Hultzsch (review). Technology and Culture, 61 (1):349–350, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/752946.

Paterson:2023:HCM

[Pat23] Mark Paterson. Horn, or The Counterside of Media by Henning Schmidgen (review). Technology and Culture, 64(4):1343–1344, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
911034.

Patrick:2024:GBB

[Pat24] Annie Y. Patrick. Giving Bodies Back to Data: Image Makers, Bricolage, and Reinvention in Magnetic Resonance Technology by Silvia Casini (review). Technology and Culture, 65(2):705–707, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
926335.

Pearson:2022:TIW

[Pea22] Jessica Lynne Pearson. Technological Internationalism and World Order: Aviation, Atomic Energy, and the Search for International Peace, 1920–1950 by Waqar H. Zaidi (review). Technology and Culture, 63(3):891–892, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/859747.

Pelgrims:2024:VCH

[Pel24] Claire Pelgrims. Velocipedomania: A Cultural History of the Velocipede in France by Corry Cropper and Seth Whidden (review). Technology and Culture, 65(1):360–361, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/920530.

Peng:2022:CFC

[Pen22] Xin Peng. Chinatown Film Culture: The Appearance of Cinema in San Francisco’s Chinese Neighborhood by Kim K. Fahlstedt (review). Technology and Culture, 63(4):1221–1223, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/868073.

Pereira:2021:RIR

[Per21a] Hugo Silveira Pereira. Railway imperialism revisited: The failed line from Macao to Guangzhou. Technology and Culture, 62(1): 82–104, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/783906.

Perlman:2021:FCH

[Per21b] Allison Perlman. The Future of Change: How Technology Shapes Social Revolutions by Ray Brescia (review). Technology and Culture, 62(2):650–651, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/794385.

Pereira:2023:TRC

[Per23a] Hugo Silveira Pereira. Transnational Railway Cultures: Trains in Music, Literature, Film, and Visual Art ed. by Benjamin Fraser and Steven Spalding (review). Technology and Culture, 64(1): 271–272, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/879604.

Pereira:2023:CRF

[Per23b] Hugo Silveira Pereira. The camera and the railway: Framing the Portuguese Empire and technological landscapes in Angola and Mozambique, 1880s–1910s. Technology and Culture, 64(3):737– 759, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
903971.

Peters:2024:PIL

[Pet24a] Mario Peters. Pensar las infraestructuras en Latinoamérica [Thinking about infrastructures in Latin America] ed. by Dhan Zunino Singh, Valeria Gruschetsky, and Melina Piglia (review). Technology and Culture, 65(2):693–695, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/926330.

Petricek:2024:LRA

[Pet24b] Tomas Petricek. Language and the Rise of the Algorithm by Jeffrey M. Binder (review). Technology and Culture, 65(1):427–429, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920566.

Pieters:2022:HDP

[Pie22] Toine Pieters. That High Design of Purest Gold: A Critical History of the Pharmaceutical Industry, 1880–2020 by Graham Dutfield (review). Technology and Culture, 63(2):593–594, April 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/854044.

Pike:2021:CRM

[Pik21] David Pike. Cities, Railways, Modernities: London, Paris, and the Nineteenth Century by Carlos López Galviz (review). Technology and Culture, 62(2):626–627, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/794397.

Pineda:2021:TLA

[Pin21] Yovanna Pineda. Technology in Latin American History ed. by David Pretel and Helge Wendt (review). Technology and Culture, 62(3):902–903, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/803144.

Plotnick:2020:ATI

[Plo20] Rachel Plotnick. Archaeologies of Touch: Interfacing with Haptics from Electricity to Computers by David Parisi (review). Technology and Culture, 61(4):1229–1231, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/777634.

Plouviez:2023:TVB

[Plo23] David Plouviez. Turret versus Broadside: An Anatomy of British Naval Prestige, Revolution and Disaster, 1860–1870 by Howard J. Fuller (review). Technology and Culture, 64(3):963–965, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/903997.

Prieto-Nanez:2022:THI

[PÑ22] Fabian Prieto-Ñañez. A Transnational History of the Internet in Central America, 1985–2000: Networks, Integration, and Development by Ignacio Siles (review). Technology and Culture, 63 (2):582–583, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/854038.

Polonia:2021:MSB

[Pol21] Amélia Polónia. The Merchant Ship in the British Atlantic, 1600– 1800: Continuity and Innovation in a Key Technology by Phillip Reid (review). Technology and Culture, 62(3):933–935, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/803122.

Pollock:2024:NRA

[Pol24] Ethan Pollock. Nuclear Russia: The Atom in Russian Politics and Culture by Paul R. Josephson (review). Technology and Culture, 65(1):388–390, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/920545.

Pomeroy:2020:BBD

[Pom20] Steven A. Pomeroy. Battleship Bismarck: A Design and Operational History by William H. Garzke Jr. et al. (review). Technology and Culture, 61(3):971–973, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/763605.

Pooley:2023:NBN

[Poo23] Colin Pooley. No Bicycle, No Bus, No Job: The Making of Workers’ Mobility in the Netherlands, 1920–1990 by Patrick Bek (review). Technology and Culture, 64(3):958–960, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/903994.

Pope:2020:RKS

[Pop20] Norris Pope. The reception of kodachrome sheet film in American commercial photography. Technology and Culture, 61(1):1–41, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
752959.

Popplow:2024:TMG

[Pop24] Marcus Popplow. Technology in Modern German History: 1800 to the Present by Karsten Uhl (review). Technology and Culture, 65(2):744–745, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/926356.

Post:2023:GAT

[Pos23] Bob Post. The Great American Transit Disaster: a Century of Austerity, Auto-Centric Planning, and White Flight by Nicholas Dagen Bloom (review). Technology and Culture, 64(4):1335–1336, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
911030.

Pouillard:2023:LHE

[Pou23a] Violette Pouillard. L’animal et l’homme: De l’exploitation à la sauvegarde ed. by Anne-Marie Flambard–Héricher and François Blary (review). Technology and Culture, 64(2):582–584, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893051.

Poulsen:2023:HCW

[Pou23b] René Taudal Poulsen. A History of Cold War Industrialisation: Finnish Shipbuilding between East and West by Saara Matala (review). Technology and Culture, 64(1):280–281, January 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/879619.

Pouillard:2024:DDS

[Pou24] Véronique Pouillard. Les débuts du système suisse des brevets d’invention (1873–1914) by Nicolas Chachereau (review). Technology and Culture, 65(1):401–403, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/920552.

Prasad:2020:TIR

[Pra20] Amit Prasad. The Technological Indian by Ross Bassett (review). Technology and Culture, 61(1):354–357, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/752938.

Prasad:2023:LCB

[Pra23] Amit Prasad. At the Limits of Cure by Bharat Jayram Venkat (review). Technology and Culture, 64(4):1301–1302, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/911012.

PresasIPuig:2021:TTN

[Pre21] Albert Presas I Puig. Taking on Technocracy: Nuclear Power in Germany, 1945 to the Present by Dolores L. Augustine (review). Technology and Culture, 62(1):273–274, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/783924.

PresasIPuig:2023:LMV

[Pre23a] Albert Presas I Puig. Une longue marche vers l’indépendance et la transparence: L’histoire de l’Autorité de sûreté nucléaire française by Philippe Saint Raymond, and: Confessions of a Rogue Nuclear Regulator by Gregory B. Jaczko (review). Technology and Culture, 64(2):622–625, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893073.

Pretel:2023:HCG

[Pre23b] David Pretel. Hidden connections: The global history of jungle commodities. Technology and Culture, 64(1):202–219, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
879587.

Prytherch:2023:JIR

[Pry23] David Prytherch. Justice and the Interstates: The Racist Truth about Urban Highways ed. by Ryan Reft, Amanda Phillips de Lucas, and Rebecca Retzlaff (review). Technology and Culture, 64 (4):1327–1329, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/911026.

Puig:2022:AEM

[Pui22] Nuria Puig. Angels of Efficiency: A Media History of Consulting by Florian Hoof (review). Technology and Culture, 63(1):264–266, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
844187.

Prescod-Weinstein:2024:BSW

[PW24] Chanda Prescod-Weinstein. Brown Skins, White Coats: Race Science in India, 1920–66 by Projit Bihari Mukharji, and: Medicine, Science, and Making Race in Civil War America by Leslie A. Schwalm, and: Masters of Health: Racial Science and Slavery in U.S. Medical Schools by Christopher D. E. Willoughby (review). Technology and Culture, 65(2):701–705, April 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/926334.

Perez-Zapico:2023:FBL

[PZ23] Daniel Pérez-Zapico. Faith in a Beam of Light: Magic Lantern and Belief in Western Europe, 1860–1940 ed. by Sabine Lenk and Natalija Majsova (review). Technology and Culture, 64(3):992–994, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
904013.

Rader:2023:BMR

[Rad23] Karen A. Rader. Brain Magnet: Research Triangle Park and the Idea of the Idea Economy by Alex Sayf Cummings (review). Technology and Culture, 64(2):638–639, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/893081.

Ragas:2024:HDC

[Rag24] José Ragas. How to design a (cybernetic) revolution: Chileans remember Salvador Allende’s Techno-Utopian Project. Technology and Culture, 65(2):667–674, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/926320.

Rai:2022:RHP

[Rai22] Saurav Kumar Rai. Reaching out: How the pandemic impacted knowledge production and dissemination in India. Technology and Culture, 63(2):471–476, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/854008.

Rai:2023:HFS

[Rai23] Saurav Kumar Rai. The History of Forensic Science in India by Saumitra Basu (review). Technology and Culture, 64(2):584–585, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893052.

Ramesh:2021:WEH

[Ram21a] Aditya Ramesh. Water and the Environmental History of Modern India by Velayutham Saravanan (review). Technology and Culture, 62(4):1235–1236, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/819808.

Ramnath:2021:TMN

[Ram21b] Aparajith Ramnath. Temples of Modernity: Nationalism, Hinduism, and Transhumanism in South Indian Science by Robert Geraci (review). Technology and Culture, 62(1):311–312, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
783946.

Raman:2022:WHT

[Ram22a] Alka Raman. Weaving Histories: The Transformation of the Handloom Industry in South India, 1800–1960 by Karuna Dietrich Wielenga (review). Technology and Culture, 63(1):261–262, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
844185.

Ramesh:2022:TGC

[Ram22b] Sangaralingam Ramesh. Tata: The Global Corporation That Built Indian Capitalism by Mircea Raianu (review). Technology and Culture, 63(3):920–921, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/859762.

Ramirez:2022:PEA

[Ram22c] Susan Elizabeth Ramírez. Perspectives on Early Andean Civilization in Peru: Interaction, Authority, and Socioeconomic Organization during the First and Second Millennia B.C. ed. by Richard L. Burger et al. (review). Technology and Culture, 63(1):236–237, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
844171.

Raman:2023:HMH

[Ram23] Alka Raman. From hand to machine: How Indian cloth quality shaped British cotton spinning technology. Technology and Culture, 64(3):707–736, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/903970.

Rashidbeigi:2022:IMM

[Ras22] Samin Rashidbeigi. Iran in Motion: Mobility, Space, and the Trans-Iranian Railway by Mikiya Koyagi (review). Technology and Culture, 63(1):280–282, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/844195.

Ravanis:2020:EWM

[Rav20] Julia Ravanis. The Eye of War: Military Perception from the Telescope to the Drone by Antoine J. Bousquet (review). Technology and Culture, 61(3):973–974, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/763599.

Reynolds-Cuellar:2024:SDK

[RCCAM24] Pedro Reynolds-Cuéllar, Diego Cerna-Aragon, and Eden Medina. Seeds, dams, and khipus: Latin America’s eclectic recent history of technology. Technology and Culture, 65(2):447–472, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/926311.

Rees:2022:RTF

[Ree22] Jonathan Rees. Risk on the Table: Food Production, Health, and the Environment ed. by Angela N. H. Creager and Jean-Paul Gaudillière (review). Technology and Culture, 63(3):899–900, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
859751.

Reid:2021:SUB

[Rei21a] Phillip Reid. Science, Utility and British Naval Technology, 1793– 1815: Samuel Bentham and the Royal Dockyards by Roger Morriss (review). Technology and Culture, 62(3):937–939, July 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/803124.

Reinhardt:2021:PNC

[Rei21b] Anne Reinhardt. The Profits of Nature: Colonial Development and the Quest for Resources in Nineteenth-Century China by Peter B. Lavelle (review). Technology and Culture, 62(1):276–278, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
783926.

Reith:2021:MKK

[Rei21c] Reinhold Reith. Materielle Kultur und Konsum in der Frühen Neuzeit [Material culture and consumption in the Early Modern Era] ed. by Julia A. Schmidt-Funke (review). Technology and Culture, 62(2):595–597, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/794412.

Rentetzi:2022:RET

[Ren22] Maria Rentetzi. Radiation Evangelists: Technology, Therapy and Uncertainty at the Turn of the Century by Jeffrey Womack (review). Technology and Culture, 63(4):1230–1232, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/868078.

Reyes:2022:THC

[Rey22] Marc Reyes. A Technological History of Cold-War India, 1947– 1969: Autarky and Foreign Aid by William A. T. Logan (review). Technology and Culture, 63(4):1244–1246, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/868087.

Ribeiro:2023:NSC

[Rib23] Alyssa Ribeiro. Nuclear Suburbs: Cold War Technoscience and the Pittsburgh Renaissance by Patrick Vitale (review). Technology and Culture, 64(1):292–293, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/879630.

Riedl:2022:TSF

[Rie22] Matthias Riedl. Technik und Science-Fiction in der Vormoderne [Technology and science fiction in the premodern age] ed. by Brigitte Burrichter and Dorothea Klein (review). Technology and Culture, 63(3):916–917, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/859760.

Rikitianskaia:2021:RSC

[Rik21] Maria Rikitianskaia. Radio’s Second Century: Past, Present, and Future Perspectives ed. by John Allen Hendricks (review). Technology and Culture, 62(3):949–950, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/803130.

Rindzeviciute:2020:CT

[Rin20] Eglė Rindzevičiūtė. Chernobyl as technoscience. Technology and Culture, 61(4):1178–1187, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/777612.

Ringer:2022:GIV

[Rin22] Monica M. Ringer. Global Islam: A Very Short Introduction by Nile Green (review). Technology and Culture, 63(1):299–300, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
844205.

Rafiq:2023:STC

[RO23] Arif Ainur Rafiq and Hefri Asra Omika. Science, Technology, and the Cultural Cold War in Asia: From Atoms for Peace to Space Flight by Yuka Moriguchi Tsuchiya (review). Technology and Culture, 64(3):978–979, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/904005.

Robinson:2020:MTT

[Rob20] Francis Robinson. Modern Things on Trial: Islam’s Global and Material Reformation in the Age of Rida, 1865–1935 by Leor Halevi (review). Technology and Culture, 61(3):952–953, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
763606.

Robertson:2022:CSA

[Rob22a] Beth A. Robertson. Cesarean Section: An American History of Risk, Technology, and Consequence by Jacqueline H. Wolf (review). Technology and Culture, 63(2):552–553, April 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/854020.

Robertson:2022:GEV

[Rob22b] Jennifer Robertson. Glamorized exploitation: Visual images of Meiji-period “Factory Girls” (jokō). Technology and Culture, 63 (2):450–457, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/854005.

Robitaille:2022:MBA

[Rob22c] Anne-Julie Robitaille. Metalworking in Bronze Age China: The Lost-Wax Process by Peng Peng (review). Technology and Culture, 63(3):918–919, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/859761.

Rofheart:2020:FTC

[Rof20] Mahriana Rofheart. Fictional technologies of collaboration. Technology and Culture, 61(2S):S158–S180, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/761590.

Rogers:2021:SEE

[Rog21] Ariel Rogers. Seeing by Electricity: The Emergence of Television, 1878–1939 by Doron Galili (review). Technology and Culture, 62(3):951–952, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/803131.

Rogerson:2023:WNG

[Rog23] Ken Rogerson. When the Nerds Go Marching In: How Digital Technology Moved from the Margins to the Mainstream of Political Campaigns by Rachel K. Gibson (review). Technology and Culture, 64(2):602–604, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893062.

Romero:2022:HCA

[Rom22] Adam M. Romero. Hazardous Chemicals: Agents of Risk and Change, 1800–2000 ed. by Ernst Homburg and Elisabeth Vaupel (review). Technology and Culture, 63(1):263–264, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/844186.

Rosenthal:2020:MAW

[Ros20] Caitlin Rosenthal. Manufacturing Advantage: War, the State, and the Origins of American Industry, 1776–1848 by Lindsay Schakenbach Regele (review). Technology and Culture, 61(2):691–693, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
761585.

Rosson:2022:TAE

[Ros22a] Lois Rosson. Through Astronaut Eyes: Photographing Early Human Spaceflight by Jennifer K. Levasseur (review). Technology and Culture, 63(1):289–291, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/844200.

Rostvik:2022:TTB

[Røs22b] Camilla Mørk Røstvik. Tampon technology in Britain: Unilever’s Project Hyacinth and the “7-Day War” Campaign, 1968–1980. Technology and Culture, 63(1):61–86, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/844164.

Roy:2021:URT

[Roy21] Tirthankar Roy. Useful & reliable: Technological transformation in colonial India. Technology and Culture, 62(2):494–520, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
794374.

Roy:2024:FHW

[Roy24] Kaushik Roy. Firepower: How Weapons Shaped Warfare by Paul Lockhart (review). Technology and Culture, 65(1):431–432, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920568.

Roca-Rosell:2021:AMA

[RR21a] Antoni Roca-Rosell. Art and matter: artisans, architects, and technology (1770–1830) by Valérie Nègre (review). Technology and Culture, 62(1):292–293, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/783935.

Roca-Rosell:2021:JEG

[RR21b] Antoni Roca-Rosell. Jeux d’eaux à La Granja. Les fontaniers français en Espagne au XVIII^{e} siècle [Water games at La Granja: French plumbers in Spain in the eighteenth century] by Sophie Omère (review). Technology and Culture, 62(4):1239–1240, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
819810.

Ruamcharoen:2024:TPR

[Rua24] Boyd Ruamcharoen. Tropicalizing the portable radio: Electronics and the U.S. military’s battle against fungi in the Pacific war. Technology and Culture, 65(2):497–529, April 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/926313.

Rupcic:2024:NLH

[Rup24] Tijana Rupčić. In a New Light: Histories of Women and Energy ed. by Abigail Harrison Moore and R. W. Sandwell (review). Technology and Culture, 65(2):725–727, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/926346.

Russell:2020:TTR

[Rus20] Edmund Russell. The Train and the Telegraph: A Revisionist History by Benjamin Sidney Michael Schwantes (review). Technology and Culture, 61(3):992–993, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/763616.

Russell:2022:CMH

[Rus22] Edmund Russell. Capitalism matters: How financial and technological innovations shaped U.S. telegraphs, 1845–60. Technology and Culture, 63(1):31–60, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/844163.

Russell:2024:LEH

[Rus24a] Andrew L. Russell. Labor’s End: How the Promise of Automation Degraded Work by Jason Resnikoff (review). Technology and Culture, 65(2):707–709, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/926336.

Russell:2024:ARC

[Rus24b] Ben Russell. The American Robot: a Cultural History by Dustin A. Abnet (review). Technology and Culture, 65(2):712–714, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
926339.

Russell:2023:AHD

[RW23] Edmund Russell and Lauren Winkler. Animating history: Digitally mapping the United States telegraph system. Technology and Culture, 64(4):1211–1225, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/911001.

Ryden:2020:TCE

[Ryd20] Göran Rydén. Transnational Cultures of Expertise: Circulating State-Related Knowledge in the 18th and 19th Centuries ed. by Lothar Schilling and Jacob Vogel (review). Technology and Culture, 61(4):1242–1243, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/777641.

Sada:2024:MMA

[Sad24] Eva Rivas Sada. Making Machines of Animals: The International Livestock Exposition by Neal A. Knapp (review). Technology and Culture, 65(1):372–374, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/920537.

Santos:2020:BST

[San20] Gonçalo Santos. Birthing stories and techno-moral change across generations: Coping with hospital births and high-tech medicalization in rural South China, 1960s–2010s. Technology and Culture, 61(2):581–616, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/761577.

Sanvitale:2021:CDE

[San21] Ginevra Sanvitale. Il computer è donna. Eroine geniali e visionarie che hanno fatto la storia del’informatica [The computer is female: Brilliant and visionary heroines who made computer science history] by Carla Petrocelli (review). Technology and Culture, 62(3):925–926, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/803117.

Sarkar:2022:CLW

[Sar22] Smritikumar Sarkar. Classes of Labour: Work and Life in a Central Indian Steel Town by Jonathan Parry (review). Technology and Culture, 63(4):1201–1202, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/868061.

Shu:2020:TMC

[SC20a] Chang-Xue Shu and Thomas Coomans. Towards modern ceramics in China: Engineering sources and the manufacture céramique de Shanghai. Technology and Culture, 61(2):437–479, April 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/761572.

Slayton:2020:TIE

[SC20b] Rebecca Slayton and Brian Clarke. Trusting infrastructure: The emergence of computer security incident response, 1989–2005. Technology and Culture, 61(1):173–206, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/752964.

Scales:2020:BHA

[Sca20] Rebecca Scales. Barriers Down: How American Power and Free-Flow Policies Shaped Global Media by Diana Lemberg (review). Technology and Culture, 61(3):957–958, July 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/763608.

Scarpellini:2021:CSC

[Sca21] Emanuela Scarpellini. La cucina: storia culturale di un luogo domestico [The kitchen: The cultural history of a domestic space] by Imma Forino (review). Technology and Culture, 62(1):320–321, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
783951.

Scarpellini:2023:LIT

[Sca23] Emanuela Scarpellini. Leonardo to the Internet: Technology and Culture from the Renaissance to the Present by Thomas J. Misa (review). Technology and Culture, 64(1):220–221, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/879620.

Schatzberg:2020:WDM

[Sch20a] Eric Schatzberg. Where do models of innovation come from? Benoit Godin, Models of Innovation. Technology and Culture, 61(1):337–340, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/752935.

Schlombs:2020:WDG

[Sch20b] Corinna Schlombs. Wege in die Digitale Gesellschaft: Computernützung in der Bundesrepublik 1955–1990 [Towards a digital society: Computer use in West Germany 1955–1990] ed. by Frank Bösch (review). Technology and Culture, 61(4):1248–1250, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
777645.

Schmid:2020:CTS

[Sch20c] Sonja D. Schmid. Chernobyl the TV series: On suspending the truth or what’s the benefit of lies? Technology and Culture, 61 (4):1154–1161, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/777610.

Schultz:2020:HAA

[Sch20d] Tim Schultz. Harnessing the Airplane: American and British Cavalry Responses to a New Technology, 1903–1939 by Lori A. Henning (review). Technology and Culture, 61(1):364–365, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
752952.

Schafer:2021:UWS

[Sch21a] Dagmar Schäfer. Useful work: State demands and craftsmen’s social mobility in Fifteenth-Century China. Technology and Culture, 62(2):373–400, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/794369.

Schmitt:2021:WWC

[Sch21b] Martin Schmitt. Wie die Welt in den Computer kam: Zur Entstehung digitaler Wirklichkeit [How the world came to be in the computer: On the creation of digital reality] by David Gugerli (review). Technology and Culture, 62(2):657–658, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794399.

Schultz:2021:GDH

[Sch21c] Tim Schultz. The Good Drone: How Social Movements Democratize Surveillance by Austin Choi-Fitzpatrick, and: Unmanning: How Humans, Machines and Media Perform Drone Warfare by Katherine Chandler (review). Technology and Culture, 62(3):927– 929, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
803118.

Schwoch:2022:CET

[Sch22] James Schwoch. Communicating Europe: Technologies, Information, Events by Andreas Fickers and Pascal Griset (review). Technology and Culture, 63(1):291–293, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/844201.

Schafer:2023:AEN

[Sch23a] Valérie Schafer. Abstractions and Embodiments: New Histories of Computing and Society ed. by Janet Abbate and Stephanie Dick (review). Technology and Culture, 64(2):611–613, April 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/893067.

Schlombs:2023:BHW

[Sch23b] Corinna Schlombs. Built on the hands of women: Data, automation, and gender in West Germany’s financial industry. Technology and Culture, 64(1):63–89, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/879582.

Schuld:2023:NSI

[Sch23c] Dawna Schuld. Northern Sparks: Innovation, Technology Policy, and the Arts in Canada from Expo 67 to the Internet Age by Michael Century (review). Technology and Culture, 64(4): 1318–1319, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/911021.

Scotti:2023:SWM

[Sco23] Alina Scotti. The Stone and the Wireless: Mediating China, 1861– 1906 by Shaoling Ma (review). Technology and Culture, 64(1): 238–239, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/879617.

Suarez-Diaz:2023:IES

[SD23] Edna Suárez-Díaz. Itineraries of Expertise: Science, Technology and the Environment in Latin America’s Long Cold War ed. by Andra B. Chastain and Timothy W. Lorek (review). Technology and Culture, 64(3):974–976, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/904003.

Soca:2024:FFT

[SD24] Fernanda A. Soca and Mariana E. Di Bello. Fragile foundations: Tracing Argentina’s semiconductor saga amid institutional turmoil. Technology and Culture, 65(2):555–570, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/926315.

Seely:2020:PMT

[See20a] Bruce Seely. Power Moves: Transportation, Politics, and Development in Houston by Kyle Shelton (review). Technology and Culture, 61(1):369–371, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/752955.

Seely:2020:TNT

[See20b] Bruce Seely. Taking Nazi Technology: Allied Exploitation of German Science after the Second World War by Douglas O’Reagan (review). Technology and Culture, 61(4):1236–1238, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/777638.

Seefried:2022:PFZ

[See22] Elke Seefried. Das Projekt Futurologie: Über Zukunft und Fortschritt in der Bundesrepublik 1952–1982 [The futurology project: About future and progress in West Germany, 1952–1982] by Achim Eberspächer (review). Technology and Culture, 63 (3):914–915, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/859759.

Seitz:2024:SSC

[Sei24] John B. Seitz. Seeds of the settler colony: How peasant and Kazakh knowledge, environment, and bureaucracy shaped steppe agronomy in the Late Russian Empire. Technology and Culture, 65(2):531–554, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/926314.

Schipper:2020:SUM

[SEO20] Frank Schipper, Martin Emanuel, and Ruth Oldenziel. Sustainable urban mobility in the present, past, and future. Technology and Culture, 61(1):307–317, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/752932.

Serban:2024:ELD

[Şer24] Stelu Şerban. Engineering the Lower Danube: Technology and International Cooperation in an Imperial Borderland by Luminita Gatejel (review). Technology and Culture, 65(1):394–395, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920548.

Santos:2020:RRT

[SG20] Gonçalo Santos and Suzanne Z. Gottschang. Rethinking reproductive technologies and modernities in time and space. Technology and Culture, 61(2):549–558, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/761575.

Szczepaniak-Gillece:2021:SDM

[SG21] Jocelyn Szczepaniak-Gillece. On the Screen: Displaying the Moving Image, 1926–1942 by Ariel Rogers (review). Technology and Culture, 62(2):646–647, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/794410.

Shabtay:2022:EHT

[Sha22a] Talia Shabtay. Engineering Hollywood: Technology, Technicians, and the Science of Building the Studio System by Luci Marzola (review). Technology and Culture, 63(2):574–575, April 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/854033.

Shapiro:2022:ETI

[Sha22b] Justin Shapiro. Empire’s Tracks: Indigenous Nations, Chinese Workers, and the Transcontinental Railroad by Manu Karuka (review). Technology and Culture, 63(2):585–587, April 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/854040.

Shell:2020:HAO

[She20] Hanna Rose Shell. High altitude observatory: Conversations on the plurality of worlds. Technology and Culture, 61(1):239–259, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
752966.

Shirai:2020:HDC

[Shi20] Chiaki Shirai. Historical dynamism of childbirth in Japan: Medicalization and its normative politics, 1868–2017. Technology and Culture, 61(2):559–580, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/761576.

Siegelbaum:2022:GAE

[Sie22] Lewis H. Siegelbaum. Globalizing Automobilism: Exuberance and the Emergence of Layered Mobility, 1900–1980 by Gijs Mom (review). Technology and Culture, 63(1):278–280, January 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/844194.

Sims:2020:ALI

[Sim20] Benjamin Sims. The American Lab: An Insider’s History of the Lawrence Livermore National Laboratory by C. Bruce Tarter (review). Technology and Culture, 61(4):1238–1240, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/777639.

Simon:2024:IWM

[Sim24] Sara M. B. Simon. Immeasurable Weather: Meteorological Data and Settler Colonialism from 1820 to Hurricane Sandy by Sara J. Grossman (review). Technology and Culture, 65(2):720–722, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
926343.

Singh:2023:UIH

[Sin23] Dhan Zunino Singh. Urban Infrastructure: Historical and Social Dimensions of an Interconnected World ed. by Joseph Heathcott, Jonathan Soffer, and Rae Zimmerman (review). Technology and Culture, 64(3):946–948, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/903988.

Slavishak:2022:CCR

[Sla22] Edward Slavishak. Collision course: Rural track crossing habits and the railroad in the United States, 1915–32. Technology and Culture, 63(1):209–233, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/844169.

Slayton:2023:VSH

[Sla23] Rebecca Slayton. A Vulnerable System: The History of Information Security in the Computer Age by Andrew Stewart (review). Technology and Culture, 64(1):232–234, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/879628.

Slaughter:2024:NIP

[Sla24] Aimee Slaughter. Nuclear imaginaries and power in Oppenheimer. Technology and Culture, 65(1):319–332, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/920526.

Suresh:2021:BSI

[SM21] Pavithra Suresh and Alexander Monea. Black Software: The Internet and Racial Justice, from the AfroNet to Black Lives Matter by Charlton D. McIlwain (review). Technology and Culture, 62(3):896–897, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/803121.

Smethurst:2020:IBP

[Sme20] Paul Smethurst. Invisible Bicycle: Parallel Histories and Different Timelines ed. by Tiina Männistö-Funk, Timo Myllyntaus (review). Technology and Culture, 61(1):360–362, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/752950.

Smith:2020:EEN

[Smi20] Jenny Leigh Smith. Eurasian Environments: Nature and Ecology in Imperial Russian and Soviet History ed. by Nicholas Breyfogle (review). Technology and Culture, 61(1):352–354, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/752940.

Smithstead:2023:MWW

[Smi23] Lisa Smithstead. Movie Workers: The Women Who Made British Cinema by Melanie Bell (review). Technology and Culture, 64 (1):242–243, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/879590.

Soentgen:2021:DOS

[Soe21] Jens Soentgen. Deutschlands Ölfelder. Eine Stoffgeschichte der Kulturpflanze Raps (1897–2017) [Germany’s oil fields: A material history of the crop plant rapeseed (1897–2017)] by Sarah Waltenberger (review). Technology and Culture, 62(1):275–276, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
783925.

Soland:2024:CEG

[Sol24a] Peter Soland. Cover essay: La gente, controllers of the universe. Technology and Culture, 65(2):437–445, April 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/926310.

Solares:2024:PTM

[Sol24b] Israel G. Solares. El pensamiento sobre la técnica en México [Thinking about technology in Mexico] ed. by Irving Samadhi Aguilar Rocha and José Francisco Barrón Tovar (review). Technology and Culture, 65(2):690–691, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/926328.

Sorlin:2023:ETP

[Sör23] Sverker Sörlin. The Environmental Turn in Postwar Sweden: a New History of Knowledge by David Larsson Heidenblad (review). Technology and Culture, 64(4):1320–1321, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/911022.

Soulier:2024:LTH

[Sou24a] Philippe Soulier. L’invention de la technologie: Une histoire intellectuelle avec André Leroi-Gourhan [The invention of technology: an intellectual history with André Leroi-Gourhan] by Nathan Schlanger (review). Technology and Culture, 65(2):748– 749, April 2024. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/926358.

Sousa:2024:CCP

[Sou24b] M. Luísa Sousa. Capital Científica: Práticas da Ciência em Lisboa e a História Contemporânea de Portugal [Scientific capital:

Science practices in Lisbon and the contemporary history of Portugal] ed. by Tiago Saraiva and Marta Macedo (review). Technology and Culture, 65(2):750–751, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/926359.

Spinak:2020:QFA

[Spi20] Abby Spinak. “Not quite so freely as air”: Electrical statecraft in North America. Technology and Culture, 61(1):71–108, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
752961.

Spinak:2021:WAP

[Spi21] Abby Spinak. The World as an Architectural Project by Hashim Sarkis et al. (review). Technology and Culture, 62(4):1250–1252, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
819782.

Spinak:2024:PAF

[Spi24] Abby Spinak. Powering American Farms: The Overlooked Origins of Rural Electrification by Richard F. Hirsh (review). Technology and Culture, 65(1):378–380, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/920540.

Scott-Smith:2021:HKM

[SS21] Giles Scott-Smith. How Knowledge Moves: Writing the Transnational History of Science and Technology ed. by John Krige (review). Technology and Culture, 62(1):262–263, January 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/783918.

Stapleton:2020:ASM

[Sta20] Darwin Stapleton. Accounting for Slavery: Masters and Management by Caitlin Rosenthal (review). Technology and Culture, 61(1):346–347, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/752954.

Stadelmann:2022:WWW

[Sta22] Christian Stadelmann. Warmes Wasser — Weiße Ware: Energiewende im Badezimmer, 1880–1939 [Warm water – White goods: Energy transition in the bathroom, 1880–1939] by Nina Lorkowski (review). Technology and Culture, 63(4):1258–1259, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
868095.

Stanton:2023:DDM

[Sta23a] N. Bucky Stanton. Designs of Destruction: The Making of Monuments in the Twentieth Century by Lucia Allais (review). Technology and Culture, 64(1):294–295, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/879589.

Stark:2023:DWW

[Sta23b] James Stark. The Doctor Who Wasn’t There: Technology, History, and the Limits of Telehealth by Jeremy A. Greene (review). Technology and Culture, 64(4):1367–1369, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/911046.

Stein:2020:PSA

[Ste20] Blair Stein. Poisonous Skies: Acid Rain and the Globalization of Pollution by Rachel Emma Rothschild (review). Technology and Culture, 61(4):1263–1265, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/777653.

Stein:2022:SES

[Ste22a] Susanne Stein. The Shenzhen Experiment: The Story of China’s Instant City by Juan Du (review). Technology and Culture, 63 (2):587–589, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/854041.

Stemberk:2022:CDS

[Šte22b] Jan Štemberk. Cesty a dial’nice na Slovensku v medzivojnovom období [Roads and highways in Slovakia during the interwar period] by Michal Ďurčo (review). Technology and Culture, 63(4): 1194–1195, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/868097.

Stoneman:2022:WMW

[Sto22] Timothy H. B. Stoneman. When the Medium Was the Mission: The Atlantic Telegraph and the Religious Origins of Network Culture by Jenna Supp-Montgomerie (review). Technology and Culture, 63(1):300–302, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/844206.

Storni:2023:CEA

[Sto23] Marco Storni. Crafting Enlightenment: Artisanal Histories and Transnational Networks ed. by Lauren R. Cannady and Jennifer Ferng (review). Technology and Culture, 64(3):970–972, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/904001.

Stoneman:2024:MCT

[Sto24] Timothy H. B. Stoneman. The Marvelous Clouds: Toward a Philosophy of Elemental Media by John Durham Peters (review). Technology and Culture, 65(2):735–736, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/926351.

Stranges:2022:URE

[Str22a] Anthony N. Stranges. Unnatural Resources: Energy and Environmental Politics in Appalachia after the 1973 Oil Embargo by Michael Camp (review). Technology and Culture, 63(1):274–276, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
844192.

Streb:2022:IIP

[Str22b] Jochen Streb. Inventing Ideas: Patents, Prizes, and the Knowledge Economy by B. Zorina Khan (review). Technology and Culture, 63(1):252–253, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/844180.

Suarez:2021:RMH

[Suá21] Iria Suárez. Rise of the Modern Hospital: An Architectural History of Health and Healing, 1870–1940 by Jeanne Kisacky (review). Technology and Culture, 62(1):282–283, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/783929.

Subramanian:2022:SSO

[Sub22] Ramesh Subramanian. Shaping Science: Organizations, Decisions, and Culture on NASA’s Teams by Janet Vertesi (review). Technology and Culture, 63(1):304–306, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/844208.

Schafer:2021:TGU

[SV21] Dagmar Schäfer and Simona Valeriani. Technology is global: The useful & reliable knowledge debate. Technology and Culture, 62 (2):327–347, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/794367.

Svelch:2021:ADI

[Šve21] Jaroslav Švelch. Atari Design: Impressions on Coin-Operated Video Game Machines by Raiford Guins (review). Technology and Culture, 62(3):930–931, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/803119.

Swanson:2021:PNE

[Swa21] Drew Swanson. The Place with No Edge: An Intimate History of People, Technology, and the Mississippi River Delta by Adam Mandelman (review). Technology and Culture, 62(3):943–944, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
803127.

Sylvest:2022:DCE

[Syl22] Casper Sylvest. The Doomsday Clock at 75 ed. by Robert K. Elder and J. C. Gabel (review). Technology and Culture, 63(4): 1246–1248, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/868088.

Tan:2021:RTC

[Tan21a] Ying Jia Tan. Railroads and the Transformation of China by Elisabeth Köll (review). Technology and Culture, 62(3):962–963, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
803138.

Tanimoto:2021:IDU

[Tan21b] Masayuki Tanimoto. Introduction and diffusion: Useful and reliable knowledge in early modern industrial Japan. Technology and Culture, 62(2):423–441, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/794371.

Tapp:2023:TOW

[Tap23] Caroline Tapp. In Their Own Words: Forgotten Women Pilots of Early Aviation by Fred Erisman (review). Technology and Culture, 64(1):273–274, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/879603.

Tarr:2023:CAL

[Tar23] Joel A. Tarr. Conflicted American Landscapes by David E. Nye (review). Technology and Culture, 64(1):283–284, January 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/879622.

Tate:2021:RRP

[Tat21] Ryan Driskell Tate. Rural revolt: Power line protests and the alternative technology movement in the United States, 1970s. Technology and Culture, 62(1):1–26, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/783903.

Tatarchenko:2022:YCF

[Tat22] Ksenia Tatarchenko. Your Computer Is On Fire ed. by Thomas S. Mullaney et al. (review). Technology and Culture, 63(3):923–925, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
859764.

Taylor:2023:UCW

[Tay23] Rabun Taylor. Using and Conquering the Watery World in Greco– Roman Antiquity by Georgia L. Irby (review). Technology and Culture, 64(4):1294–1295, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/911008.

Telotte:2020:SSS

[Tel20] J. P. Telotte. Silent Serial Sensations: The Wharton Brothers and the Magic of Early Cinema by Barbara Tepa Lupack (review). Technology and Culture, 61(4):1231–1233, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/777635.

Tenner:2022:PTF

[Ten22] Edward Tenner. Prometheus Tamed: Fire, Security, and Modernities, 1400 to 1900 by Cornel Zwierlein (review). Technology and Culture, 63(3):925–926, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/859765.

Theodore:2022:CTR

[The22] David Theodore. Captivating Technology: Race, Carceral Technoscience, and Liberatory Imagination in Everyday Life ed. by Ruha Benjamin (review). Technology and Culture, 63(3):867–868, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
859734.

Thompson:2020:ANT

[Tho20] Drew Thompson. “Não há nada” (“There is nothing”): Absent headshots and identity documents in independent Mozambique. Technology and Culture, 61(2S):S104–S134, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/761588.

Thompson:2021:RPH

[Tho21] Peter Thompson. A Rainbow Palate: How Chemical Dyes Changed the West’s Relationship to Food by Carolyn Cobbold (review). Technology and Culture, 62(2):642–643, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/794387.

Thompson:2022:CAH

[Tho22] Peter Thompson. The Chemical Age: How Chemists Fought Famine and Disease, Killed Millions, and Changed Our Relationship with the Earth by Frank A. von Hippel (review). Technology and Culture, 63(3):902–904, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/859753.

Thompson:2023:AGM

[Tho23a] Peter Thompson. The Age of the Gas Mask: How British Civilians Faced the Terrors of Total War by Susan R. Grayzel (review). Technology and Culture, 64(3):932–933, July 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/903981.

Thomson:2023:VCA

[Tho23b] Jennifer Thomson. Valuing Clean Air: The EPA and the Economics of Environmental Protection by Charles Halvorson (review). Technology and Culture, 64(4):1322–1323, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/911023.

Thomas:2024:RMH

[Tho24] Pradip Ninan Thomas. Radio for the Millions: Hindi-Urdu Broadcasting Across Borders by Isabel Huacuja Alonso (review). Technology and Culture, 65(2):733–734, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/926350.

Tomory:2021:GSM

[Tom21] Leslie Tomory. Governing Systems: Modernity and the Making of Public Health in England, 1830–1910 by Tom Crook (review). Technology and Culture, 62(2):612–613, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794388.

Trischler:2020:WBS

[Tri20] Helmuth Trischler. Well-Being, Sustainability and Social Development: The Netherlands 1850–2050 by Harry Lintsen et al. (review). Technology and Culture, 61(3):963–965, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/763609.

Trudel:2024:HCF

[Tru24] Dominique Trudel. Une histoire de la cybernétique en France (1948–1975) by Ronan Le Roux (review). Technology and Culture, 65(1):426–427, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/920565.

Thebaud-Sorger:2020:CSM

[TS20] Marie Thébaud-Sorger. Changing scale to master nature: Promoting small-scale inventions in Eighteenth-Century France and Britain. Technology and Culture, 61(4):1076–1107, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/777603.

Tsukahara:2023:CMT

[Tsu23] Togo Tsukahara. Climatic Media: Transpacific Experiments in Atmospheric Control by Yuriko Furuhata (review). Technology and Culture, 64(1):285–286, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/879605.

Tunc:2022:LBE

[TT22] Tanfer Emin Tunc and Gokhan Tunc. “A light bulb in every house”: The Istanbul General Electric factory and American technology transfer to Turkey. Technology and Culture, 63(3):749–774, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
859723.

Tucker:2022:DVN

[Tuc22] Jennifer Tucker. Discourses of Vision in Nineteenth-Century Britain: Seeing, Thinking, Writing by Jonathan Potter (review). Technology and Culture, 63(2):535–536, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/854050.

Turpin:2020:WMF

[Tur20] Robert J. Turpin. Women on the Move: The Forgotten Era of Women’s Bicycle Racing by Roger Giles (review). Technology and Culture, 61(1):362–364, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/752942.

Twagira:2020:IAH

[Twa20a] Laura Ann Twagira. Introduction: Africanizing the history of technology. Technology and Culture, 61(2S):S1–S19, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/761591.

Twagira:2020:MCW

[Twa20b] Laura Ann Twagira. Machines that Cook or women who cook? Lessons from Mali on technology, labor, and women’s things. Technology and Culture, 61(2S):S77–S103, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/761587.

Tyson:2024:MAW

[Tys24] John A. Tyson. Making Art Work: How Cold War Engineers and Artists Forged a New Creative Culture by W. Patrick McCray (review). Technology and Culture, 65(2):714–716, April 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/926340.

Unger:2022:GRG

[Ung22] Corinna R. Unger. The Green Revolution in the Global South: Science, Politics, and Unintended Consequences by R. Douglas Hurt (review). Technology and Culture, 63(4):1237–1239, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
868083.

Utsumi:2021:PNP

[Uts21] Hirofumi Utsumi. Pathways into and Out of Nuclear Power in Western Europe: Austria, Denmark, Federal Republic of Germany, Italy, and Sweden ed. by Astrid Mignon Kirchhoff (review). Technology and Culture, 62(4):1283–1285, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/819800.

Valentines-Alvarez:2021:GFA

[VÁ21] Jaume Valentines-Álvarez. De la Guerra Fría al calentamiento global: Estados Unidos, España y el nuevo orden científico mundial [From Cold War to global warming: United States, Spain, and the new world scientific order] ed. by Lino Camprubí et al. (review). Technology and Culture, 62(3):904–906, July 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/803145.

Vail:2023:GSA

[Vai23] David D. Vail. Glyphosate and the Swirl: an Agroindustrial Chemical on the Move by Vincanne Adams (review). Technology and Culture, 64(4):1324–1325, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/911024.

Valeriani:2021:GBP

[Val21] Simona Valeriani. Grasping the body: Physicians, tailors, and holy people. Technology and Culture, 62(2):467–493, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794373.

VandenHeede:2020:RGH

[Van20a] Pieter Van den Heede. Respawn: Gamers, Hackers, and Technogenic Life by Colin Milburn (review). Technology and Culture, 61(4):1250–1252, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/777646.

VanDerVleuten:2020:HTA

[Van20b] Erik Van Der Vleuten. History and technology in an Age of “Grand Challenges”: Raising questions. Technology and Culture, 61(1):260–271, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/752928.

vanDijck:2021:SFS

[vD21] José van Dijck. Seeing into the Future: A Short History of Prediction by Martin van Creveld (review). Technology and Culture, 62(1):258–260, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/783916.

vandeLaar:2021:MIP

[vdL21] Paul van de Laar. Making Industrial Pittsburgh Modern: Environment, Landscape, Transportation, and Planning by Edward K. Muller and Joel A. Tarr (review). Technology and Culture, 62 (2):624–625, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/794408.

vanderLinden:2022:SSO

[vdL22] F. Robert van der Linden. Sovereign Skies: The Origins of American Civil Aviation Policy by Sean Seyer (review). Technology and Culture, 63(2):559–561, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/854025.

vanderStraeten:2022:LTL

[vdS22] Jonas van der Straeten. Let There Be Light: Engineering, Entrepreneurship and Electricity in Colonial Bengal, 1880–1945 by Suvobrata Sarkar (review). Technology and Culture, 63(2):550– 551, April 2022. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/854019.

vanEekelen:2022:HHE

[vE22] Bregje F. van Eekelen. Histories of Human Engineering: Tact and Technology by Maarten Derksen (review). Technology and Culture, 63(1):270–272, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/844190.

Veenis:2022:TLM

[Vee22] Milena Veenis. The Things of Life: Materiality in Late Soviet Russia by Alexey Golubev (review). Technology and Culture, 63 (1):306–308, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/844209.

Venkataraman:2024:WAB

[Ven24] Viswanathan Venkataraman. When the Air Became Important: A Social History of the New England and Lancashire Textile Industries by Janet Greenlees (review). Technology and Culture, 65(1):362–363, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/920531.

Vergara:2021:APM

[Ver21] Germán Vergara. Apostle of Progress: Modesto C. Rolland, Global Progressivism, and the Engineering of Revolutionary Mexico by Justin J. Castro (review). Technology and Culture, 62(3):900–902, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
803143.

Veraart:2022:KKR

[Ver22a] Frank Veraart. Kalkstein als ’Kritischer’ Rohstoff: Eine Stoffgeschichte der Industrialisierung, 1840–1930 [Limestone as a ’critical’ raw material: A material history of industrialization, 1840–1930] by Sebastian Haumann (review). Technology and Culture, 63(2):595–596, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/854045.

Verschueren:2022:IEX

[Ver22b] Pierre Verschueren. Ingénieurs et Entreprises, XIX^{e}–XXI^{e} siècle [Engineers and companies, 19th–21th centuries] ed. by Florent le Bot and Alain P. Michel (review). Technology and Culture, 63 (1):272–274, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/844191.

Vermij:2023:LPR

[Ver23] Rienk Vermij. Loath to Print: The Reluctant Scientific Author, 1500–1750 by Nicole Howard (review). Technology and Culture, 64(1):258–259, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/879612.

Vest:2020:PMP

[Ves20] J. Martin Vest. Popular Music in the Post Digital Age ed. by Ewa Mazierska, Les Gillon, Ton Rigg (review). Technology and Culture, 61(1):382–383, January 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/752951.

Vest:2023:MPY

[Ves23a] J. Martin Vest. “Mr. Phonograph, are you there?”: Mechanical failure and technological determinism, 1877–1900. Technology and Culture, 64(1):172–201, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/879586.

Vester:2023:ACM

[Ves23b] Katharina Vester. Artificial Color: Modern Food and Racial Fiction by Catherine Keyser (review). Technology and Culture, 64 (1):299–301, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/879614.

Vetter:2023:HFT

[Vet23] Florian Vetter. A History of Financial Technology and Regulation: From American Incorporation to Cryptocurrency and Crowdfunding by Seth C. Oranburg (review). Technology and Culture, 64 (2):615–616, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/893069.

Vincent:2024:HPB

[VF24] Jane Vincent and Leopoldina Fortunati. How prepaid billing in Italy helped shape the global diffusion of mobile phones. Technology and Culture, 65(1):293–314, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/920524.

Vikstrom:2020:PEL

[Vik20] Hanna Vikström. Producing electric light: How resource scarcity affected light bulbs, 1880–1914. Technology and Culture, 61 (3):901–922, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/763593.

Virdi:2022:UST

[Vir22] Jaipreet Virdi. The Untold Story of the Talking Book by Matthew Rubery (review). Technology and Culture, 63(1):243–244, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
844175.

vanMeer:2020:GIC

[vM20] Elisabeth van Meer. Gaming the Iron Curtain: How Teenagers and Amateurs in Communist Czechoslovakia Claimed the Medium of Computer Games by Jaroslav Švelch (review). Technology and Culture, 61(2):702–704, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/761566.

Volf:2022:CSS

[Vol22] Darina Volf. Collaboration in Space and the Search for Peace on Earth by Andrew L. Jenks (review). Technology and Culture, 63(3):885–886, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/859744.

Wilson:2023:LMC

[WAB^{+}23] Daniel C. S. Wilson, Mariona Coll Ardanuy, Kaspar Beelen, Barbara McGillivray, and Ruth Ahnert. The living machine: A computational approach to the Nineteenth-Century language of technology. Technology and Culture, 64(3):875–902, July 2023. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/903976.

Wachelder:2022:RFM

[Wac22] Joseph Wachelder. Renaissance Fun: The Machines behind the Scenes by Philip Steadman (review). Technology and Culture, 63(1):239–241, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/844173.

Wachelder:2023:MSS

[Wac23] Joseph Wachelder. The Moving Statues of Seventeenth-Century Amsterdam: Automata, Waxworks, Fountains, Labyrinths by Angela Vanhaelen (review). Technology and Culture, 64(4):1297– 1299, October 2023. CODEN TECUA3. ISSN 0040-165X (print),

1097-3729 (electronic). URL https://muse.jhu.edu/pub/1/
article/911010.

Wagner:2023:RKR

[Wag23] Hans-Ulrich Wagner. Radiobilder: Eine Kulturgeschichte des Radios in Österreich by Christine Ehardt (review). Technology and Culture, 64(3):996–998, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/904015.

Wallace:2021:ICL

[Wal21a] Harold D. Wallace, Jr. Inventing in a crisis: Lighting the United States after the 1973 Oil Embargo. Technology and Culture, 62 (4):1143–1171, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/819774.

Walton:2021:MME

[Wal21b] Steven A. Walton. The Medieval Military Engineer: From the Roman Empire to the Sixteenth Century by Peter Purton (review). Technology and Culture, 62(1):294–295, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/783936.

Walker:2022:RCP

[Wal22] Barbara Walker. From Russia with Code: Programming Migrations in Post-Soviet Times ed. by Mario Biagioli and Vincent Antonin Lépinay (review). Technology and Culture, 63(1):266–268, January 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
844188.

Wang:2021:CTM

[Wan21] Hsien-Chun Wang. Contested tracks to modernity: Negotiating narratives at Taiwan’s Railway Department Park. Technology and Culture, 62(2):573–583, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/794379.

Waanggren:2023:DMW

[Wån23] Lena Wånggren. Dissensuous Modernism: Women Writers, the Senses, and Technology by Allyson DeMaagd (review). Technology and Culture, 64(3):998–1000, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/904016.

Ward:2020:CMT

[War20] Jacob Ward. Computer models and thatcherist futures: From monopolies to markets in British telecommunications. Technology and Culture, 61(3):843–870, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/763585.

Ward:2023:CCS

[War23] Jacob Ward. Cyber and the City: Securing London’s Banks in the Computer Age by Ashley Sweetman (review). Technology and Culture, 64(2):613–615, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/893068.

Weiss:2020:MEV

[Wei20] Sean Weiss. Making engineering visible: Photography and the politics of drinking water in Modern Paris. Technology and Culture, 61(3):739–771, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/763579.

Wellum:2020:LPP

[Wel20] Caleb Wellum. Landscapes of Power: Politics of Energy in the Navajo Nation by Dana E. Powell (review). Technology and Culture, 61(4):1256–1257, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/777649.

Weldon:2021:ETA

[Wel21] Stephen P. Weldon. Existential Threats: American Apocalyptic Beliefs in the Technological Era by Lisa Vox (review). Technology and Culture, 62(1):313–314, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/783947.

Wendland:2020:UMS

[Wen20] Anna Veronika Wendland. Ukrainian memory spaces and nuclear technology: The musealization of Chornobyl’s disaster. Technology and Culture, 61(4):1162–1177, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/777611.

Wendt:2022:MLR

[Wen22] Helge Wendt. Mining Language: Racial Thinking, Indigenous Knowledge, and Colonial Metallurgy in the Early Modern Iberian World by Allison Margaret Bigelow (review). Technology and Culture, 63(2):533–534, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/854049.

Wendt:2024:FME

[Wen24] Helge Wendt. Fueling Mexico: Energy and Environment, 1850– 1950 by Germán Vergara, and: Electrifying Mexico: Technology and the Transformation of a Modern City by Diana J. Montaño (review). Technology and Culture, 65(2):685–687, April 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/926326.

Whitney:2023:TKE

[Whi23] Elspeth Whitney. Technical Knowledge in Europe, 1200–1500 AD ed. by Ricardo Córdoba de la Llave and Javier López Rider (review). Technology and Culture, 64(4):1295–1297, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/911009.

Wieser:2021:BHB

[Wie21] Bernhard Wieser. A Business History of the Bicycle Industry: Shaping Marketing Practices by Carlo Mari (review). Technology and Culture, 62(3):969–971, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/803142.

Wijfjes:2022:RPN

[Wij22] Huub Wijfjes. Reluctant Power: Networks, Corporations, and the Struggle for Global Governance in the Early 20th Century by Rita Zajácz (review). Technology and Culture, 63(2):561–563, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
854026.

Williams:2020:UGS

[Wil20] Rosalind Williams. An Underground Guide to Sewers, or: Down, Through & Out in Paris, London, New York &c by Stephen Halliday (review). Technology and Culture, 61(3):965–966, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/763607.

Williams:2021:CEA

[Wil21a] Rosalind Williams. Crisis: The emergence of another hazardous concept. Technology and Culture, 62(2):521–546, April 2021. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/794375.

Williot:2021:MAE

[Wil21b] Jean-Pierre Williot. Menaces sur l’alimentation. Emballages, colorants et autres contaminants alimentaires XIX^{e}–XXI^{e} siècles [Threats to food: Packaging, colorants and other food contaminants, 19th–21st centuries] by Florence Hachez-Leroy (review). Technology and Culture, 62(3):945–946, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/803128.

Williams:2022:SUF

[Wil22a] Amrys O. Williams. Supermarket USA: Food and Power in the Cold War Farms Race by Shane Hamilton (review). Technology and Culture, 63(3):895–896, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/859749.

Williamson:2022:MDN

[Wil22b] Bess Williamson. Measuring Difference, Numbering Normal: Setting the Standards for Disability in the Interwar Period by Coreen McGuire (review). Technology and Culture, 63(2):527–528, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
854035.

Williams:2023:DMP

[Wil23] Nicola Williams. Do microscopes have politics? Gendering the electron microscope in laboratory biological research. Technology and Culture, 64(4):1159–1183, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/910999.

Wirten:2023:FCV

[Wir23] Eva Hemmungs Wirtén. The Filing Cabinet: A Vertical History of Information by Craig Robertson (review). Technology and Culture, 64(1):222–223, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/879625.

Wisnioski:2022:EEI

[Wis22] Matthew Wisnioski. Economics and Engineering: Institutions, Practices, and Cultures ed. by Pedro Garcia Duarte and Yann Giraud (review). Technology and Culture, 63(2):523–525, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/854013.

Wittner:2021:MTA

[Wit21] David Wittner. Making Time: Astronomical Time Measurement in Tokugawa Japan by Yulia Frumer (review). Technology and Culture, 62(2):600–602, April 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/794395.

Wittner:2022:TTM

[Wit22] David Wittner. A tale of two mills: Socio-technological integration in Meiji Japan, 1868–1912. Technology and Culture, 63(2):349–376, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
854001.

Welk-Joerger:2023:MB

[WJ23] Nicole Welk-Joerger. Maintaining bovine. Technology and Culture, 64(4):1141–1158, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/910998.

Wolters:2023:DNP

[Wol23] Timothy S. Wolters. The Development of Nuclear Propulsion in the Royal Navy, 1946–1975 by Gareth Michael Jones (review). Technology and Culture, 64(4):1339–1340, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/911032.

Worthen:2024:SPP

[Wor24] W. B. Worthen. Staging, Playing, Pyrotechnics and Magic: Conventions of Performance in Early English Theatre by Philip Butterworth (review). Technology and Culture, 65(1):424–425, January 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
920564.

Wright:2021:NLC

[Wri21] Gavin Wright. NASA and the Long Civil Rights Movement ed. by Brian C. Odom and Stephen P. Waring (review). Technology and Culture, 62(1):301–302, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/783940.

Wu:2023:PNC

[Wu23] Jiarui Wu. Prototype Nation: China and the Contested Promise of Innovation by Silvia M. Lindtner (review). Technology and Culture, 64(3):980–981, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/904006.

Yamada:2023:DPE

[Yam23a] Keisuke Yamada. Decomposed: The Political Ecology of Music by Kyle Devine, and: Audible Infrastructures: Music, Sound, Media ed. by Kyle Devine and Alexandrine Boudreault-Fournier (review). Technology and Culture, 64(1):247–250, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/879600.

Yamada:2023:CEV

[Yam23b] Keisuke Yamada. Cover essay: Visual images in sound studies. Technology and Culture, 64(2):303–307, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/893039.

Yamada:2023:MCE

[Yam23c] Keisuke Yamada. Mobilizing citizens’ ears: Aural training as civil defense, 1941–45. Technology and Culture, 64(2):359–378, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893088.

Yang:2020:EVH

[Yan20] Daqing Yang. Electrified Voices: How the Telephone, Phonograph, and Radio Shaped Modern Japan, 1868–1945 by Kerim Yasar (review). Technology and Culture, 61(4):1233–1234, October 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/777636.

Yang:2022:DHA

[Yan22] Yimin Yang. Development History of Ancient Chinese Glass Technology ed. by Fuxi Gan (review). Technology and Culture, 63 (2):597–598, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/854047.

Yao:2021:IFP

[Yao21] Liang Yao. Information Fantasies: Precarious Mediation in Postsocialist China by Xiao Liu (review). Technology and Culture, 62(1):308–309, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/783944.

Yi:2022:CMP

[Yi22] Yuan Yi. Crafted for mass production: Imported spinning machinery on the shop floor, China, 1910s–1920s. Technology and Culture, 63(4):979–1004, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/868050.

Yoo:2020:IPT

[Yoo20] Sangwoon Yoo. Innovation in practice: The “Technology Drive Policy” and the 4Mb DRAM R&D consortium in South Korea in the 1980s and 1990s. Technology and Culture, 61(2):385–415, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
761570.

Yu:2021:MVC

[Yu21] Sarah Yu. Mass Vaccination: Citizens’ Bodies and State Power in Modern China by Mary Augusta Brazelton (review). Technology and Culture, 62(4):1237–1238, October 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/819809.

Zachary:2023:CAP

[Zac23] G. Pascal Zachary. Chasing Automation: The Politics of Technology and Jobs from the Roaring Twenties to the Great Society by Jerry Prout (review). Technology and Culture, 64(2):640–641, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
893082.

Zaidi:2020:PPC

[Zai20] Waqar Zaidi. Peace and Power in Cold War Britain: Media, Movements and Democracy, c. 1945–68 by Christopher R. Hill (review). Technology and Culture, 61(1):367–369, January 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/752941.

Zaidi:2023:BAI

[Zai23] Waqar H. Zaidi. The British Aircraft Industry and American-Led Globalisation, 1943–1982 by Takeshi Sakade (review). Technology and Culture, 64(3):955–956, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/903992.

Zakaria:2023:TFQ

[Zak23] Faizah Zakaria. Timber and Forestry in Qing China: Sustaining the Market by Meng Zhang (review). Technology and Culture, 64(2):647–648, April 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/893086.

Zanier:2022:ECL

[Zan22] Valeria Zanier. Enduring Change: The Labor and Social History of One Third-Front Industrial Complex in China from the 1960s to the Present by Ju Li (review). Technology and Culture, 63 (4):1203–1204, October 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/868062.

Zapico:2020:SRX

[Zap20] Daniel Pérez Zapico. Superrealistas. Rayos X y vanguardias artísticas by Ana Lamata Manuel (review). Technology and Culture, 61(3):975–976, July 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/763611.

Zdrodowska:2021:AAH

[Zdr21] Magdalena Zdrodowska. Accessible America: A History of Disability and Design by Bess Williamson (review). Technology and Culture, 62(3):954–956, July 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/803134.

Zeavin:2022:WUF

[Zea22] Hannah Zeavin. “This is womenspace”: USENET and the fight for a digital backroom, 1983–86. Technology and Culture, 63(3):634– 664, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
859719.

Zeller:2022:BTR

[Zel22] Thomas Zeller. Builders of the Third Reich: The Organisation Todt and Nazi Forced Labour by Charles Dick (review). Technology and Culture, 63(2):529–530, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.

jhu.edu/pub/1/article/854046.

Zeller:2023:ILR

[Zel23] Thomas Zeller. Imaging landscapes, roads, race, and power. Technology and Culture, 64(4):1261–1273, October 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/911005.

Zemmin:2021:RIC

[Zem21] Florian Zemmin. Rediscovering the Islamic Classics: How Editors and Print Culture Transformed an Intellectual Tradition by Ahmed El Shamsy (review). Technology and Culture, 62(1): 314–316, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/

pub/1/article/783948.

Zeng:2022:RCW

[Zen22] Zhaojin Zeng. Recharging China in War and Revolution, 1882– 1955 by Ying Jia Tan (review). Technology and Culture, 63(3):921– 923, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
859763.

Zhang:2020:FWJ

[Zha20] Lei Zhang. Foreign Wells: Japanese well drilling in Beijing, 1900– 10. Technology and Culture, 61(2):416–436, April 2020. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/761571.

Zhang:2021:UCH

[Zha21] Jun Zhang. Unending Capitalism: How Consumerism Negated China’s Communist Revolution by Karl Gerth (review). Technology and Culture, 62(1):318–319, January 2021. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://

muse.jhu.edu/pub/1/article/783950.

Ziakkas:2023:SSC

[Zia23] Dimitrios Ziakkas. Stealth: The Secret Contest to Invent Invisible Aircraft by Peter Westwick (review). Technology and Culture, 64(1):252–253, January 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.

edu/pub/1/article/879631.

Zeng-Jian:2022:LKC

[ZJ22] Guan Zeng-Jian. Lost Knowledge: The Concept of Vanished Technologies and Other Human Histories by Benjamin B. Olshin (review). Technology and Culture, 63(1):237–239, January 2022. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/844172.

Zumthurm:2022:CMC

[ZK22] Tizian Zumthurm and Stefan Krebs. Collecting middle-class memories? The pandemic, technology, and crowdsourced archives. Technology and Culture, 63(2):483–493, April 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL

https://muse.jhu.edu/pub/1/article/854010.

Zou:2022:ESA

[Zou22] Dongxin Zou. Economizing socialist aid: China’s failed surgical plant in Algeria, 1973–80. Technology and Culture, 63(3):718–748, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
859722.

Zhihui:2024:CSP

[ZP24] Zhang Zhihui and Cheng Peng. China’s Space Programme: From the Era of Mao Zedong to Xi Jinping by S. Chandrashekar (review). Technology and Culture, 65(2):731–733, April 2024. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/926349.

Zuleta:2024:PSI

[Zul24] María Cecilia Zuleta. Con los pies en el surco: Instituciones estatales y actores de la ciencia agropecuaria en La Pampa (1958– 1983) [With feet in the furrow: State institutions and actors of agricultural science in La Pampa (1958–1983)] by Federico Martocci (review). Technology and Culture, 65(2):679–681, April 2024. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (elec-

tronic). URL https://muse.jhu.edu/pub/1/article/926323.

Zylstra:2020:CTM

[Zyl20] Geoff D. Zylstra. Colored Travelers: Mobility and the Fight for Citizenship before the Civil War by Elizabeth Stordeur Pryor (review). Technology and Culture, 61(2):693–694, April 2020. CO-DEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic).

URL https://muse.jhu.edu/pub/1/article/761586.

Zylstra:2022:MMA

[Zyl22] Geoff D. Zylstra. Modern Mobility Aloft: Elevated Highways, Architecture, and Urban Change in Pre-Interstate America by Amy D. Finstein (review). Technology and Culture, 63(3):878–879, July 2022. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL https://muse.jhu.edu/pub/1/article/
859740.